### v1

# Vojta’s abc conjecture for entire curves in toric varieties highly ramified over the boundary

Abstract.
We prove Vojta’s abc conjecture for projective space ${\mathbb{P}}^{n}({\mathbb{C}})$, assuming that the entire curves in ${\mathbb{P}}^{n}({\mathbb{C}})$ are highly ramified over the coordinate hyperplanes. This extends the results of Guo Ji and the second-named author for the case $n=2$ (see [9]). We also explore the corresponding results for projective toric varieties. Consequently, we establish a version of Campana’s orbifold conjecture for finite coverings of projective toric varieties.
2020 Mathematics Subject Classification: Primary 30D35; Secondary 32Q45 and 32H30
The research of Min Ru is supported in part by Simon Foundations grant award #531604 and #521604. The second-named author was supported in part by Taiwan’s NSTC grant 113-2115-M-001-011-MY3.
1. Introduction
Let $X$ be an algebraic variety, and $D$ an effective divisor (possibly empty). The set $X\backslash D$ is Brody hyperbolic if every entire curve $f:\mathbb{C}\to X\backslash D$ is constant. It is Brody quasi-hyperbolic if every entire curve $f:\mathbb{C}\to X\backslash D$ is degenerate, meaning its image is contained a proper Zariski-closed subset $Z$ of $X$. We call such $Z$ the exceptional set. If $Z$ is independent of $f$, we call $X\backslash D$ Brody strong quasi-hyperbolic. This leads us to the Griffiths-Lang’s conjecture.
Griffiths-Lang’s conjecture. Let $X$ be a smooth complex projective variety, $D$ a normal crossing divisor on $X$, and $K_{X}$ the canonical divisor on $X$. If $K_{X}+D$ is big, then $X\backslash D$ is Brody quasi-hyperbolic.
A more precise quantitative version, in terms of the Nevanlinna’s functions, can be formulated as follows.
Griffiths’ conjecture. For a holomorphic map $f:\mathbb{C}\to X$ with Zariski dense image, the inequality
$T_{K_{X}+D,f}(r)\leq_{\rm exc}N_{f}(D,r)+\epsilon T_{A,f}(r)$
holds for any $\epsilon>0$, where notation $\leq_{\rm exc}$ means the inequality holds for all $r\in(0,\infty)$ except for a set of finite measure.
Here, for each positive integer $n$, $N_{f}^{(n)}(D,r)$ is the $n$-truncated counting function with respect to $D$ given by
$N_{f}^{(n)}(D,r)=\sum_{0<|z|<r}\mathop{\mathrm{min}}\{{\rm ord}_{z}f^{*}D,n\}\log\frac{r}{|z|}+\mathop{\mathrm{min}}\{{\rm ord}_{0}f^{*}D,n\}\log r,$
$N_{f}(D,r)$ is the counting function with respect to $D$ given by $N_{f}^{(\infty)}(D,r)$, and $T_{D,f}(r)$ is the (Nevanlinna) height function relative to the divisor $D$ (referring to [25, Section 12]).
Paul Vojta formulated the following general abc conjecture (see [25, Conjecture 15.2] and [25, Conjecture 23.4]), which extends the above Griffiths’ conjecture. It is also viewed as the quantitive version of Campana’s orbifold conjecture which will be stated later in the introduction.
Vojta’s abc conejcture. For any $\epsilon>0$, there exists a proper Zariski-closed subset $Z$ of $X$ such that for any entire curve $f:\mathbb{C}\to X$ whose image is not contained in $Z$, the inequality
$T_{K_{X}+D,f}(r)\leq_{\rm exc}N_{f}^{(1)}(D,r)+\epsilon T_{A,f}(r)$
holds, where $A$ is an ample divisor on $X$.
Note that Vojta’s abc conjecture strengthens Griffiths’ conjecture in two key ways: (i) the exceptional set is independent of the map $f$, and (ii) the counting function is truncated at level one. Proving these two features is quite challenging, even if the aforementioned version of Griffiths’ conjecture could be established. In this paper, we first prove Vojta’s abc conjecture for entire curves in ${\mathbb{P}}^{n}({\mathbb{C}})$, assuming the map is highly ramified over the coordinate hyperplanes. The precise statement of our main result is as follows. Additionally, we highlight that the exceptional set obtained in the theorem can be determined explicitly-this is a challenging issue in the study of Nevanlinna theory and complex hyperbolicity.
Main Theorem.
Let $G$ be a non-constant homogeneous polynomial in $\mathbb{C}[x_{0},\dots,x_{n}]$ with no monomial factors and no repeated factors. Assume that $[G=0]$ is in general position with the coordinate hyperplanes $H_{i}=[x_{i}=0]$ in $\mathbb{P}^{n}(\mathbb{C})$, $0\leq i\leq n$. Then, for any $\epsilon>0$, there exist an effectively computable positive integer $\ell$ and a proper Zariski closed subset $Z$ of $\mathbb{P}^{n}({\mathbb{C}})$ that can be explicitly determined, such that the following two inequalities hold:
  • (i)
    $N_{G(\mathbf{f})}(0,r)-N^{(1)}_{G(\mathbf{f})}(0,r)\leq_{\operatorname{exc}}\epsilon T_{f}(r)$, and
  • (ii)
    $(\deg G-\epsilon)T_{f}(r)\leq_{\operatorname{exc}}N^{(1)}_{G(\mathbf{f})}(0,r)$,
for any non-constant holomorphic map $f:{\mathbb{C}}\rightarrow{\mathbb{P}}^{n}({\mathbb{C}})$ whose image is not contained in $Z$, and where ${\rm mult}_{z_{0}}(f^{*}H_{i})\geq\ell$ for all $i$ with $0\leq i\leq n$ and for every $z_{0}\in{\mathbb{C}}$ with $f(z_{0})\in H_{i}$. Here, the evaluation of $G$ is taken at $\mathbf{f}=(f_{0},\ldots,f_{n})$ in a reduced form of $f$, meaning that $f_{0},\ldots,f_{n}$ are entire functions without common zeros.
Note that the case when $n=2$ is due to Ji Guo and the second named author ([9, Theorem 1.4]). The new idea in this paper is motivated by the recent paper [2] by considering the parabolic Riemann surface $\mathcal{Y}_{f}:=\mathbb{C}\setminus{f^{-1}(D)}$, reducing the problem to the omitting case. To achieve this, we need to extend the results in [9] and [7] to the parabolic setting.
We state some consequences of our Main Theorem, starting with the introduction of Campana’s orbifolds. Instead of considering the complement case in Brody’s hyperbolicity, we can also examine an entire curve $f:\mathbb{C}\to X$ that is ramified over $D$. According to Campana [1], an orbifold divisor is defined as $\Delta:=\sum_{Y\subset X}(1-m_{\Delta}^{-1}(Y))\cdot Y$, where $Y$ ranges over all irreducible divisors of $X$, and $m_{\Delta}(Y)\in[1,\infty)\cap{\mathbb{Q}}$, with all but finitely many values equal to one. An orbifold entire curve $f:{\mathbb{C}}\rightarrow(X,\Delta)$, with $\Delta=\sum_{i}(1-m_{i}^{-1})D_{i}$, is a homomorphic map such that $f({\mathbb{C}})\not\subset\mbox{Supp}(\Delta)$ and ${\rm mult}_{z_{0}}(f^{*}D_{i})\geq m_{i}$ for all $i$ and for every $z_{0}\in{\mathbb{C}}$ with $f(z_{0})\in D_{i}$.
Campana’s orbifold conjecture. If $(X,\Delta$) is an orbifold pair such that $K_{X}+\Delta$ is big, then there exists a proper closed subvariety $Z\subset X$ containing all the non-constant orbifold entire curves $f:{\mathbb{C}}\rightarrow(X,\Delta)$.
Remarks.
  1. (i)
    The Campana’s conjecture is a consequence of the above Vojta’s strong abc conecutre.
  2. (ii)
    Let $D:={\rm Supp}(\Delta)=D_{1}+\cdots+D_{q}$. Recall that the pair $(X,D)$ is said to be of log-general type if the divisor $K_{X}+D$ is big. It’s clear that $(X,D)$ is of log-general type if $(X,\Delta)$ is of general type, i.e. $K_{X}+\Delta$ is big. On the other hand, the condition that $(X,D)$ is of log-general type implies that $(X,\Delta)$ is of general type if $(X,\Delta)$ has a sufficiently large multiplicity along each $D_{i}$, $1\leq i\leq q$. ((((See [12, Corollary 2.2.24].))))
The Main Theorem leads to the following two theorems concerning Campana’s orbifold conjecture; however, we will not include the proofs, as they can be adapted from those in [9] for the case $n=2$. We begin with the result concerning $n+2$ components.
Theorem 1.
Let $\Delta_{0}$ be an orbifold divisor of ${\mathbb{P}}^{n}({\mathbb{C}})$ and let $H_{0},H_{1},\dots,H_{n}$ be the coordinate hyperplanes in ${\mathbb{P}}^{n}({\mathbb{C}})$, such that $\Delta_{0}$ and $H_{0},H_{1},\dots,H_{n}$ are in general position. Let $m_{i}\in(1,\infty)\cap{\mathbb{Q}}$, $0\leq i\leq n$, and $\Delta=\Delta_{0}+(1-{1\over m_{0}})H_{0}+(1-{1\over m_{1}})H_{1}+\cdots+(1-{1\over m_{n}})H_{n}$. Assume that $\deg\Delta>n+1$. Then there exist a proper Zariski closed subset $W$ of ${\mathbb{P}}^{n}({\mathbb{C}})$ and an effectively computable positive integer $\ell$ such that the image of any orbifold entire curve $f:{\mathbb{C}}\rightarrow({\mathbb{P}}^{n}({\mathbb{C}}),\Delta)$ with $\mathop{\mathrm{min}}\{m_{0},m_{1},\dots,m_{n}\}\geq\ell$ must be contained in $W$.
The next case involves $n+1$ components, not all lines, in $\mathbb{P}^{n}({\mathbb{C}})$.
Theorem 2.
Let $F_{i}$, $11\leq i\leq n+$ 1, be homogeneous irreducible polynomials of positive degrees in $\mathbb{C}[x_{0},\ldots,x_{n}]$. Assume that $D_{i}:=[F_{i}=0]\subset\mathbb{P}^{n}(\mathbb{C})$, $11\leq i\leq n+$ 1, intersect transversally. Let $\Delta=(1-{1\over m_{1}})D_{1}+\cdots+(1-{1\over m_{n+1}})D_{n+1}$ be an orbifold divisor of $\mathbb{P}^{n}({\mathbb{C}})$. Assume that $\deg\Delta>n+1$. Then, there exist an effectively computable positive integer $\ell$ and a proper Zariski closed subset $W$ of $\mathbb{P}^{n}({\mathbb{C}})$ such that the image of any orbifold entire curve $f:{\mathbb{C}}\rightarrow({\mathbb{P}}^{n}({\mathbb{C}}),\Delta)$ with $\mathop{\mathrm{min}}\{m_{1},\dots,m_{n+1}\}\geq\ell$ must be contained in $W$.
Note that $\mathbb{P}^{n}({\mathbb{C}})$ is a toric variety when we identify ${\mathbb{G}}_{m}^{n}$ with ${\mathbb{P}}^{n}({\mathbb{C}})\backslash\cup_{i=0}^{n}H_{i}$, where $H_{0},\dots,H_{n}$ are the $n+1$ coordinate hyperplanes. A toric variety $X$ is defined as a variety over ${\mathbb{C}}$ that contains ${\mathbb{G}}_{m}^{n}$ as a dense open subvariety, with the action of ${\mathbb{G}}_{m}^{n}$ on itself extends to an algebraic action of ${\mathbb{G}}_{m}^{n}$ on $X$. We refer to [5] for more details on toric varieties. Using the arguments in [13] and [6], we can extend our Main Theorem to projective toric varieties. Recall from [13] that an “admissible pair” is a couple $(X,V)$ where $V$ is a nonsingular variety embedded in a nonsingular projective variety $X$ in such a way that $D_{0}=X\setminus V$ is a normal crossings divisor. A Campana orbifold $(X,\Delta)$ is said to be associated to the admissible pair $(X,V)$ if $D_{0}$ is the support of $\Delta$.
Theorem 3.
Let $X$ be a nonsingular projective toric variety and let $D$ be an effective reduced divisor on the admissible pair $(X,\mathbb{G}_{m}^{n})$, such that $D$ is in general position with the boundary $D_{0}=X\setminus\mathbb{G}_{m}^{n}$. Let $A$ be a big divisor on $X$. For every $\epsilon>0$ there is a positive integral number $\ell$ and a proper Zariski-closed subset $Z$ of $X$ such that, for any Campana orbifold $(X,\Delta)$ associated to the admissible pair $(X,\mathbb{G}_{m}^{n})$ with multiplicities at least $\ell$ along the boundary $D_{0}$, the following two inequalities hold:
  1. (i)
    $N_{f}(D,r)-N_{f}^{(1)}(D,r)\leq_{\rm exc}\epsilon T_{A,f}(r)$, and
  2. (ii)
    $T_{D,f}(r)\leq_{\rm exc}N_{f}^{(1)}(D,r)+\epsilon T_{A,f}(r)$
for any non-constant entire orbifold curve $f$ on $(X,\Delta)$ such that the image of $f$ is not contained in $Z$.
Theorem 3 implies Campana’s conjecture for an orbifold pair that is a finite covering of a toric variety, as follows.
Theorem 4.
Let $X$ be a nonsingular projective toric variety, and let $D_{0}=X\setminus\mathbb{G}_{m}^{n}$. Let $Y$ be a nonsingular complex projective variety with a finite morphism $\pi:Y\to X$. Let $H:=\pi^{*}(D_{0})$ and $R\subset Y$ the ramification divisor of $\pi$ omitting components from the support of $H$. Assume that $\pi(R)$ and $D_{0}$ are in general position on $X$. Then there exists a positive integer $\ell$ such that if the orbifold pair $(Y,\Delta)$ is of general type with ${\rm Supp}(\Delta)={\rm Supp}(H)$ and multiplicity at least $\ell$ for each component, then there exists a proper Zariski closed subset $W$ of $Y$ such that the image of any non-constant orbifold entire curve $f:{\mathbb{C}}\rightarrow(Y,\Delta)$ must be contained in $W$.
In Section 2, we prove a version of the Main Theorem for rational maps and explain the motivation for using parabolic Riemann surfaces. Section 3 reviews basic results from Nevanlinna theory relevant to parabolic settings, as well as the key theorems from [2]. Furthermore, we extend the GCD theorem developed in [14] and in [9] to parabolic setting. In Section 4, we establish the abc theorem for parabolic Riemann surfaces, building on the methods developed in [7]. Section 5 presents the proof of the Main Theorem, while Section 6 extends results to toric varieties.
2. Motivation through the rational case
In [7], a version of Vojta’s abc conjecture for algebraic tori over algebraic function fields is established. As we will see, the algebraic reduction method presented in this article is a key technical component in proving the Main Theorem. The core idea behind the proof is also inspired by the function field approach. In this section, we will prove a version of the Main Theorem for rational functions and explain the role of parabolic Riemann surfaces in this context.
2.1. The Rational Case
The statements and arguments apply verbatim to algebraic function fields. Here, let $K=\mathbb{C}(t)$ be the function field of $\mathbb{P}^{1}({\mathbb{C}})$. Let $S$ be a finite subset of $\mathbb{P}^{1}(\mathbb{C})$. We denote by ${\mathcal{O}}_{S}^{*}$ the set of $S$-units, which consists of rational functions with no zeros and poles outside of $S$. For a rational function $f=\frac{P}{Q}$, where $P$ and $Q$ are coprime polynomials, we define $\deg f:=\max\{\deg P,\deg Q\}$ and denote by ${\rm ord}_{\mathbf{p}}^{+}(f)$ the order of the zero of $f$ at $\mathbf{p}\in\mathbb{P}^{1}(\mathbb{C})$. We use the notation $N_{S}(f)$ (respectively $N_{S}^{(1)}(f)$) to count the number of zeros of $f$ with multiplicity (respectively, without multiplicity) outside of $S$. When $S$ is empty, we simply write $N(f)$ (respectively $N^{(1)}(f)$) to count the number of zeros with multiplicity (respectively, without multiplicity) on $\mathbb{P}^{1}(\mathbb{C})$.
Recall a simplified version of the main technical theorem in [7].
Theorem 5 ([7, Theorem 4]).
Let $G\in\mathbb{C}[x_{1},\ldots,x_{n}]$ be a non-constant homogeneous polynomial with neither monomial factors nor repeated factors. Let $S$ be a finite subset of $\mathbb{P}^{1}(\mathbb{C})$. Then, for any $\epsilon>0$, there exist a positive real number $c_{0}$ and a proper Zariski closed subset $Z$ of $\mathbb{A}^{n}$ such that, for all $\mathbf{u}:=(u_{1},\ldots,u_{n})\in({\mathcal{O}}_{S}^{*})^{n}\setminus Z$, we have either
  1. (i)
    the inequality
    $\max_{1\leq i\leq n}\{\deg(u_{i})\}\leq c_{0}\max\{1,|S|-2\}$
    holds,
  2. (ii)
    or the following two statements hold
    1. (a)
      $N_{S}(G(\mathbf{u}))-N_{S}^{(1)}(G(\mathbf{u}))\leq\epsilon\max_{1\leq i\leq n}\{\deg u_{i}\}$ if $G(u_{1},\ldots,u_{n})\neq 0$.
    2. (b)
      If $G(0,\ldots,0)\neq 0$ and $\deg_{X_{i}}G=\deg G=d$ for $1\leq i\leq n$, then
      $N_{S}^{(1)}(G(\mathbf{u}))\geq(\deg G-\epsilon)\max_{1\leq i\leq n}\{\deg u_{i}\}.$
Here $c_{0}$ can be effectively bounded from above in terms of $\epsilon$, $n$, and the degree of $G$. Moreover, the exceptional set $Z$ can be expressed as the zero locus of a finite set $\Sigma\subset\mathbb{C}[x_{1},\ldots,x_{n}]$ with the following properties: (Z1) $\Sigma$ depends on $\epsilon$ and $G$ and can be determined explicitly and (Z2) $|\Sigma|$ and the degree of each polynomial in $\Sigma$ can be effectively bounded from above in terms of $\epsilon$, $n$, and the degree of $G$.
We will demonstrate how to apply the above theorem to show the following version of Main Theorem for rational case.
Theorem 6.
Let $G$ be a non-constant homogeneous polynomial in $\mathbb{C}[x_{0},\cdots,x_{n}]$ with no monomial factors and no repeated factors. Assume $[G=0]$ is in general position with the coordinate hyperplanes $H_{i}=[x_{i}=0]$, $0\leq i\leq n$. Then, for any $\epsilon>0$, there exist effectively computable positive integers $\ell$ and a proper Zariski closed subset $Z$ of $\mathbb{P}^{n}({\mathbb{C}})$ such that the following two inequalities hold:
  • (i)
    $N(G(\mathbf{f}))-N^{(1)}(G(\mathbf{f}))\leq\epsilon\deg\mathbf{f}$, and
  • (ii)
    $N^{(1)}(G(\mathbf{f}))\geq(\deg G-\epsilon)\deg\mathbf{f}$,
for any non-constant entire rational curve $\mathbf{f}=(f_{0},\ldots,f_{n})$, where $f_{0},\ldots,f_{n}\in\mathbb{C}[z]$ are polynomials without common zeros, provided that $\mathbf{f}(\mathbb{C})\not\subset Z$ and ${\rm ord}_{\mathbf{p}}(f_{i})\geq\ell$ if ${\rm ord}_{\mathbf{p}}(f_{i})>0$ for $0\leq i\leq n$, and for all $\mathbf{p}\in\mathbb{P}^{1}({\mathbb{C}})$. Here, $\deg\mathbf{f}=\max\{\deg f_{0},\ldots,\deg f_{n}\}$.
In addition to Theorem 5, we also require the following result to prove Theorem 6.
Proposition 7 ([7, Corollary 6]).
Let $G=\sum_{\mathbf{i}\in I_{G}}a_{\mathbf{i}}{\mathbf{x}}^{\mathbf{i}}\in\mathbb{C}[x_{1},\ldots,x_{n}]$ be a non-constant polynomial, where $a_{\mathbf{i}}\neq 0$ if $\mathbf{i}\in I_{G}$. Assume that $G(0,\dots,0)\neq 0$. Let $Z$ be the Zariski closed subset that is the union of hypersurfaces of $\mathbb{A}^{n}$ of the form $\sum_{\mathbf{i}\in I}a_{\mathbf{i}}{\mathbf{x}}^{\mathbf{i}}=0$ where $I$ is a non-empty subset of $I_{G}$. Let $S$ be a finite subset of $\mathbb{P}^{1}(\mathbb{C})$. Then, for all $(u_{1},\ldots,u_{n})\in(\mathcal{O}_{S}^{*})^{n}\setminus Z$, we have
$\sum_{\mathbf{p}\in S}{\rm ord}_{\mathbf{p}}^{+}(G(u_{1},\ldots,u_{n}))\leq\tilde{c}\max\{1,|S|-2\}$
where $\tilde{c}=\frac{1}{2}\binom{n+\deg G}{n}(\binom{n+\deg G}{n}+1)$.
Proof of Theorem 6.
Denote by $\ell$ a sufficiently large integer to be determined later. Let $\mathbf{f}=(f_{0},\ldots,f_{n})$, where $f_{0},\ldots,f_{n}\in\mathbb{C}[z]$ are polynomials without common zeros such that ${\rm ord}_{\mathbf{p}}(f_{i})\geq\ell$ if ${\rm ord}_{\mathbf{p}}(f_{i})>0$ for $0\leq i\leq n$ and $\mathbf{p}\in{\mathbb{P}}^{1}(\mathbb{C})$. Let
(2.1) $\tilde{S}_{\mathbf{f}}:=\{\mathbf{p}\in\mathbb{P}^{1}(\mathbb{C})\,|\,{\rm ord}_{\mathbf{p}}(f_{i})>0\quad\text{for some $0\leq i\leq n$}\}\cup\{\infty\}.$
Then
(2.2) $|\tilde{S}_{\mathbf{f}}|\leq\sum_{i=0}^{n}N_{S}^{(1)}(f_{i})+1\leq\frac{n+1}{\ell}\max_{0\leq i\leq n}\{\deg f_{i}\}+1,$
and $f_{i}\in\mathcal{O}_{\tilde{S}_{\mathbf{f}}}^{*}$ for $0\leq i\leq n$. Let $\epsilon>0$ be given. We now apply Theorem 5 to find a positive integer $c_{0}$ and a proper Zariski closed subset $Z_{0}$ independent of $\tilde{S}_{\mathbf{f}}$ of $\mathbb{A}^{n}$ such that we have either the inequality
(2.3) $\max_{0\leq i\leq n}\{\deg f_{i}\}(=\deg\mathbf{f})\leq c_{0}\max\{1,|\tilde{S}_{\mathbf{f}}|-2\}$
holds or the following two inequalities hold:
(2.4) $N_{\tilde{S}_{\mathbf{f}}}(G_{0}(u_{1},\ldots,u_{n}))-N_{\tilde{S}_{\mathbf{f}}}^{(1)}(G_{0}(u_{1},\ldots,u_{n}))\leq\epsilon\deg\mathbf{f},$
(2.5) $N_{\tilde{S}_{\mathbf{f}}}^{(1)}(G_{0}(u_{1},\ldots,u_{n}))\geq(\deg G-\epsilon)\cdot\deg\mathbf{f},$
if $\mathbf{f}(\mathbb{C})\not\subset\tilde{Z}_{0}$, where $\tilde{Z}_{0}$ is the projective closure of $Z_{0}$ in $\mathbb{P}^{n}({\mathbb{C}})$. Here $G_{0}(x_{1},\ldots,x_{n}):=G(1,x_{1},\ldots,x_{n})$ and $u_{i}=\frac{f_{i}}{f_{0}}\in\mathcal{O}_{\tilde{S}_{\mathbf{f}}}^{*}$, $1\leq i\leq n$. Since the constant $c_{0}$ and the exceptional set $Z$ in Theorem 5 do not depend on $S$, it is important to note that $c_{0}$ and $Z_{0}$ do not depend on $\mathbf{f}$ as well. We first show that (2.3) does not hold if $\ell>2c_{0}n$. Indeed, if (2.3) holds with $\ell>2c_{0}n$, then it follows from (2.2) and the choice of $\ell$ that (2.3) yields
$\max_{0\leq i\leq n}\{\deg f_{i}\}\leq\frac{(n+1)c_{0}}{\ell}\max_{0\leq i\leq n}\{\deg f_{i}\}+c_{0}<\frac{1}{2}\max_{0\leq i\leq n}\{\deg f_{i}\}+c_{0}.$
Therefore, $\max_{0\leq i\leq n}\{\deg f_{i}\}<2c_{0}.$ This is not possible since not all the polynomials $f_{i}$ are constant, the assumption that ${\rm ord}_{\mathbf{p}}(f_{i})\geq\ell$ if ${\rm ord}_{\mathbf{p}}(f_{i})>0$ for $0\leq i\leq n$ and $\mathbf{p}\in{\mathbb{P}}^{1}(\mathbb{C})$ implies that $2nc_{0}<\ell<\max_{0\leq i\leq n}\{\deg f_{i}\}$.
We continue to let $\ell>2c_{0}n$. As we have shown that (2.3) does not hold, it follows that (2.4) and (2.5) both hold if $\mathbf{f}(\mathbb{C})\not\subset\tilde{Z}_{0}$. Since $N^{(1)}(G_{0}(u_{1},\ldots,u_{n}))\geq N_{\tilde{S}_{\mathbf{f}}}^{(1)}(G_{0}(u_{1},\ldots,u_{n}))$, it follows trivially from (2.5) that
(2.6) $N^{(1)}(G(\mathbf{f}))\geq N^{(1)}(G_{0}(u_{1},\ldots,u_{n}))\geq(\deg G-\epsilon)\cdot\deg\mathbf{f}.$
This shows (ii) holds.
It remains to show (i). We now compute zeros for points in $\tilde{S}_{\mathbf{f}}$. We will use Propostion 7. Let $W$ be the Zariski closed subset of $\mathbb{P}^{n}(\mathbb{C})$ containing all hypersurfaces defined by all possible subsums (including $G$) in the expansion of $G$. Assume that $\mathbf{f}(\mathbb{C})\not\subset W$. Since $G_{0}(0,\dots,0)=G(1,0,\ldots,0)\neq 0$, we now apply Propositon 7 to get
$\sum_{\mathbf{p}\in\tilde{S}_{\mathbf{f}}}{\rm ord}^{+}_{\mathbf{p}}(G_{0}(u_{1},\ldots,u_{n}))\leq\tilde{c}\max\{1,|\tilde{S}_{\mathbf{f}}|-2\},$
where $\tilde{c}=\frac{1}{2}\binom{n+degG}{n}(\binom{n+\deg G}{n}+1)$. Together with (2.2) and that $\ell\leq\max_{0\leq i\leq n}\{\deg f_{i}\}$, it yields
$m_{\tilde{S}_{\mathbf{f}}}(G_{0}(u_{1},\ldots,u_{n})):=\sum_{\mathbf{p}\in\tilde{S}_{\mathbf{f}}}{\rm ord}^{+}_{\mathbf{p}}(G_{0}(u_{1},\ldots,u_{n}))\leq\frac{(n+1)\tilde{c}}{\ell}\max_{0\leq i\leq n}\{\deg f_{i}\}\leq\epsilon\deg\mathbf{f}$
provided $\ell>(n+1)\tilde{c}_{0}\epsilon^{-1}.$ Combining with (2.4), we have
(2.7) $N(G_{0}(u_{1},\ldots,u_{n}))-N^{(1)}(G_{0}(u_{1},\ldots,u_{n}))\leq 2\epsilon\deg\mathbf{f}.$
We now repeat the above arguments for $(\frac{f_{0}}{f_{i}},\cdots,\frac{f_{n}}{f_{i}})$, $0\leq i\leq n$. Then we find a Zariski closed subset $Z$ and a positive constant $c$, independent of $\mathbf{f}$, such that for each $0\leq i\leq n$ we have
(2.8) $N\left(G(\frac{f_{0}}{f_{i}},\cdots,\frac{f_{n}}{f_{i}})\right)-N^{(1)}\left(G(\frac{f_{0}}{f_{i}},\cdots,\frac{f_{n}}{f_{i}})\right)\leq 2\epsilon\deg\mathbf{f}.$
Since $f_{0},\ldots,f_{n}$ are polynomials without common zeros, for each $z\in\mathbb{C}$ there exists at least one $f_{i}$ such that $f_{i}(z)\neq 0$. Then
${\rm ord}^{+}_{z}(G(f_{0},\ldots,f_{n}))={\rm ord}_{z}\left(G(\frac{f_{0}}{f_{i}},\cdots,\frac{f_{n}}{f_{i}})\right).$
This shows that
$N(G(\mathbf{f}))-N^{(1)}(G(\mathbf{f}))$ $\leq\sum_{i=0}^{n}\left(N\left(G(\frac{f_{0}}{f_{i}},\cdots,\frac{f_{n}}{f_{i}})\right)-N^{(1)}\left(G(\frac{f_{0}}{f_{i}},\cdots,\frac{f_{n}}{f_{i}})\right)\right)$
[…]
The assertion (i) can be derived by adjusting $\epsilon$. ∎
2.2. Motivation for the proof of the Main Theorem
As in Section 2.1, the starting point of the proof of Theorem 6 is as the follows: For a given rational map, we define a set $S$ (depending on the map) that contains the zeros and poles of each coordinate function of the map. In this way, all the coordinate functions become $S$-units, and so we can apply Theorem 5. Importantly, the constants and the exceptional set obtained in Theorem 5 do not depend on $S$, allowing this method to be used for all rational maps. To apply this view point to the analytic case, we consider $f=(f_{0},\ldots,f_{n})$, where $f_{i}$, $0\leq i\leq n$, are entire functions without common zeros. Let $\mathcal{E}_{f}$ denote the set of all the zeros of $f_{i}$, $0\leq i\leq n$, and define $\mathcal{Y}_{f}:=\mathbb{C}\setminus\mathcal{E}_{f}$. Then ${\mathcal{Y}_{f}}$ is a parabolic open Riemann surface. On ${\mathcal{Y}_{f}}$, the restriction $\tilde{f}_{i}:=f_{i}|_{\mathcal{Y}_{f}}$ becomes a unit on ${\mathcal{Y}}$ (i.e. it has neither zeros nor poles). This perspective was also adopted in [2], where they laid important foundational work. For our purposes, we will need to develop an analogue of Theorem 5 over parabolic open Riemann surfaces.
3. Nevanlinna theory and the GCD theorem on parabolic Riemann surfaces
3.1. Parabolic Riemann surfaces
[…]
1. For a hypersurface $D=\{Q=0\}$ in $\mathbb{P}^{n}({\mathbb{C}})$, where $Q$ is a homogeneous polynomial of degree $d$, the proximity function is defined as
[…]
where $\lambda_{D}$ is the Weil-function on ${\mathbb{P}}^{n}({\mathbb{C}})$ given by
$\lambda_{D}({\bf x})=\log{\|{\bf x}\|^{d}\over|Q({\bf x})|},~{}~{}{\bf x}\in\mathbb{P}^{n}({\mathbb{C}}).$
2. The characteristic (order) function is defined by
$T_{f,{\mathcal{Y}}}(r)=\int_{1}^{r}{dt\over t}\int_{B(t)}f^{*}\omega_{FS}.$
3. The counting function and the $k$-truncated counting function are defined by
$N_{f,{\mathcal{Y}}}(D,r):=\int_{1}^{r}\sum_{z\in B(t)}{\rm ord}_{z}f^{*}D{dt\over t},~{}~{}~{}~{}~{}~{}~{}\mbox{and}~{}~{}N_{f,{\mathcal{Y}}}^{(k)}(D,r):=\int_{1}^{r}\sum_{z\in B(t)}\mathop{\mathrm{min}}\{k,{\rm ord}_{z}f^{*}D\}{dt\over t}.$
4. The counting function extends naturally to the sub-schemes, as well as the notion $N_{\text{gcd},{\mathcal{Y}}}(f,g)(r)$, where $f,g$ are meromorphic functions on $\mathcal{Y}$. See [14].
Parabolic First Main Theorem. Let $D$ be a hypersurface in ${\mathbb{P}}^{n}({\mathbb{C}})$ and $f:{\mathcal{Y}}\rightarrow\mathbb{P}^{n}({\mathbb{C}})$ be a holomorphic map with $f({\mathcal{Y}})\not\subset D$. Then
$T_{f,{\mathcal{Y}}}(r)=m_{f,{\mathcal{Y}}}(D,r)+N_{f,{\mathcal{Y}}}(D,r)+O(1).$
It is important to note that, when we write $O(g(x))$ for some function $g$, we mean that $Cg(x)$ where $C$ is a constant independent of ${\mathcal{Y}}$ throughout the paper.
Fix a nowhere vanishing global holomorphic vector field $\xi\in\Gamma({\mathcal{Y}},T_{\mathcal{Y}})$ (we denote it by ${\partial\over\partial z}$) on ${\mathcal{Y}}$. Such a vector field exists because ${\mathcal{Y}}$ is non-compact and consequently its holomorphic tangent bundle $T_{\mathcal{Y}}$ is holomorphically trivial. From Lemma 4.4 in [10],
$\int_{S_{r}}\log\left|d\sigma\left({\partial\over\partial z}\right)\right|^{2}d\mu_{r}=-\int_{1}^{r}\chi_{\sigma}(t)\frac{dt}{t}+2\varsigma\log r+O(1),$
where $\chi_{\sigma}(r)$ is the Euler characteristic of $B(r)$, and
[…]
which is independent of $r$ for $r$ being sufficiently large. Denote by
$\mathfrak{X}^{+}_{\sigma}(r):=\int_{S(r)}\log^{+}\left|d\sigma\left({\partial\over\partial z}\right)\right|^{2}d\mu_{r}.$
[…]
(3.1) $\hat{\tau}:=\log\hat{\sigma}:=\log^{+}|z|+\sum_{j\geq 1}r_{j}\log^{+}{r_{j}\over|z-a_{j}|}.$
Note that $\hat{\tau}$ takes values in $[0,+\infty)$, is continuous on ${\mathcal{Y}}$ and is smooth (indeed harmonic) outside the circle $S(0,1):=\{|z|=1\}$ and the disjoint circles $S(a_{j},r_{j}):=\{|z-a_{j}|=r_{j}\}$. By the Poincare-Lelong formula, it is clear that
$dd^{c}\hat{\tau}={1\over 2}\nu(0,1)+{r_{j}\over 2}\sum_{j=1}^{\infty}\left(\nu(a_{j},r_{j})-\delta_{a_{j}}\right)$
is a distribution of order 00 and locally of finite mass, where $\nu(a_{j},r_{j})$ is the Haar measure on the circle $S(a_{j},r_{j})$. The smoothing process was done in [2] as follows: as in the Appendix in [2], let
$h(r)=\begin{cases}0&\text{ for $r\leq{3\over 4}$}\\ {1\over 4\pi}{1\over 1+e^{{r-1\over(r-1)^{2}-1/16}}}&\text{for ${3\over 4}<r<{5\over 4}$}\\ {1\over 4\pi}&\text{for $r\geq{5\over 4}$}\end{cases}$
and
$H(r):=4\pi\int_{0}^{r}h(s-c){ds\over s},$
where the constant $c\in(0,1/2)$ is chosen such that $H$ agrees with $\log^{+}r$ outside $[{1\over 2},{3\over 2}]$. Note that $H$ is an absolute function independent of $a_{j}$. This function $H(r)$ is smooth of $\log^{+}r$ with $0\leq H(r)-\log^{+}r\leq\log^{+}{3\over 2}<{1\over 2}$ and $H(r)=\log^{+}r$ when $H(r)\geq\log{3\over 2}$. Define
$\tau:=\log\sigma:=H(z)+\sum_{j=1}^{\infty}r_{j}H\left({r_{j}\over|z-a_{j}|}\right).$
Then $\sigma\geq\hat{\sigma}$ and the difference $\sigma-\hat{\sigma}$ is supported on
$\text{Supp}(\sigma-\hat{\sigma})\subset U:=\left(A\left(0,{1\over 2},{3\over 2}\right)\backslash{\mathcal{E}}\right)\bigcup_{j=1}^{\infty}A\left(a_{j},{1\over 2}r_{j},{3\over 2}r_{j}\right),$
where $A(a_{j},{1\over 2}r_{j},{3\over 2}r_{j}):=\{z\in{\mathcal{Y}}:{1\over 2}r_{j}\leq|z-a_{j}|\leq{3\over 2}r_{j}\}$ are pairwise disjoint annuli. For $z\not\in\cup_{j=1}^{\infty}\overline{{\mathbb{D}}(a_{j},{3\over 2}r_{j})}$, $\log^{+}\left|{r_{j}\over z-a_{j}}\right|=H\left(\left|{r_{j}\over z-a_{j}}\right|\right)=0$ for each $j$. Thus $\tau(z)=H(|z|)$ and $\hat{\tau}(z)=\log^{+}|z|$, which are eqaual when $\tau(z)\geq\log{3\over 2}$. Let $B_{r}^{\sigma}:=\{z\in{\mathcal{Y}}~{}|~{}\sigma(z)<r\}$ be the $\sigma$-ball of radius $r$. Then the above discussion implies that
(3.2) $B_{t}^{\sigma}\subset{\mathbb{D}}_{t}~{}~{}~{}~{}(t\geq 1)~{}~{}~{}\mbox{and}~{}~{}~{}~{}~{}{\mathbb{D}}_{t}\backslash\left(\bigcup_{j=1}^{\infty}{\mathbb{D}}(a_{j},{3\over 2}r_{j})\right)\subset B_{t}^{\sigma}~{}~{}~{}~{}(t\geq{3\over 2}).$
We now show that the smooth 2-form $dd^{c}\sigma$ defines an order 0 distribution of finite mass on ${\mathcal{Y}}$ (see Lemma 3.2 in [2]) so that Parabolic Logarithmic Derivative Lemma (Lemma 9) holds for $\sigma$. We use the formula that, in polar coordinates $z=a+re^{i\theta}$ for some $a\in{\mathbb{C}}$, for a smooth function $\phi$, one has
$d^{c}\phi={1\over 4\pi}\left(r{\partial\phi\over\partial r}d\theta-{1\over r}{\partial\phi\over\partial\theta}dr\right).$
The smooth 2-form $dd^{c}\sigma$ is supported on $U$ since $\log\sigma$ is harmonic elsewhere. On the annulus $A(0,{1\over 2},{3\over 2})$, in polar coordinates $z=re^{i\theta}$ one has (see the Appendix in [2])
$dd^{c}H(|z|)=d\left({r\over 4\pi}{\partial H(r)\over\partial r}d\theta\right)=d(h(r-c)d\theta)=O(1)dr\wedge d\theta=O(1)rdr\wedge d\theta$
since ${1\over 2}\leq r\leq{3\over 2}$ is bounded. Thus $dd^{c}H(|z|)$ is of finite mass on $A(0,{1\over 2},{3\over 2})\backslash{\mathcal{E}}$. On the annulus $A(a_{j},{1\over 2}r_{j},{3\over 2}r_{j})$,
$dd^{c}\left(r_{j}H\left({r_{j}\over|z-a_{j}|}\right)\right)$ $=d\left({r\over 4\pi}{\partial r_{j}H\left({r_{j}\over r}\right)\over\partial r}d\theta\right)$
$=d\left(-r_{j}h\left({r_{j}\over r}-c\right)d\theta\right)=O(r_{j})dr\wedge d\theta=O(1)rdr\wedge d\theta.$
Since $\sum_{j=1}^{\infty}r^{2}_{j}\leq\sum_{j=1}^{\infty}r_{j}<+\infty$, the support $U$ is of finite Lebesgue measure. We conclude that $dd^{c}\log\sigma$ is of finite mass.
Using the above computations with ${\mathcal{E}}:=f^{-1}(D)$, we recall the following important result due to Z. Chen, D.T. Huynh, R. Sun and S.Y. Xie [2].
Proposition 8 ([2]).
Let $f:{\mathbb{C}}\rightarrow{\mathbb{P}}^{n}({\mathbb{C}})$ be a holomorphic map. Let $D$ be a hypersurface in ${\mathbb{P}}^{n}({\mathbb{C}})$. Let ${\mathcal{Y}}_{f}:={\mathbb{C}}\backslash f^{-1}(D)$ with the smooth exhaustion function $\tau:=\log\sigma$ given above. Then
  1. (i)
    $0\leq T_{f}(r)-T_{\widetilde{f},\mathcal{Y}_{f}}(r)\leq\log r+O(1),$
    for $r\geq 1$, where $\widetilde{f}$ the restriction of $f$ on $\mathcal{Y}_{f}$.
  2. (ii)
    For $r>1$,
    $\mathfrak{X}^{+}_{\sigma}(r)\leq N^{(1)}_{f}(r,D)+C\log r,$
    where $C$ is a constant independent of $f$ and $D$.
We outline a proof here.
Proof.
Let ${\mathcal{E}}:=f^{-1}(D)$ and write :=${\mathcal{E}}:=\{a_{j}\}_{j=1}^{\infty}$. Denote by $B_{r}^{\sigma}$ and $S_{r}^{\sigma}$ the $\sigma$-ball of radius $r$ and its boundary. From (3.2), $B_{t}^{\sigma}\subset{\mathbb{D}}_{t}$ ($t\geq 1$) and
${\mathbb{D}}_{t}\backslash\left(\bigcup_{j=1}^{\infty}D(a_{j},{3\over 2}r_{j})\right)\subset B_{t}^{\sigma}~{}~{}~{}~{}~{}~{}~{}~{}(t\geq{3\over 2}).$
Thus
$0\leq T_{f}(r)-T_{\widetilde{f},\mathcal{Y}_{f}}(r)\leq\int_{{3\over 2}}^{r}{dt\over t}\int_{\left(\bigcup_{j=1}^{\infty}{\mathbb{D}}(a_{j},{3\over 2}r_{j})\right)\cap{\mathbb{D}}_{t}}f^{*}\omega_{FS}+O(1)$
$\leq\int_{1}^{r}{dt\over t}\int_{\bigcup_{j=1}^{\infty}{\mathbb{D}}(a_{j},{3\over 2}r_{j})}f^{*}\omega_{FS}+O(1)$
$=\int_{1}^{r}{dt\over t}\left(\sum_{j=1}^{\infty}\int_{{\mathbb{D}}(a_{j},{3\over 2}r_{j})}f^{*}\omega_{FS}\right)+O(1),~{}~{}~{}~{}(r\geq 1).$
Recall in (3.1) that the radius $r_{j}>0$ can be chosen arbitrarily small. For our purpose, for each $j\geq 1$, we choose $r_{j}>0$ sufficiently small so that $\int_{{\mathbb{D}}(a_{j},{3\over 2}r_{j})}f^{*}\omega_{FS}<2^{-j}$. Hence the above estimate yields
$0\leq T_{f}(r)-T_{\widetilde{f},\mathcal{Y}_{f}}(r)\leq\log r+O(1),~{}~{}~{}~{}(r\geq 1).$
This proves (i).
To prove (ii), using the above computations and following the argument in [16, Proposition 3.3, pp. 32-33] (see also [2, Line 3 on page 12]), the weighted Euler characteristic satisfies
$\mathfrak{X}^{+}_{\sigma}(r)=\mathfrak{X}_{\sigma}(r)=\int_{1}^{r}\#\{j:|a_{j}|<t\}{dt\over t}+O(\log r).$
Note that #$\#\{{\mathbb{D}}_{t}\cap{\mathcal{E}}\}$ is exactly $n_{f}^{(1)}(D,t)$, which is finite. Since $B^{\sigma}_{r}\subset{\mathbb{D}}_{r}$ for $r\geq 1$,
$\mathfrak{X}^{+}_{\sigma}(r)\leq N^{(1)}_{f}(r,D)+C\log r.$
This completes the proof of (ii). ∎
3.2. Nevanlinna theory on parabolic Riemann surfaces
Let ${\mathcal{Y}}$ be a parabolic open Riemann surface. The following is the starting point of the parabolic Nevanlinna theory.
Lemma 9 (Parabolic Logarithmic Derivative Lemma, Theorem 3.7 in [16] and Proposition 5.1 in [10]).
Let $f$ be a meromorphic function on ${\mathcal{Y}}$. For any $\delta>0$, one has
$m_{f^{\prime}/f,{\mathcal{Y}}}(r)\leq_{\rm exc}{1\over 2}((1+\delta)^{2}\log T_{f,{\mathcal{Y}}}(r)+\delta\log r+\mathfrak{X}^{+}_{\sigma}(r))+O(1).$
In particular, it gives the following
Lemma 10.
If $u$ is a nowhere vanishing holomorphic function on $\mathcal{Y}$ ((((i.e. $u$ is a unit)))), then, for any $\delta>0$, one has
$T_{u^{\prime}/u,{\mathcal{Y}}}(r)\leq_{\rm exc}{1\over 2}((1+\delta)^{2}\log T_{u,{\mathcal{Y}}}(r)+\delta\log r+\mathfrak{X}^{+}_{\sigma}(r))+O(1).$
We recall the following theorem which is an extension of H. Cartan’s theorem. It was from Theorem 6.3 in [10], although the statement is slightly different.
Theorem 11 (Theorem 6.3 in [10]).
Let $f=(f_{0},\dots,f_{n}):{\mathcal{Y}}\rightarrow{\mathbb{P}}^{n}({\mathbb{C}})$ be a holomorphic map whose image is not contained in any proper linear subspaces of ${\mathbb{P}}^{n}({\mathbb{C}})$, where $f_{0},\dots,f_{n}$ are holomorphic functions on ${\mathcal{Y}}$ without common zeros. Let $H_{1},\dots,H_{q}$ be hyperplanes on $\mathbb{P}^{n}(\mathbb{C})$. Then, for $\epsilon>0$,
$\int_{S(r)}\max_{J}\sum_{k\in J}$ $\lambda_{H_{k}}(f)d\mu_{r}+N_{W,\mathcal{Y}}(0,r)$
$\leq_{\rm exc}(n+1+\epsilon)T_{f,{\mathcal{Y}}}(r)+\left(\frac{n(n+1)}{2}+2\epsilon\right)(\mathfrak{X}^{+}_{\sigma}(r)+\delta\log r)+O(1),$
where $W$ is the Wronskian of $f$, and the maximum is taken over all subsets $J$ such that $H_{j},j\in J$, are in general position.
The following theorem can be derived from the proof of [18, Theorem 2.1] and Lemma 9.
Theorem 12.
Let $f_{0},\dots,f_{n}$ be entire functions on $\mathcal{Y}$ without common zeros. Assume that $f_{n+1}$ is the holomorphic function on ${\mathcal{Y}}$ such that $f_{0}+\dots+f_{n}+f_{n+1}=0$. If $\sum_{i\in I}f_{i}\neq 0$ for any proper subset $I\subset\{0,\dots,n+1\}$, then, for any $\epsilon>0$, we have
$(1-\epsilon)T_{f_{j}/f_{i},{\mathcal{Y}}}(r)\ \leq_{\operatorname{exc}}\sum_{i=0}^{n+1}N_{f_{i},{\mathcal{Y}}}^{(n)}(0,r)+\frac{n(n+1)}{2}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)$
for any pair $0\leq i,j\leq n$.
3.3. The moving GCD theorem for parabolic Riemann surfaces
One of the key components in [9] is the extension of the GCD theorem in [14] to the case where the coefficients of two coprime polynomials are allowed to be meromorphic functions of the same growth. In this subsection, we aim to extend this line of argument to the setting of parabolic Riemann surfaces. Before we proceed, let us briefly outline the development of the GCD theorem.
The foundational arguments in [14] are based on results initially established in [13] for the number field case. In [22], the GCD results concerning $\mathbb{G}_{m}^{n}$ from [13] were generalized to more general varieties by combining the approach in [21] with Diophantine approximation results for arbitrary varieties introduced by the first named author and Vojta [20]. More recently, a further generalization was achieved in [11], which applied Vojta’s results on integral points on semiabelian varieties for the number field case, as well as [15] for the complex case.
To proceed, we will first derive a version of the second main theorem with moving targets of the same growth, which will serve our purpose.
Let $\mathcal{Y}$ be a parabolic open Riemann surface. Let $a_{0},\ldots,a_{n}$ be meromorphic functions, not all zeros, on $\mathcal{Y}$, and let $L:=a_{0}X_{0}+\dots+a_{n}X_{n}$. Then $L$ defines a hyperplane $H$ in $\mathbb{P}^{n}$, over the field generated by $a_{0},\ldots,a_{n}$ over $\mathbb{C}$. We note that $H(z)$ is the hyperplane determined by the linear form $L(z)=a_{0}(z)X_{0}+\dots+a_{n}(z)X_{n}$ for $z\in\mathcal{Y}$ where $a_{0},\cdots,a_{n}$ are entire and there is no common zero of $a_{0},\cdots,a_{n}$ at $z$. The definition of the Weil function, proximity function and counting function can be easily extended to moving hyperplanes. For example,
(3.3) $\lambda_{H(z)}(P)=-\log\frac{|a_{0}(z)x_{0}+\cdots+a_{n}(z)x_{n}|}{\max\{|x_{0}|,\dots,|x_{n}|\}\max\{|a_{0}(z)|,\dots,|a_{n}(z)|\}},$
for $z\in\mathcal{Y}$ which is not a common zero of $a_{0},\cdots,a_{n}$, or a pole of any $a_{k}$, $0\leq k\leq n$. The First Main Theorem for a moving hyperplane $H$ can be stated as
(3.4) $T_{\mathbf{f},\mathcal{Y}}(r)=N_{\mathbf{f},\mathcal{Y}}(H,r)+m_{\mathbf{f},\mathcal{Y}}(H,r)+T_{\mathbf{a},\mathcal{Y}}(r)+O(1),$
where ${\mathbf{a}}:=[a_{0}:\cdots:a_{n}]$.
Without loss of generality, we will normalize the linear forms $L_{j}$, $1\leq j\leq q$, such that for each $1\leq j\leq q$, there exists $0\leq j^{\prime}\leq n$ with $a_{ij^{\prime}}=1$. Let $t$ be a positive integer and let $V(t)$ be the complex vector space spanned by the elements
$\left\{\prod a_{jk}^{n_{jk}}:n_{jk}\geq 0,\,\sum n_{jk}=t\right\},$
where the products and sums run over $1\leq j\leq q$ and $0\leq k\leq n$. Let $1=b_{1},b_{2},\cdots,b_{u}$ be a basis of $V(t)$ and $b_{1},\cdots,b_{w}$ a basis of $V(t+1)$. It’s clear that $u\leq w$. Moreover, we have
(3.5) $\liminf_{t\to\infty}\dim V(t+1)/\dim V(t)=1.$
Definition 13.
Let $V$ be a $\mathbb{C}$-vector space spanned by finitely many meromorphic functions. We say that a holomorphic map ${\mathbf{f}}=(f_{0},\dots,f_{n}):\mathcal{Y}\to\mathbb{P}^{n}({\mathbb{C}})$ is linearly non-degenerate over $V$ if whenever we have a linear combination $\sum_{i=1}^{m}a_{i}f_{i}=0$ with $a_{i}\in V$, then $a_{i}=0$ for each $i$. Otherwise we say that ${\mathbf{f}}$ is linearly degenerate over $V$.
The following formulation of the second main theorem with moving targets follows from the strategy in [23] and the proof of [17, Theorem A6.2.4], by adding the Wronskian term and computing the error term explicitly when applying the second main theorem (Theorem 11).
Theorem 14.
[…]
(3.6) $\int_{S(r)}\max_{J}\sum_{k\in J}$ $\lambda_{H_{k}(z)}({\mathbf{f}}(z))d\mu_{r}+\frac{1}{u}N_{W}(0,r)$
(3.7) $\leq_{\operatorname{exc}}\left(\frac{w}{u}(n+1)+\epsilon\right)T_{\mathbf{f},\mathcal{Y}}(r)+\frac{w}{u}(n+1)(t+2)\max_{1\leq j\leq q}T_{L_{j},{\mathcal{Y}}}(r)$
(3.8) $+{1\over 2}(n+1)^{2}{w^{2}\over u}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(1),$
where $T_{L_{j},{\mathcal{Y}}}(r)=T_{[a_{j0}:\cdots:a_{jn}],{\mathcal{Y}}}(r)$, $w=\dim V(t+1)$, $u=\dim V(t)$, and the maximum is taken over all subsets $J$ of $\{1,\dots,q\}$ such that $H_{j}(z)$, $j\in J$, are in general position.
Using the above theorem, and through the same argument from the proofs of [14, Theorem 5.7], [8, Theorem 26], and [9, Theorem 4.4], we can get the following GCD theorem (note we now consider the case that $u_{1},\ldots,u_{n}$ are units on $\mathcal{Y}$). We first fix some notations: Let $K$ be a subfield of the field of meromorphic functions on $\mathcal{Y}$. Let $F,G$ be coprime polynomials in $n$ variables of the same degree $d>0$ over $K$. Assume that one of the coefficients in each expansion of $F$ and $G$ is 1. For every positive integer $t$, we denote by $V_{F,G}(t)$ the (finite-dimensional) $\mathbb{C}$-vector space spanned by $\prod_{\alpha}\alpha^{n_{\alpha}}$, where $\alpha$ runs through all non-zero coefficients of $F$ and $G$, $n_{\alpha}\geq 0$ and $\sum n_{\alpha}=t$; we also put $d_{t}:=\dim V_{F,G}(t)$. For every integer $m\geq d$, we let $M=M_{m,n,d}=2\binom{m+n-d}{n}-\binom{m+n-2d}{n}$. We now define the characteristic function of a polynomial over $K$. Let $Q=\sum_{{\mathbf{i}}\in I_{Q}}a_{\mathbf{i}}\cdot{\mathbf{x}}^{\mathbf{i}}\in K[x_{1},\ldots,x_{n}]$, where ${\mathbf{i}}=(i_{1},\ldots,i_{n})$, ${\mathbf{x}}^{\mathbf{i}}=x_{1}^{i_{1}}\ldots x_{n}^{i_{n}}$, and $a_{\mathbf{i}}\neq 0$ if ${\mathbf{i}}\in I_{Q}$. We let
(3.9) $\|Q\|_{z}:=\max_{{\mathbf{i}}\in I_{Q}}\{|a_{\mathbf{i}}(z)|\},$
and define the characteristic function of $Q$ as
(3.10) $T_{Q,\mathcal{Y}}(r):=T_{[\cdots:a_{\mathbf{i}}:\cdots],\mathcal{Y}}(r),$
where $a_{\mathbf{i}}$ is taken for every ${\mathbf{i}}\in I_{Q}$.
Theorem 15 (GCD theorem on parabolic Riemann surfaces evaluating at units).
Let $n\geq 2$ be a positive integer. Let $u_{1},\ldots,u_{n}$ be units on $\mathcal{Y}$, i.e. holomorphic functions on $\mathcal{Y}$ without zeros and let ${\bf u}=(u_{1},\ldots,u_{n})$. Let $F,G$ be coprime polynomials in $n$ variables of the same degree $d>0$ over $K$. Assume that one of the coefficients in each expansion of $F$ and $G$ is 1. Then there exists a positive constant $c$ and a positive integer $b$ such that if the set $\{u_{1}^{i_{1}}\cdots u_{n}^{i_{n}}:i_{1}+\cdots+i_{n}\leq m\}$ is linearly independent over $V_{F,G}(Mb+1)$, then, for any $\delta>0$, we have
$N_{\rm gcd,{\mathcal{Y}}}(F({\bf u}),G({\bf u}),r)$ $\leq_{\rm exc}\left(\frac{M^{\prime}}{M}+\frac{w}{u}-1+\frac{\epsilon}{M}\right)mn\max_{1\leq i\leq n}T_{u_{i},{\mathcal{Y}}}(r)$
$+c^{\prime}(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r))+c^{\prime\prime}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(1),$
where $w=d_{Mb}$, $u=d_{M(b-1)}$, $c^{\prime}=\frac{w}{u}M(b+2)+1$, and $c^{\prime\prime}=M\frac{w^{2}}{u}$, and $M^{\prime}=\binom{m+n}{n}-M$.
We outline a proof here.
Proof.
As in the proof of [9, Theorem 4.3], we assume that the ideal $(F,G)$ is proper in $K[x]$. Let $(F,G)_{m}:=K[x]_{m}\cap(F,G)$, where we choose $\{\phi_{1},\dots,\phi_{M}\}$ to be a basis of the $K-$vector space $(F,G)_{m}$ consisting of elements of the form $F{\bf x}^{\bf i},G{\bf x}^{\bf i}$. Let $\Phi:=(\phi_{1},\dots,\phi_{M})$. For each $z\in{\mathcal{Y}}$, we can construct a basis $B_{z}$ of $V_{m}:=K[{\bf x}]_{m}/(F,G)_{m}$ with monomial representatives ${\bf x}^{{\bf i}_{1}},\dots,{\bf x}^{{\bf i}_{M^{\prime}}}$ as in the proof of Theorem 5.7 in [14]. Let $I_{z}:=\{{\bf i}_{1},\dots,{\bf i}_{M^{\prime}}\}$. For each ${\bf i}\not\in I_{z}$ with $|{\bf i}|=m$, we have
${\bf x}^{\bf i}+\sum_{{\bf j}\in I_{z}}c_{z,{\bf j}}{\bf x}^{\bf j}\in(F,G)_{m}$
for some choice $c_{z,{\bf j}}\in K$. Then, for each such ${\bf i}$, there is a linear form over $K$ such that
$\tilde{L}_{z,{\bf i}}(\phi_{1},\dots,\phi_{M})={\bf x}^{\bf i}+\sum_{{\bf j}\in I_{z}}c_{z,{\bf j}}{\bf x}^{\bf j}.$
We note that since there are correspondingly only finitely many choices of a monomial basis of $V_{m}$, there are only finitely many choices of $c_{z,{\bf j}}$, even as $z$ ranges over all of $\mathbb{C}$. Following the explicit construction in the proof of [9, Theorem 4.3], we can find a nonzero element $c_{z}\in V_{F,G}(M)$ such that the normalized linear form $L_{z,{\bf i}}:=c_{z}\tilde{L}_{z,{\bf i}}$ satisfies
$L_{z,{\bf i}}(\phi_{1},\dots,\phi_{M})=c_{z}\left({\bf x}^{\bf i}+\sum_{{\bf j}\in I_{z}}c_{z,{\bf j}}{\bf x}^{\bf j}\right)$
and
(3.11) $T_{L_{z,{\bf i}},{\mathcal{Y}}}(r)\leq M(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r)),$
by [9, (4.12)]. Moreover, the explicit construction also demonstrates that there are only a finitely many choice for $c_{z}$, even though $z$ runs through all of $\mathbb{C}$
Following the steps in the proof of Theorem 4.3 in [9], we then obtain
(3.12) $\log|L_{z,{\bf i}}(\Phi({\bf u}))|-\log\|L_{z,{\bf i}}\|_{z}\leq\log|{\bf u}(z)^{\bf i}|+\log\|F\|_{z}+\log\|G\|_{z}+O(1).$
Consequently, we have
(3.13) $\int_{S(r)}\sum_{|{\bf i}|=m,{\bf i}\not\in I_{z}}(\log|L_{z,{\bf i}}(\Phi({\bf u})|-\log\|L_{z,{\bf i}}\|)d\mu_{r}\leq M^{\prime}mnT_{{\bf u},\mathcal{Y}}(r)+M(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r))+O(1).$
Also, from (3.3) in the proof of [14, Theorem 5.7], we have
(3.14) $\int_{S(r)}\log\|\Phi({\bf u})\|d\mu_{r}=T_{\Phi({\bf u}),{\mathcal{Y}}}(r)+N_{\rm gcd,{\mathcal{Y}}}(F({\bf u}),G({\bf u}),r)+O(1).$
Since the set $\{u_{1}^{i_{1}}\cdots u_{n}^{i_{n}}:i_{1}+\cdots+i_{n}\leq m\}$ is linearly independent over $V_{F,G}(Mb+1)$, we have that $\Phi({\bf u}):{\mathbb{C}}\rightarrow{\mathbb{P}}^{M-1}$ is linearly nondegenerate over $V_{F,G}(Mb)=V(b)$. We note that, when computing Weil functions, we may assume that the coefficients of $L_{z,{\bf i}}$ are entire functions without common zeros. By this construction, for each $z\in{\mathcal{Y}}$, the hyperplanes defined by evaluating the linear forms $L_{z,{\bf i}}$ at $z$ with $|{\bf i}|=m$ and ${\bf i}\not\in I_{z}$ are in general position. Additionally, we observe that (3.11) and (3.13) remain valid under this choice of linear forms. Applying Theorem 14 to $\Phi({\bf u})$ and the set of the linear forms $\{L_{z,{\bf i}}:|{\bf i}|=m,{\bf i}\not\in I_{z}\}$, we have the following inequality:
(3.15) $\int_{S(r)}\sum_{|{\bf i}|=m,{\bf i}\not\in I_{z}}\lambda_{L_{z,{\bf i}}}(\Phi({\bf u}))d\mu_{r}+\frac{1}{u}N_{W}(0,r)$
(3.16) $\leq_{\operatorname{exc}}\left(\frac{w}{u}M+\epsilon\right)T_{\Phi({\bf u}),\mathcal{Y}}(r)+\frac{w}{u}M^{2}(b+2)(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r))$
(3.17) $+M^{2}{w^{2}\over u}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(1),$
where
$\lambda_{L_{z,{\bf i}}}(\Phi({\bf u}))=-\log|L_{z,{\bf i}}(\Phi({\bf u})|+\log\|\Phi({\bf u})\|+\log\|L_{z,{\bf i}}\|+O(1).$
By combining (3.15) with (3.13), (3.14), we get
$MN_{\rm gcd,{\mathcal{Y}}}(F({\bf u}),G({\bf u}),r)+\frac{1}{u}N_{W}(0,r)$
$\leq_{\operatorname{exc}}\left(\frac{w}{u}M-M+\epsilon\right)T_{\Phi({\bf u}),\mathcal{Y}}(r)+M^{\prime}mnT_{{\bf u},\mathcal{Y}}(r)$
$+\left(\frac{w}{u}M^{2}(b+2)+M\right)(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r))+M^{2}{w^{2}\over u}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(1),$
which proves the theorem. ∎
Remark.
Let $\epsilon$ be a sufficiently small positive number. By (3.5), we find $t$ such that $\frac{w}{u}\leq 1+\epsilon$. An explicit upper bound for $t$ can be found in the proof of [3, Proposition 5.3] in term of $\epsilon$ and the number of generator for $V(1)$, which is less than $qn$. More precisely, if $\epsilon<\frac{1}{16}$, then we can find
(3.18) $t\leq\frac{3qn}{\epsilon}\log\frac{1}{\epsilon}$
such that $\frac{w}{u}\leq 1+\epsilon$. Furthermore, for such $t$, we have
(3.19) $u:=\dim V(t)\leq\left(\frac{1}{\epsilon}\right)^{qn}.$
Consequently,
(3.20) $w:=\dim V(t+1)\leq(1+\epsilon)\left(\frac{1}{\epsilon}\right)^{qn}.$
Therefore, the constants appeared in (3.6) depend only on $\epsilon$ and the number of hyperplanes.
Theorem 16 (The key theorem).
Let $n\geq 2$ be a positive integer. Let $K$ be a subfield of the field of meromorphic functions on $\mathcal{Y}$. Let $F$ and $G$ be coprime polynomials in $K[x_{1},\ldots,x_{n}]$. Then, for any $\epsilon>0$, there exist a positive integer $m$ and positive reals $c_{i}$, $1\leq i\leq 4$, all depending only on $\epsilon$, $n$, and the degree of $F$ and $G$ such that for all $n$-tuple $(u_{1},\ldots,u_{n})$ of units on $\mathcal{Y}$ satisfying
(3.21) $\max_{1\leq i\leq n}T_{u_{i},\mathcal{Y}}(r)\geq c_{1}(\mathfrak{X}^{+}_{\sigma}(r)+\log r),$
we have either
(3.22) $T_{u_{1}^{m_{1}}\cdots u_{n}^{m_{n}},\mathcal{Y}}(r)\leq_{\rm exc}c_{2}(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r))+c_{3}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)$
for some non-trivial tuple of integers $(m_{1},\ldots,m_{n})$ with $|m_{1}|+\cdots+|m_{n}|\leq 2m$, or
(3.23) $N_{\rm gcd,\mathcal{Y}}(F(u_{1},\ldots,u_{n}),G(u_{1},\ldots,u_{n}),r)\leq_{\rm exc}\epsilon\max_{1\leq i\leq n}T_{u_{i},\mathcal{Y}}(r)+c_{4}(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r)).$
Furthermore, if $\epsilon>0$ is sufficiently small, then we may choose $m=C_{1}\epsilon^{-1}$, where $C_{1}$ is a positive real independent of $\epsilon$.
Proof.
We note that $c_{i},1\leq i\leq 4,$ will be determined during the course of the proof. As in the proof of [9, Theorem 4.1], we will assume that one of the coefficients in each expansion of $F$ and $G$ is 1. Recall that $M:=M_{m}:=2\binom{m+n-d}{n}-\binom{m+n-2d}{n}$ and $M^{\prime}:=M^{\prime}_{m}:=\binom{m+n}{n}-M=O(m^{n-2})$.
Let $\epsilon>0$. To establish (3.22) or (3.23), we can assume that $\epsilon$ is sufficiently small. Based on the above estimates, we can choose a real $C_{1}\geq 1$ independent of $\epsilon$ and the $u_{i},1\leq i\leq n,$ such that $m=C_{1}\epsilon^{-1}\geq 2d$, and
(3.24) $\frac{M^{\prime}mn}{M}\leq\frac{\epsilon}{4}.$
By (3.5) we may then choose a sufficiently large integer $b\in\mathbb{N}$ such that
(3.25) $\frac{w}{u}-1\leq\frac{\epsilon}{4mn},$
where $w:=\dim_{{\bf k}}V_{F,G}(Mb)$ and $u:=\dim_{{\bf k}}V_{F,G}(Mb-M)$.
Let ${\bf u}=(u_{1},\ldots,u_{n})$. Suppose that the set $\{u_{1}^{i_{1}}\cdots u_{n}^{i_{n}}:i_{1}+\cdots+i_{n}\leq m\}$ is linearly independent over $V_{F,G}(Mb+1)$. Then, by Theorem 15, we have
(3.26) $N_{\rm gcd,{\mathcal{Y}}}(F({\bf u}),G({\bf u}),r)$ $\leq_{\rm exc}\left(\frac{M^{\prime}}{M}+\frac{w}{u}-1+\frac{\epsilon}{M}\right)mn\max_{1\leq i\leq n}T_{u_{i},\mathcal{Y}}(r)$
(3.27) $+c^{\prime}(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r))+c^{\prime\prime}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(1).$
Let $c_{1}=4c^{\prime\prime}\epsilon^{-1}$. Then (3.21) reads
$c^{\prime\prime}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(1)\leq\frac{\epsilon}{3}\max_{1\leq i\leq n}T_{u_{i},\mathcal{Y}}(r).$
Together with (3.24) and (3.25), we derive from (3.26) and the above inequality that
$N_{\rm gcd,{\mathcal{Y}}}(F({\bf u}),G({\bf u}),r)\leq_{\rm exc}\epsilon\max_{1\leq i\leq n}T_{u_{i},\mathcal{Y}}(r)+c^{\prime}(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r)).$
Finally, if the set $\{u_{1}^{i_{1}}\cdots u_{n}^{i_{n}}:i_{1}+\cdots+i_{n}\leq m\}$ is linearly dependent over $V_{F,G}(Mb+1)$, i.e. there is a non-trivial relation
(3.28) $\sum_{\mathbf{i}}\alpha_{\mathbf{i}}\mathbf{u}^{\mathbf{i}}=0,$
where the sum runs over those $\mathbf{i}\in\mathbb{Z}_{\geq 0}^{n}$ and $|\mathbf{i}|\leq m$, and $\alpha_{\mathbf{i}}\in V_{F,G}(Mb+1)$ for each $\mathbf{i}$ such that $\alpha_{\mathbf{i}_{0}}\neq 0$ for some $\mathbf{i}_{0}$. Then we have
(3.29) $\sum_{\mathbf{i}\neq\mathbf{i}_{0}}\frac{\alpha_{\mathbf{i}}}{\alpha_{\mathbf{i}_{0}}}\mathbf{u}^{\mathbf{i}-\mathbf{i}_{0}}=-1.$
Since $\alpha_{\mathbf{i}}\in V_{F,G}(Mb+1)$, we have
(3.30) $T_{\frac{\alpha_{\mathbf{i}}}{\alpha_{\mathbf{i}_{0}}},\mathcal{Y}}(r)\leq T_{\alpha_{\mathbf{i}},\mathcal{Y}}(r)+T_{\alpha_{\mathbf{i_{0}}},\mathcal{Y}}(r)\leq 2(Mb+1)(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r)).$
By Theorem 12 and (3.30), the relation (3.29) implies
$T_{\frac{\alpha_{\mathbf{i}}}{\alpha_{\mathbf{i}_{0}}}\mathbf{u}^{\mathbf{i}-\mathbf{i}_{0}},\mathcal{Y}}(r)\leq_{\operatorname{exc}}C_{2}\cdot(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r))+c_{3}(\mathfrak{X}^{+}_{\sigma}(r)+\log r),$
where $C_{2}=4\binom{n+m}{n}(Mb+1)$ and $c_{3}={\binom{n+m}{n}}\left({\binom{n+m}{n}}+1\right)$. Then
(3.31) $T_{\mathbf{u}^{\mathbf{i}-\mathbf{i}_{0}},\mathcal{Y}}(r)$ $\leq T_{\frac{\alpha_{\mathbf{i}}}{\alpha_{\mathbf{i}_{0}}}\mathbf{u}^{\mathbf{i}-\mathbf{i}_{0}},\mathcal{Y}}(r)+T_{\frac{\alpha_{\mathbf{i}_{0}}}{\alpha_{\mathbf{i}}},\mathcal{Y}}(r)$
(3.32) $\leq_{\rm exc}c_{2}(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r))+c_{3}(\mathfrak{X}^{+}_{\sigma}(r)+\log r),$
where $c_{2}=C_{2}+2(Mb+1)$. ∎
For the convenience of later application, we state the following result for $n=1$, which can be easily obtained from the theory of resultant (see [19, Proposition 10]).
Proposition 17.
Let $K$ be a subfield of the field of meromorphic functions on $\mathcal{Y}$. Let $F,G\in K[x_{0},x_{1}]$ be nonconstant coprime homogeneous polynomials. Then there exists a positive constant $c$ such that for any holomorphic functions $g_{0}$ and $g_{1}$ on ${\mathcal{Y}}$ with no common zeros, the following holds:
$N_{\rm gcd,\mathcal{Y}}(F(g_{0},g_{1}),G(g_{0},g_{1}),r)\leq c(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r)).$
4. The abc theorem for parabolic Riemann surfaces
We use the above key theorem (Theorem 16) in the following manner: Let $\mathbf{u}=(u_{1},\ldots,u_{n})$, where $u_{i}$, $1\leq i\leq n$, are holomorphic functions on ${\mathcal{Y}}$ without zeros (i.e. $u_{i}$ are the units on ${\mathcal{Y}}$). Let $\mathbb{K}_{\mathbf{u}}$ be the subfield of the meromorphic function field such that, for any element $a\in\mathbb{K}_{\mathbf{u}}$, there exists a nonnegative real constant $c_{a}$ (which may depend on $a$) satisfying
$T_{a,{\mathcal{Y}}}(r)\leq_{\rm exc}c_{a}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(\max_{1\leq i\leq n}\{\log T_{u_{i},{\mathcal{Y}}}(r)\}).$
Notably, in our application, the constant $c_{a}$ can be effectively computed. We will use Theorem 16 for polynomials in the ring $\mathbb{K}_{\mathbf{u}}[x_{1},\ldots,x_{n}]$. Furthermore, we observe that $u_{i}^{\prime}/u_{i}\in\mathbb{K}_{\mathbf{u}}$ for $1\leq i\leq n$, since
(4.1) $T_{u_{i}^{\prime}/u_{i},{\mathcal{Y}}}(r)\leq_{\rm exc}{1\over 2}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+2\log T_{u_{i},{\mathcal{Y}}}(r)$
by Lemma 10.
Let $\mathbf{x}:=(x_{1},\ldots,x_{n})$. For $\mathbf{i}=(i_{1},\ldots,i_{n})\in\mathbb{Z}^{n}$, we let $\mathbf{x^{i}}:=x_{1}^{i_{1}}\cdots x_{n}^{i_{n}}$ and $\mathbf{u^{i}}:=u_{1}^{i_{1}}\cdots u_{n}^{i_{n}}$. For a non-constant polynomial $F(\mathbf{x})=\sum_{\mathbf{i}}a_{\mathbf{i}}\mathbf{x}^{\mathbf{i}}\in\mathbb{K}_{\mathbf{u}}[\mathbf{x}]:=\mathbb{K}_{\mathbf{u}}[x_{1},\dots,x_{n}]$, we define
(4.2) $D_{\mathbf{u}}(F)(\mathbf{x}):=\sum_{\mathbf{i}}\frac{(a_{\mathbf{i}}\mathbf{u}^{\mathbf{i}})^{\prime}}{\mathbf{u}^{\mathbf{i}}}\mathbf{x}^{\mathbf{i}}=\sum_{\mathbf{i}}\left(a_{\mathbf{i}}\cdot\sum_{j=1}^{n}i_{j}\frac{u_{j}^{\prime}}{u_{j}}\right)\mathbf{x}^{\mathbf{i}}\in\mathbb{K}_{\mathbf{u}}[\mathbf{x}].$
A direct computation shows that
(4.3) $F(\mathbf{u})^{\prime}=D_{\mathbf{u}}(F)(\mathbf{u}),$
and that the following product rule
(4.4) $D_{\mathbf{u}}(FG)=D_{\mathbf{u}}(F)G+FD_{\mathbf{u}}(G)$
holds for $F,G\in\mathbb{K}_{\mathbf{u}}[\mathbf{x}]$.
The following lemma shares the same proof of Lemma 5.1 in [9].
Lemma 18.
Let $\mathbf{u}=(u_{1},\ldots,u_{n})$, where $u_{i}$, $1\leq i\leq n$, are the units on ${\mathcal{Y}}$. Let $G$ be a non-constant polynomial in $\mathbb{K}_{\mathbf{u}}[\mathbf{x}]$ with no monomial factors and no repeated factors. Then $F:=D_{\mathbf{u}}(G)$ and $G$ are coprime in $\mathbb{K}_{\mathbf{u}}[\mathbf{x}]$ unless there exists a non-trivial tuple of integers $(m_{1},\ldots,m_{n})$ with $\sum_{i=1}^{n}|m_{i}|\leq 2\deg G$ such that $T_{u_{1}^{m_{1}}\cdots u_{n}^{m_{n}},{\mathcal{Y}}}\leq T_{G,{\mathcal{Y}}}(r).$
Proof.
The first step of the proof of [9, Lemma 5.1] shows that we may assume that $G$ is irreducible. Write $G=\sum_{\mathbf{i}}b_{\mathbf{i}}\mathbf{x}^{\mathbf{i}}\in\mathbb{K}_{\mathbf{u}}[\mathbf{x}]$, which contains at least two distinct terms, since $G$ has no monomial factors. Then $D_{\mathbf{u}}(G)/G\in\mathbb{C}^{*}$ since $D_{\mathbf{u}}(G)$ is not zero and $\deg D_{\mathbf{u}}(G)=\deg G$. Comparing the coefficients of $G$ and $D_{\mathbf{u}}(G)$, for nonzero $b_{\mathbf{i}}$ and $b_{\mathbf{j}}$ in $\mathbb{K}_{\mathbf{u}}$ ($\mathbf{i}\neq\mathbf{j}$), we have from (4.2) that
$\frac{(b_{\mathbf{i}}\mathbf{u}^{\mathbf{i}})^{\prime}}{b_{\mathbf{i}}\mathbf{u}^{\mathbf{i}}}=\frac{(b_{\mathbf{j}}\mathbf{u}^{\mathbf{j}})^{\prime}}{b_{\mathbf{j}}\mathbf{u}^{\mathbf{j}}},$
where $\mathbf{i}=(i_{1},\ldots,i_{n})$, and $\mathbf{j}=(j_{1},\ldots,j_{n})$. It implies that $\frac{b_{\mathbf{i}}\mathbf{u}^{\mathbf{i}}}{b_{\mathbf{j}}\mathbf{u}^{\mathbf{j}}}\in\mathbb{C}$. Therefore,
$T_{u_{1}^{i_{1}-j_{1}}\cdots u_{n}^{i_{n}-j_{n}},\mathcal{Y}}\leq T_{G,{\mathcal{Y}}}(r).$
As an application of the key theorem (Theorem 16) and using the above lemma, we obtain the following theorem.
Theorem 19.
Let $\mathbf{u}=(u_{1},\ldots,u_{n})$, where $u_{i}$, $1\leq i\leq n$, are units on ${\mathcal{Y}}$. Let $G$ be a non-constant polynomial in $\mathbb{K}_{\mathbf{u}}[\mathbf{x}]$ with no monomial factors and no repeated factors. Then, for any $\epsilon>0$, there exists a positive integer $m$ and a positive constants $\tilde{c}_{1}$ and $\tilde{c}_{2}$ depending only on $\epsilon$, $n$, and the degree and the coefficients of $G$, but independent of $\mathbf{u}$, such that the following holds: If
(4.5) $\max_{1\leq i\leq n}T_{u_{i},\mathcal{Y}}(r)\geq\tilde{c}_{1}(\mathfrak{X}^{+}_{\sigma}(r)+\log r),$
then we have either
(4.6) $T_{u_{1}^{m_{1}}\cdots u_{n}^{m_{n}},\mathcal{Y}}(r)\leq_{\rm exc}\tilde{c}_{2}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(\max_{1\leq j\leq n}\{\log T_{u_{j},{\mathcal{Y}}}(r)\})$
for a non-trivial $n$-tuple $(m_{1},\ldots,m_{n})$ of integers with $\sum_{i=0}^{n}|m_{i}|\leq 2m$, or
$N_{G(\mathbf{u}),{\mathcal{Y}}}(0,r)-N^{(1)}_{G(\mathbf{u}),{\mathcal{Y}}}(0,r)\leq_{\rm exc}\epsilon\max_{1\leq j\leq n}\{T_{u_{j},{\mathcal{Y}}}(r)\}.$
Proof.
We first note that the constant ${\tilde{c}}_{1},{\tilde{c}}_{2}$ will be determined during the course of the proof. Let $z_{0}\in\mathcal{Y}$. If $v_{z_{0}}(G(\mathbf{u}))\geq 2$, then it follows from (4.3) that $v_{z_{0}}(D_{\mathbf{u}}(G)(\mathbf{u}))=v_{z_{0}}(G(\mathbf{u}))-1$. Hence,
$\mathop{\mathrm{min}}\{v_{z_{0}}^{+}(G(\mathbf{u})),v_{z_{0}}^{+}(D_{\mathbf{u}}(G)(\mathbf{u}))\}\geq v_{z_{0}}^{+}(G(\mathbf{u}))-\mathop{\mathrm{min}}\{1,v_{z_{0}}^{+}(G(\mathbf{u}))\}.$
Consequently,
(4.7) $N_{\gcd,{\mathcal{Y}}}(G(\mathbf{u}),D_{\mathbf{u}}(G)(\mathbf{u}),r)\geq N_{G(\mathbf{u}),{\mathcal{Y}}}(0,r)-N^{(1)}_{G(\mathbf{u}),{\mathcal{Y}}}(0,r).$
Note that there exists a nonnegative real $c_{G}$ such that
(4.8) $T_{G,{\mathcal{Y}}}(r)\leq_{\rm exc}c_{G}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(\max_{1\leq i\leq n}\{\log T_{u_{i},{\mathcal{Y}}}(r)\}),$
since the coefficients of $G$ are in $\mathbb{K}_{\mathbf{u}}$. Furthermore,
(4.9) $T_{D_{\mathbf{u}}(G),{\mathcal{Y}}}(r)$ $\leq 2T_{G,{\mathcal{Y}}}(r)+\deg G\sum_{i=1}^{n}T_{u_{i}^{\prime}/u_{i},{\mathcal{Y}}}(r)$
(4.10) $\leq_{\rm exc}(2c_{G}+n\deg G)(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(\max_{1\leq i\leq n}\{\log T_{u_{i},{\mathcal{Y}}}(r)\}),$
and
(4.11) $T_{G,{\mathcal{Y}}}(r)+T_{D_{\mathbf{u}}(G),{\mathcal{Y}}}(r)\leq_{\rm exc}(3c_{G}+n\deg G)(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(\max_{1\leq i\leq n}\{\log T_{u_{i},{\mathcal{Y}}}(r)\})$
by (4.8) and (4.9). By Lemma 18 and (4.8), $G$ and $D_{\mathbf{u}}(G)$ are either coprime or
$T_{u_{1}^{m_{1}}\cdots u_{n}^{m_{n}},\mathcal{Y}}(r)\leq_{\rm exc}T_{G,{\mathcal{Y}}}(r)\leq_{\rm exc}c_{G}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(\max_{1\leq i\leq n}\{\log T_{u_{i},{\mathcal{Y}}}(r)\})$
for a non-trivial $n$-tuple $(m_{1},\ldots,m_{n})$ of integers with $\sum_{i=0}^{n}|m_{i}|\leq 2\deg G$, i.e., either $G$ and $D_{\mathbf{u}}(G)$ are coprime or (4.6) holds.
It remains to consider where $G$ and $D_{\mathbf{u}}(G)$ are coprime. Let $\epsilon>0$ be given, which may assume to be sufficiently small. We apply Theorem 16 and find $m$ and $c_{i}>0$, $1\leq i\leq 4$, such that if (4.5) holds with $\tilde{c}_{1}\geq c_{1}$ , then we have either
(4.12) $T_{u_{1}^{m_{1}}\cdots u_{n}^{m_{n}},\mathcal{Y}}(r)\leq_{\rm exc}c_{2}(T_{G,{\mathcal{Y}}}(r)+T_{D_{\mathbf{u}}(G),{\mathcal{Y}}}(r))+c_{3}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)$
for some non-trivial tuple of integers $(m_{1},\ldots,m_{n})$ with $|m_{1}|+\cdots+|m_{n}|\leq 2m$, or
(4.13) $N_{\rm gcd,\mathcal{Y}}(G({\mathbf{u}}),D_{\mathbf{u}}(G)({\mathbf{u}}),r)\leq_{\rm exc}\frac{\epsilon}{2}\max_{1\leq i\leq n}T_{u_{i},\mathcal{Y}}(r)+c_{4}(T_{G,{\mathcal{Y}}}(r)+T_{D_{\mathbf{u}}(G),{\mathcal{Y}}}(r)).$
We can conclude the proof by combining (4.11), (4.12), and (4.13), and letting
(4.14) $\tilde{c}_{1}=\max\{c_{1},2c_{4}(3c_{G}+n\deg G)\epsilon^{-1}\},~{}~{}~{}\mbox{and}~{}~{}\tilde{c}_{2}=\max\{c_{G},c_{3}+3c_{2}c_{G}+nc_{2}\deg G\}.$
The proof of [7, Theorem 4] can be adapted to the current situation, leading to the following result.
Theorem 20.
Let $G$ be a non-constant polynomial in $\mathbb{C}[x_{1},\ldots,x_{n}]$ with no monomial factors and no repeated factors. Then, for any $\epsilon>0$, there exist a positive real number $c_{0}$ and a proper Zariski closed subset $Z$ of $\mathbb{A}^{n}(\mathbb{C})$ such that for all $n$-tuple $\mathbf{u}=(u_{1},\ldots,u_{n})$ of units on ${\mathcal{Y}}$ with $\mathbf{u}(\mathcal{Y})\not\subset Z$ and
(4.15) $\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}}(r)\}\geq c_{0}(\mathfrak{X}^{+}_{\sigma}(r)+\log r),$
we have
(4.16) $N_{G(\mathbf{u}),\mathcal{Y}}(0,r)-N^{(1)}_{G(\mathbf{u}),\mathcal{Y}}(0,r))\leq_{\rm exc}\epsilon\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}}(r)\}$
if $G(\mathbf{u})\neq 0$. Here, $c_{0}$ can be effectively bounded from above in terms of $\epsilon$, $n$, and the degree of $G$. Moreover, the exceptional set $Z$ can be expressed as the zero locus of a finite set $\Sigma\subset\mathbb{C}[x_{1},\ldots,x_{n}]$ with the following properties: (Z1) $\Sigma$ depends on $\epsilon$ and $G$, but not $\mathcal{Y}$ and can be determined explicitly, (Z2) $|\Sigma|$ and the degree of each polynomial in $\Sigma$ can be effectively bounded from above in terms of $\epsilon$, $n$, and the degree of $G$.
We outline a proof here.
Proof.
We follow the proof of [7, Theorem 4]. Let $u_{i},1\leq i\leq n$, be units on ${\mathcal{Y}}$. We apply Theorem 19 through several reduction steps to derive a contradiction. Assume that (4.16) does not hold and let $c_{0}\geq\tilde{c}_{1}$, where $\tilde{c}_{1}$ is the constant from (4.5) in Theorem 19. Then, there exists a positive integer $M_{1}$, depending only on $\epsilon$, $n$, and the degree of $G$, but independent of $\mathbf{u}$, such that
(4.17) $T_{u_{1}^{m_{1}}\cdots u_{n}^{m_{n}},\mathcal{Y}}(r)\leq_{\rm exc}\tilde{c}_{2}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O\left(\max_{1\leq j\leq n}\{\log T_{u_{j},{\mathcal{Y}}}(r)\}\right)$
for a non-trivial $n$-tuple $(m_{1},\ldots,m_{n})$ of integers with $\sum|m_{i}|\leq M_{1}$. Thus, we have
(4.18) $\lambda_{1}:={u_{1}}^{m_{1}}\cdots{u_{n}}^{m_{n}}\in\mathbb{K}_{\mathbf{u}}.$
We may assume $\gcd(m_{1},\ldots,m_{n})=1$. By Lemma 14 in [7], $(m_{1},\ldots,m_{n})$ extends to a basis $(m_{1},\ldots,m_{n})$, $(a_{21},\ldots,a_{2n}),\ldots,(a_{n1},\ldots,a_{nn})$ of $\mathbb{Z}^{n}$ with
(4.19) $|a_{i1}|+\cdots+|a_{in}|\leq M_{1}+n\quad\text{ for }2\leq i\leq n.$
Consider the change of variables
(4.20) $\Lambda_{1}:=x_{1}^{m_{1}}\cdots x_{n}^{m_{n}},\quad\text{and }\quad X_{1,i}:=x_{1}^{a_{i1}}\cdots x_{n}^{a_{in}}\quad\text{for }2\leq i\leq n$
and put
(4.21) $\beta_{1,i}=u_{1}^{a_{i1}}\cdots u_{n}^{a_{in}}\quad\text{for }2\leq i\leq n.$
Let $A_{1}$ denote the $n\times n$ matrix whose rows are the above basis of $\mathbb{Z}^{n}$. Then we formally express the above identities as
(4.22) $(\Lambda_{1},X_{1,2},\ldots,X_{1,n})=(x_{1},\ldots,x_{n})^{A_{1}}\quad\text{and}\quad(\lambda_{1},\beta_{1,2},\ldots,\beta_{1,n})=(u_{1},\ldots,u_{n})^{A_{1}}.$
Let $B_{1}=A_{1}^{-1}$. The entries of $B_{1}$ can be bounded from above in terms of $M_{1}$ and $n$. We have
(4.23) $(x_{1},\ldots,x_{n})=(\Lambda_{1},X_{1,2},\ldots,X_{1,n})^{B_{1}}\quad\text{and}\quad(u_{1},\ldots,u_{n})=(\lambda_{1},\beta_{1,2},\ldots,\beta_{1,n})^{B_{1}}.$
Let $G_{1}(\Lambda_{1},X_{1,2},\ldots,X_{1,n})\in{\mathbb{C}}[\Lambda_{1},X_{1,2},\ldots,X_{1,n}]$ with no monomial factors and
(4.24) $G((\Lambda_{1},X_{1,2},\ldots,X_{1,n})^{B_{1}})=\Lambda_{1}^{d_{1}}X_{1,2}^{d_{2}}\cdots X_{1,n}^{d_{n}}G_{1}(\Lambda_{1},X_{1,2},\ldots,X_{1,n})$
for some integers $d_{i}$, $1\leq i\leq n$. Since the transformations in (4.22) and (4.23) are inverses of each other, and $G$ has no repeated irreducible factors, it follows that $G_{1}$ also has no repeated irreducible factors. The coefficients of $G_{1}$ are the same as the coefficients of $G$ and $\deg(G_{1})$ can be bounded from above explicitly in terms of $M_{1}$, $n$, and $\deg(G)$. Consider $G_{1}(\lambda_{1},X_{1,2},\ldots,X_{1,n})\in\mathbb{C}(\lambda_{1})[X_{1,2},\ldots,X_{1,n}]\subset{\mathbb{K}_{\mathbf{u}}}[X_{1,2},\ldots,X_{1,n}]$.
For the particular change of variables in (4.22), (4.23), and (4.24) (that depends on the matrix $A_{1}$), we apply the Lemma 16 in [7] to find a nonconstant polynomial $H_{1}^{\prime}\in\mathbb{C}[x_{1},\ldots,x_{n}]$ such that $G_{1}(\lambda_{1},X_{1,2},\ldots,X_{1,n})$ has neither monomial nor repeated irreducible factors if $H_{1}^{\prime}(u_{1},\ldots,u_{n})\not\equiv 0$. It is important to note that the polynomial $H_{1}^{\prime}$ is not only independent of $\mathbf{u}$, but also independent of $\mathcal{Y}$. We now take $H_{1}$ to be the product of all such $H_{1}^{\prime}$ where $A_{1}$ ranges over the finitely many elements of $\operatorname{GL}_{n}(\mathbb{Z})$ with $\|A_{1}\|_{\infty}\leq M_{1}+n$. From Lemma 16 in [7]  we know that $\deg H_{1}$ depends only on $\epsilon$, $n$ and $\deg G$.
Since the $u_{i}$’s, $\lambda_{1}$, and $\beta_{1,j}$’s are the units on $\mathcal{Y}$, we have
(4.25) $\begin{split}N_{G(\mathbf{u}),\mathcal{Y}}(0,r)-N_{G(\mathbf{u}),\mathcal{Y}}^{(1)}(0,r)=N_{G_{1}(\lambda_{1},\beta_{1,2},\ldots,\beta_{1,n}),\mathcal{Y}}(0,r)-N_{G_{1}(\lambda_{1},\beta_{1,2},\ldots,\beta_{1,n}),\mathcal{Y}}^{(1)}(0,r).\end{split}$
On the other hand, by taking $\tilde{c}_{1}\geq 2\tilde{c}_{2}\epsilon^{-1}$, we deduce from (4.15) and (4.17) that
(4.26) $T_{{u_{1}}^{m_{1}}\cdots{u_{n}}^{m_{n}},\mathcal{Y}}(r)\ \leq_{\rm exc}\epsilon\max_{1\leq j\leq n}\{T_{u_{j},{\mathcal{Y}}}(r)\}.$
Then, by (4.22), (4.23) and (4.26), we have:
$\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}}(r)\}={\rm O}\left(\max\{T_{\lambda_{1},\mathcal{Y}}(r),T_{\beta_{1,2},\mathcal{Y}}(r),\ldots,T_{\beta_{1,n},\mathcal{Y}}(r)\}\right)={\rm O}\left(\max_{2\leq i\leq n}\{T_{\beta_{1,i},\mathcal{Y}}(r)\}\right).$
More precisely, we find positive real numbers $b_{1,1}$ and $b_{1,2}$ such that
(4.27) $b_{1,1}\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}}(r)\}\leq\max_{2\leq i\leq n}\{T_{\beta_{1,i},\mathcal{Y}}(r)\},\quad\text{and }\quad\max_{2\leq i\leq n}\{T_{\beta_{1,i},\mathcal{Y}}(r)\}\leq b_{1,2}\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}}(r)\}.$
In conclusion, at the end of this step we have that
$\max_{2\leq i\leq n}\{T_{\beta_{1,i},\mathcal{Y}}(r)\}\geq b_{1,1}c_{0}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)$
holds, and the inequality
(4.28) $N_{G_{1}(\lambda_{1},\beta_{1,2},\ldots,\beta_{1,n})}(0,r)-N_{G_{1}(\lambda_{1},\beta_{1,2},\ldots,\beta_{1,n})}^{(1)}(0,r)$ $\leq_{\rm exc}\frac{\epsilon}{b_{1,2}}\max_{2\leq i\leq n}\{T_{\beta_{1,i},\mathcal{Y}}(r)\}$
fails to hold under the assumption that $H_{1}(u_{1},\ldots,u_{n})\not\equiv 0$.
The above is the first step. There are $(n-1)$ many steps in total as in the proof of [7, Theorem 4]. Let $12\leq s\leq n-$ 1 and suppose that we have completed the step $s-1$. This includes the construction of $H_{s-1}\in{\mathbb{C}}[x_{1},\ldots,x_{n}]$ with degree depends on $\epsilon$, $n$ and $\deg G$ only. The last one is Step $(n-1)$ resulting in $H_{n-1}\in{\mathbb{C}}[x_{1},\ldots,x_{n}]$. We now define $H=H_{1}\cdots H_{n-1}$. Then $\deg H$ depends only on $\epsilon$, $n$ and $\deg G$, since each $H_{i}$ does so. Moreover, the polynomial $H$ is not only independent of $\mathbf{u}$, but also independent of $\mathcal{Y}$.
Suppose $H(u_{1},\ldots,u_{n})\not\equiv 0$. Assume we go through all the $(n-1)$ steps to get the polynomial
$P(X_{n-1,n}):=G_{n-1}(\lambda_{1},\ldots,\lambda_{n-1},X_{n-1,n})\in{\mathbb{K}_{\mathbf{u}}}[X_{n-1,n}]$
such that its degree can be bounded explicitly in terms of $M_{n-1},\ldots,M_{1}$, $n$, and $\deg(G)$. At the end of the step $(n-1)$, we have that $\beta:=\beta_{n-1,n}$, which is a (non-constant) unit on ${\mathcal{Y}}$, such that
(4.29) $T_{\beta,\mathcal{Y}}(r)\geq b_{n-1,1}c_{0}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)$
holds and the inequality
(4.30) $N_{P(\beta)}(0,r)-N^{(1)}_{P(\beta)}(0,r)\leq_{\rm exc}\frac{\epsilon}{b_{n-1,2}}T_{\beta,\mathcal{Y}}(r)$
fails to hold. Here, $b_{n-1,1}$ and $b_{n-1,2}$ are computable positive real numbers. On the other hand, since $H_{n-1}(u_{1},\ldots,u_{n})\not\equiv 0$, the polynomial $P(X_{n-1,n})$ has neither monomial nor repeated irreducible factors, and thus, similar to (4.7), we have
(4.31) $N_{\gcd,{\mathcal{Y}}}(P(\beta),D_{\beta}(P)(\beta),r)\geq N_{P(\beta)}(0,r)-N^{(1)}_{P(\beta)}(0,r).$
By Proposition 17, there is a positive constant $C_{1}$ independent of $\beta$ such that
(4.32) $N_{\gcd,{\mathcal{Y}}}(P(\beta),D_{\beta}(P)(\beta),r)\leq C_{1}(T_{P,{\mathcal{Y}}}(r)+T_{D_{\beta}(P),{\mathcal{Y}}}(r)).$
Similar to the arguments for (4.9), we find a positive constant $C_{2}$ independent of $\beta$ such that
(4.33) $T_{P,{\mathcal{Y}}}(r)+T_{D_{\beta}(P),{\mathcal{Y}}}(r)\leq_{\rm exc}C_{2}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(\log T_{\beta,\mathcal{Y}}(r)).$
Then by (4.29), (4.31), (4.32), and (4.33), we have
(4.34) $N_{P(\beta)}(0,r)-N^{(1)}_{P(\beta)}(0,r)\leq_{\rm exc}\frac{C_{1}C_{2}}{b_{n-1,1}c_{0}}T_{\beta,\mathcal{Y}}(r)+O(\log T_{\beta,\mathcal{Y}}(r)),$
which contradicts (4.30) by taking $c_{0}\geq 2C_{1}C_{2}b_{n-1,2}b_{n-1,1}^{-1}\epsilon^{-1}$. ∎
Theorem 21.
Let $G=\sum_{\mathbf{i}\in I_{G}}\alpha_{\mathbf{i}}{\mathbf{x}}^{\mathbf{i}}\in$ be a non-constant polynomial in $\mathbb{C}[x_{1},\cdots,x_{n}]$. Assume that $G(0,\ldots,0)\neq 0$ and $\deg_{x_{i}}G=\deg G=d$ for $1\leq i\leq n$. Let $Z_{1}$ be the Zariski closed subset that is the union of hypersurfaces of $\mathbb{A}^{n}$ of the form $\sum_{\mathbf{i}\in J}\alpha_{\mathbf{i}}{\mathbf{x}}^{\mathbf{i}}=0$, where $J$ is a non-empty subset of $I_{G}$. Let $u_{1},\ldots,u_{n}$ be the units on $\mathcal{Y}$ such that the image of ${\bf u}:=(u_{1},\ldots,u_{n}):\mathcal{Y}\to(\mathbb{C}^{*})^{n}$ is not contained in $Z_{1}$. Then
(4.35) $N_{G({\bf u}),\mathcal{Y}}(0,r)\geq_{\operatorname{exc}}(\deg G-\epsilon)\cdot\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}}(r)\}-c(\mathfrak{X}^{+}_{\sigma}(r)+\log r),$
where $c=\frac{1}{2}\binom{n+d}{n}(\binom{n+d}{n}+1)$.
Proof.
Since $G(0,\ldots,0)\neq 0$ and $\deg_{x_{i}}G=\deg G=d$ for $1\leq i\leq n$, we may assume without loss of generality that $G(0,\ldots,0)=1$, and write
(4.36) $G(\mathbf{u})=1+\sum_{1\leq i\leq n}\alpha_{\mathbf{i}_{i}}u_{i}^{d}+\sum_{\mathbf{i}\in I_{G}\setminus I}\alpha_{\mathbf{i}}{\mathbf{u}}^{\mathbf{i}}$
where $\alpha_{\mathbf{i}_{i}}\neq 0$ for $0\leq i\leq n$ and $I=\{\mathbf{i}_{0}:=(0,0,\ldots,0),\mathbf{i}_{1}:=(d,0,\ldots,0),\ldots,\mathbf{i}_{n}:=(0,\ldots,0,d)\}$. If ${\mathbf{u}}({\mathbb{C}})$ is not contained in $Z_{1}$, we may use Theorem 12 to show that
$(d-\epsilon)\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}}(r)\}\leq N_{G(\mathbf{u}),\mathcal{Y}}(0,r)+c(\mathfrak{X}^{+}_{\sigma}(r)+\log r),$
where $c=\frac{1}{2}\binom{n+d}{n}(\binom{n+d}{n}+1)$. ∎
Theorem 22 (The abc Theorem on parabolic Riemann surfaces).
Let $G$ be a non-constant polynomial in $\mathbb{C}[x_{1},\ldots,x_{n}]$ with no monomial factors and no repeated factors. Then, for any $\epsilon>0$, there exist a positive real number $c_{0}$ and a proper Zariski closed subset $W$ of $\mathbb{A}^{n}(\mathbb{C})$ such that for all holomorphic maps $\mathbf{u}=(u_{1},\ldots,u_{n}):\mathcal{Y}\to(\mathbb{C}^{*})^{n}$ with $\mathbf{u}(\mathcal{Y})\not\subset W$ and satisfying
(4.37) $\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}}(r)\}\geq c_{0}(\mathfrak{X}^{+}_{\sigma}(r)+\log r),$
we have the following two statements hold
  1. (a)
    $N_{G(\mathbf{u}),\mathcal{Y}}(0,r)-N^{(1)}_{G(\mathbf{u}),\mathcal{Y}}(0,r))\leq_{\rm exc}\epsilon\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}}(r)\}$.
  2. (b)
    If $G(0,\ldots,0)\neq 0$ and $\deg_{x_{i}}G=\deg G=d$ for $1\leq i\leq n$, then
    $N^{(1)}_{G(\mathbf{u}),\mathcal{Y}}(0,r))\geq_{\rm exc}(\deg G-\epsilon)\cdot\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}}(r)\}.$
Here, $c_{0}$ can be effectively bounded from above in terms of $\epsilon$, $n$, and the degree of $G$. Moreover, the exceptional set $W$ can be expressed as the zero locus of a finite set $\Sigma\subset K[x_{1},\ldots,x_{n}]$ with the following properties: (Z1) $\Sigma$ depends on $\epsilon$ and $G$ (not on $\mathcal{Y}$) and can be determined explicitly, (Z2) $|\Sigma|$ and the degree of each polynomial in $\Sigma$ can be effectively bounded from above in terms of $\epsilon$, $n$, and the degree of $G$.
Proof.
The proof follows from Theorem 20 and Theorem 6.4. Let $W=Z\cup Z_{1}\cup[G=0]$, where $Z$ and $Z_{1}$ are the exceptional set from Theorem 20 and Theorem 6.4 for $\frac{\epsilon}{2}$ respectively. Let $\mathbf{u}=(u_{1},\ldots,u_{n}):\mathcal{Y}\to(\mathbb{C}^{*})^{n}$ with $\mathbf{u}(\mathcal{Y})\not\subset W$. Theorem 20 implies that
$N_{G(\mathbf{u}),\mathcal{Y}}(0,r)-N^{(1)}_{G(\mathbf{u}),\mathcal{Y}}(0,r))\leq\frac{\epsilon}{2}\max_{1\leq i\leq n}\{T_{u_{i}}(r)\}.$
Together with (4.35) in Theorem 6.4, we derive
(4.38) $N^{(1)}_{G(\mathbf{u}),\mathcal{Y}}(0,r))\geq_{\operatorname{exc}}\left(\deg G-\frac{\epsilon}{2}\right)\max_{1\leq i\leq n}\{T_{u_{i}}(r)\}-c(\mathfrak{X}^{+}_{\sigma}(r)+\log r).$
By replacing the constant $c_{0}$ in (4.37) by $c_{1}=\max\{c_{0},2c\epsilon^{-1}\}$, we may derive from (4.38) that
$N^{(1)}_{G(\mathbf{u}),\mathcal{Y}}(0,r))\geq_{\operatorname{exc}}(\deg G-\epsilon)\cdot\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}}(r)\}.$
5. Proof of the main theorem
Proof of the Main Theorem.
Let $f=(f_{0},\ldots,f_{n}):\mathbb{C}\mapsto\mathbb{P}^{n}(\mathbb{C})$, where $f_{0},\ldots,f_{n}$ are entire functions without common zero. We may assume $f_{i}$ is not identically zero by letting $Z$ contain all coordinate hyperplanes. When $f$ is rational, the assertion follows directly from Theorem 6, as in this case we have
$N_{G(f)}(0,r)=N(G(f))\log r,\quad N^{(1)}_{G(f)}(0,r)=N^{(1)}(G(f))\log r,\quad\text{ and }T_{f}(r)=(\deg f)\log r$
for $r$ sufficiently large. Therefore, it remains to consider the situation when $f$ is not rational. In this case, we have
$\log r=o(T_{f}(r)).$
Let $D:=\sum_{i=0}^{n}H_{i}$, where $H_{i}:=[x_{i}=0]$, $0\leq i\leq n$, are the coordinate hyperplanes. We consider ${\mathcal{Y}_{f}}:={\mathbb{C}}\backslash f^{-1}(D)$. By (ii) of Proposition 8,
(5.1) $\mathfrak{X}_{\sigma}^{+}(r)\leq\sum_{i=0}^{n}N_{f}^{(1)}(H_{i},r)+O(\log r).$
Since ${\rm mult}_{z_{0}}(f^{*}H_{i})\geq\ell$ for all $i$ with $0\leq i\leq n$ and for every $z_{0}\in{\mathbb{C}}$, we have $N_{f}^{(1)}(H_{i},r)\leq{1\over\ell}T_{f}(r)$ for $i=0,1,\dots,n$. Thus
(5.2) $\mathfrak{X}_{\sigma}^{+}(r)\leq\frac{n+1}{\ell}T_{f}(r)+O(\log r),$
where $\ell$ is a positive integer to be determined. Let $\tilde{f}_{i}$ be the restriction of $f_{i}$ to $\mathcal{Y}_{f}$ and let $u_{i}:=\frac{\tilde{f}_{i}}{\tilde{f}_{0}}$. Then $u_{i},1\leq i\leq n$, are units on $\mathcal{Y}_{f}$. Denote by ${\tilde{u}}=(1,u_{1},\ldots,u_{n})$ and $\mathbf{u}=(u_{1},\ldots,u_{n})$. Then $\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}_{f}}(r)\}=T_{\tilde{u},\mathcal{Y}_{f}}(r)+O(1)=T_{{\tilde{f}},\mathcal{Y}_{f}}(r)+O(1),$ where $\tilde{f}$ is the restriction of $f$ to $\mathcal{Y}_{f}$.
We verify that
(5.3) $\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}_{f}}(r)\}\geq c_{0}(\mathfrak{X}_{\sigma}^{+}(r)+\log r)$
holds, where $c_{0}$ is a sufficient large constant determined in Theorem 22. Indeed, if it is not true, then there exists a constant $c>0$ and a sequence $r_{n}\rightarrow+\infty$ such that
$\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}_{f}}(r_{n})\}\leq c(\mathfrak{X}_{\sigma}^{+}(r_{n})+\log r_{n}).$
Using (ii) of Proposition 8, (5.2) and the fact that $\log r=o(T_{f}(r))$, we have that
$T_{\tilde{u},\mathcal{Y}_{f}}(r_{n})\leq{c(n+2)\over\ell}T_{\tilde{u},\mathcal{Y}_{f}}(r_{n})+o(T_{\tilde{u},\mathcal{Y}_{f}}(r_{n})),$
which gives a contradiction when $\ell>c(n+2)$. Hence (5.3) holds. Therefore, by Theorem 22, there exist a proper Zariski closed subset $W$ of $\mathbb{A}^{n}$ and a positive constant $c$, which are independent of $f$, such that the following two statements hold
(5.4) $N_{\tilde{G}({\mathbf{u}}),\mathcal{Y}_{f}}(0,r)-N^{(1)}_{\tilde{G}({\mathbf{u}}),\mathcal{Y}_{f}}(0,r))\leq_{\operatorname{exc}}\frac{\epsilon}{3}\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}_{f}}(r)\}.$
(5.5) $N^{(1)}_{\tilde{G}({\mathbf{u}}),\mathcal{Y}_{f}}(0,r))\geq_{\operatorname{exc}}\left(\deg G-\frac{\epsilon}{2}\right)\cdot\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}_{f}}(r)\}.$
Here, $\tilde{G}\in\mathbb{C}[x_{1},\ldots,x_{n}]$ such that $\tilde{G}(x_{1},\ldots,x_{n})=G(1,x_{1},\ldots,x_{n})$. Therefore, $\tilde{G}({\mathbf{u}})=G(\tilde{u})$. Since $\widetilde{f}$ is the restriction of $f$ to $\mathcal{Y}_{f}=\mathbb{C}\setminus{f^{-1}(D)}$ and $\tilde{f}_{i}$ has no zero on $\mathcal{Y}_{f}$, it is obvious that
(5.6) $N^{(1)}_{\tilde{G}({\mathbf{u}}),\mathcal{Y}_{f}}(0,r)=N^{(1)}_{G(\widetilde{f}),\mathcal{Y}_{f}}(0,r)\leq N^{(1)}_{G(f)}(0,r).$
Then it follows immediately from (5.5) that
$N^{(1)}_{G(f)}(0,r)\geq_{exc}\left(\deg G-\frac{\epsilon}{2}\right)\cdot\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}_{f}}(r)\}\geq(\deg G-\epsilon)T_{f}(r)$
by (ii) of Proposition 8 and the fact that $\log r=o(T_{f}(r))$. This proves (ii).
To show (i), we apply (5.6) again and obtain the following:
$N_{G(f)}(0,r)-N^{(1)}_{G(f)}(0,r)$ $\leq N_{G(f)}(0,r)-N^{(1)}_{\tilde{G}({\mathbf{u}}),\mathcal{Y}_{f}}(0,r))$
$=N_{G(f)}(0,r)-N_{\tilde{G}(\mathbf{u}),\mathcal{Y}_{f}}(0,r)+N_{\tilde{G}({\mathbf{u}}),\mathcal{Y}_{f}}(0,r)-N^{(1)}_{\tilde{G}(\mathbf{u}),\mathcal{Y}_{f}}(0,r))$
$\leq_{\operatorname{exc}}\deg G\cdot T_{f}(r)-\left(\deg G-\frac{5\epsilon}{6}\right)\cdot\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}_{f}}(r)\},$
by the First Main Theorem, Theorem 6.4, (5.4) and (5.5). Again as $\log r=o(T_{f}(r))$, it yieids
(5.7) $N_{G(f)}(0,r)-N^{(1)}_{G(f)}(0,r)\leq\epsilon T_{f}(r)$
by (i) of Proposition 8. Finally, we can conclude the proof by taking $Z$ to be the Zariski closure of $W$ in $\mathbb{P}^{n}({\mathbb{C}})$. ∎
6. Toric Varieties
Proof of Theorem 3.
We recall the following setup of finding a natural finite open covering of $X$ from the proof of [13, Theorem 4.4]. Let $\Sigma$ be the fan corresponding to the smooth projective toric variety $X$. Then there is a finite affine covering $\{X_{\eta}\}$ of $X$, where $\eta\in\Sigma$ is a n-dimensional smooth cone with an isomorphism $i_{\eta}:X_{\eta}\to\mathbb{A}^{n}$. This isomorphism restricts to an automorphism of $\mathbb{G}_{m}^{n}$, where we identify $\mathbb{G}_{m}^{n}\subset X_{\eta}$ naturally as a subset of $X$ and $\mathbb{G}_{m}^{n}\subset\mathbb{A}^{n}$ in the standard way such that $\mathbb{A}^{n}\setminus\mathbb{G}_{m}^{n}$ consists of the affine coordinate hyperplanes $\{x_{i}=0\}$, $1\leq i\leq n$. Under this identification, an entire orbifold curve $f$ into $(X,\Delta)$ corresponds to a collection of $n$-tuples of meromorphic functions ${i}_{\eta}(f)=(u_{\eta,1},\ldots,u_{\eta,n})$ such that the zero multiplicity of each $u_{\eta,i}$, $1\leq i\leq n$, is at least $\ell$ if not zero. If $z_{0}\in\mathbb{C}$ is a pole of some $u_{\eta,i}$, $1\leq i\leq n$, then $f(z_{0})$ is in a component of $D_{0}$, the boundary of $X$, with multiplicity
$-\mathop{\mathrm{min}}\{0,v_{z_{0}}(u_{\eta,1}),\ldots,v_{z_{0}}(u_{\eta,n})\}\geq\ell.$
Let $\widetilde{{i}_{\eta}}:X_{\eta}\to\mathbb{P}^{n}({\mathbb{C}})$ be the composition of $i_{\eta}$ with the natural embedding of $\mathbb{A}^{n}$ to $\mathbb{P}^{n}({\mathbb{C}})$. Then $\widetilde{{i}_{\eta}}(f)=(f_{\eta,0},\ldots,f_{\eta,n})=[1:u_{\eta,1}:\cdots:u_{\eta,n}]:\mathbb{C}\to\mathbb{P}^{n}({\mathbb{C}})$, where $f_{\eta,0},\ldots,f_{\eta,n}$ are entire functions with no common zero. Furthermore, the pullback $(\widetilde{{i}_{\eta}}^{-1})^{*}(D|_{X_{\eta}})$ of $D$ to $\mathbb{P}^{n}({\mathbb{C}})$ is defined by some nonconstant homogeneous polynomial $F_{\eta}\in\mathbb{C}[x_{0},\ldots,x_{n}]$ with no monomial and repeated factors and $\deg_{X_{i}}F_{\eta}=\deg F_{\eta}$ for each $0\leq i\leq n$ since $D$ is reduced and is in general position with the boundary of $\mathbb{G}_{m}^{n}$ in $X$. By [25, Proposition 12.11], there exist a non-zero constant $c_{\eta,A}$ and a proper closed subset $Z_{\eta,A}\subset X_{\eta}$, depending on $\eta$ and $A$ such that
(6.1) $T_{\widetilde{{i}_{\eta}}(f)}(r)\leq c_{\eta,A}T_{A,f}(r)+O(1)$
if the image of ${i}_{\eta}(f)$ is not contained in $Z_{\eta,A}$. Let $M$ be the number of $X_{\eta}$ in $\Sigma$. We may now apply the Main Theorem to each $X_{\eta}$. Let $\epsilon>0$. Then there exists a proper Zariski closed subset $W_{\eta}\subset X_{\eta}$ for each $\eta\in\Sigma$ such that for sufficiently large integer $\ell$, we have
(6.2) $N_{F_{\eta}(\widetilde{{i}_{\eta}}(f))}(0,r)-N^{(1)}_{F_{\eta}(\widetilde{{i}_{\eta}}(f))}(0,r)\leq_{\rm exc}\frac{\epsilon}{Mc_{\eta,A}}T_{\widetilde{{i}_{\eta}}(f)}(r),$
and
(6.3) $N^{(1)}_{F_{\eta}(\widetilde{{i}_{\eta}}(f))}(0,r)\geq_{\rm exc}(\deg F_{\eta}-\frac{\epsilon}{Mc_{\eta,A}})T_{\widetilde{{i}_{\eta}}(f)}(r),$
for all $f$ such that the image of $f$ restricting to $X_{\eta}$ is not contained in $W_{\eta}$. Since $\deg_{X_{i}}F_{\eta}=\deg F_{\eta}$ for each $0\leq i\leq n$, it implies that $\deg F_{\eta}\cdot T_{\widetilde{{i}_{\eta}}(f)}(r)=T_{F_{\eta}(\widetilde{{i}_{\eta}}(f))}(r)+O(1).$ Then by the first main theorem and (6.3), we have
(6.4) $m_{F_{\eta}(\widetilde{{i}_{\eta}}(f))}(0,r)\leq_{\rm exc}\frac{\epsilon}{Mc_{\eta,A}}T_{\widetilde{{i}_{\eta}}(f)}(r).$
Let $W$ be the union of $W_{\eta}$ and $Z_{\eta,A}$ for $\eta\in\Sigma$. Then
(6.5) $N_{f}(D,r)-N_{f}^{(1)}(D,r)\leq\sum_{\eta\in\Sigma}N_{F_{\eta}(\widetilde{{i}_{\eta}}(f))}(0,r)-N^{(1)}_{F_{\eta}(\widetilde{{i}_{\eta}}(f))}(0,r)\leq_{\rm exc}\epsilon T_{A,f}(r),$
and
(6.6) $m_{f}(D,r)\leq\sum_{\eta\in\Sigma}m_{F_{\eta}(\widetilde{{i}_{\eta}}(f))}(0,r)\leq_{\rm exc}\epsilon T_{A,f}(r)$
for all orbifold curve $f:\mathbb{C}\to(X,\Delta)$ with multiplicity at least $\ell$ along $D_{0}$ and $f(\mathbb{C})\not\subset W$. Finally, the assertion (ii) follows from (6.5) and the following inequality:
$N_{f}(D,r)=T_{D,f}(r)-m_{f}(D,r)\geq_{\rm exc}T_{D,f}(r)-\epsilon T_{A,f}(r)$
by (6.6). ∎
Proof of Theorem 4.
Let $\pi:Y\to X$ be a finite morphism. We recall from [4, Lemma 1] and the proof of [6, Theorem 6] that $R$ is a big divisor since $(Y,\Delta)$ is of general type. Let $R_{0}$ be an irreducible component of $R$. Without loss of generality, we let $R_{0}=R$. Otherwise, we simply repeat the following steps for each irreducible component. Let $D=\pi(R)$, which is in general position with $D_{0}$ by assumption. Now, consider $f:\mathbb{C}\to(Y,\Delta)$ as a nonconstant orbifold entire curve. Since $\pi^{*}D$ has multiplicity at least 2 along $R$, we have
(6.7) $N_{f}(R,r)\leq N_{\pi\circ f}(D,r)-N^{(1)}_{\pi\circ f}(D,r).$
We are now in a position to apply Theorem 3 for a small $\epsilon>0$ (to be determined later) and an ample divisor $A$ on $X$. Consequently, there exists a positive integer $\ell_{1}$ and a proper Zariski closed subset $W_{1}$ of $X$ such that if the orbifold $(X,\Delta_{0})$ has multiplicities at least $\ell_{1}$ along $D_{0}$, then
(6.8) $N_{\pi\circ f}(D,r)-N^{(1)}_{\pi\circ f}(D,r)\leq_{\operatorname{exc}}\epsilon T_{A,\pi\circ f}(r),$
and
(6.9) $T_{D,\pi\circ f}(r)\leq_{\rm exc}N_{\pi\circ f}^{(1)}(D,r)+\epsilon T_{A,\pi\circ f}(r)$
if the image of $\pi\circ f$ is not contained in $W_{1}$. Assuming that $\ell\geq\ell_{1}$ and the image of $f$ is not contained in $\pi^{-1}(W_{1})$, then by (6.7) and (6.8), we have
(6.10) $N_{f}(R,r)\leq_{\operatorname{exc}}\epsilon T_{A,\pi\circ f}(r).$
Since $R\leq\pi^{*}(D)$ (as divisors), the functorial property implies that
(6.11) $m_{f}(R,r)\leq m_{\pi\circ f}(D,r)+O(1)=T_{D,\pi\circ f}(r)-N_{\pi\circ f}(D,r)+O(1)$
Together with (6.9), we have:
(6.12) $m_{f}(R,r)\leq\epsilon T_{A,\pi\circ f}(r)+O(1).$
Combining the two equations (6.10) and (6.12) and using the functorial property of characteristic functions, we have:
(6.13) $T_{R,f}(r)\leq 2\epsilon T_{\pi^{*}A,f}(r)+O(1).$
Since $R$ is a big divisor on $Y$, by [25, Proposition 12.11], there exists a constant $c$ and a proper Zariski-closed subset $W_{0}$ of $Y$, depending only on $A$ and $R$, such that such that
(6.14) $T_{\pi^{*}A,f}(r)\leq cT_{R,f}(r)+O(1),$
if the image of $\mathbf{f}$ is not contained in $W_{0}$. Combining (6.14) with (6.13), we obtain
$T_{R,f}(r)\leq_{\operatorname{exc}}2c\epsilon T_{R,f}(r)+O(1).$
This is not possible as $c$ is independent of $\epsilon$, and $\epsilon$ can be taken sufficiently small. In conclusion, we have found a sufficiently large integer $\ell$ such that if $(Y,\Delta)$ has multiplicities at least $\ell$ along $D_{0}$, then the image of any orbifold entire curve $f$ on $(Y,\Delta)$ is contained in $\pi^{-1}(W_{1})\cup W_{0}$. ∎

### v2

@@
 Consequently, we establish a version of Campana’s orbifold conjecture for finite coverings of projective toric varieties.
-2020 Mathematics Subject Classification: Primary 30D35; Secondary 32Q45 and 32H30
-The research of Min Ru is supported in part by Simon Foundations grant award #531604 and #521604. The second-named author was supported in part by Taiwan’s NSTC grant 113-2115-M-001-011-MY3.
+2020 Mathematics Subject Classification: Primary 32H30; Secondary 32Q45 and 30D35
+The research of Min Ru is supported in part by Simon Foundations grant award #531604 and #521604.
+The research of Julie Tzu-Yueh Wang was supported in part by Taiwan’s NSTC grant 113-2115-M-001-011-MY3.
 1. Introduction
-Let $X$ be an algebraic variety, and $D$ an effective divisor (possibly empty).
-The set $X\backslash D$ is Brody hyperbolic if every entire curve $f:\mathbb{C}\to X\backslash D$ is constant.
-It is Brody quasi-hyperbolic if every entire curve $f:\mathbb{C}\to X\backslash D$ is degenerate, meaning its image is contained a proper Zariski-closed subset $Z$ of $X$.
-We call such $Z$ the exceptional set.
-If $Z$ is independent of $f$, we call $X\backslash D$ Brody strong quasi-hyperbolic.
-This leads us to the Griffiths-Lang’s conjecture.
-Griffiths-Lang’s conjecture.
-Let $X$ be a smooth complex projective variety, $D$ a normal crossing divisor on $X$, and $K_{X}$ the canonical divisor on $X$.
-If $K_{X}+D$ is big, then $X\backslash D$ is Brody quasi-hyperbolic.
+[…]
+[…]
 A more precise quantitative version, in terms of the Nevanlinna’s functions, can be formulated as follows.
@@
 $T_{K_{X}+D,f}(r)\leq_{\rm exc}N_{f}(D,r)+\epsilon T_{A,f}(r)$
-holds for any $\epsilon>0$, where notation $\leq_{\rm exc}$ means the inequality holds for all $r\in(0,\infty)$ except for a set of finite measure.
+holds for any $\epsilon>0$, where $A$ is an ample divisor on $X$, and the notation $\leq_{\rm exc}$ means the inequality holds for all $r\in(0,\infty)$ except for a set of finite measure.
 Here, for each positive integer $n$, $N_{f}^{(n)}(D,r)$ is the $n$-truncated counting function with respect to $D$ given by
 $N_{f}^{(n)}(D,r)=\sum_{0<|z|<r}\mathop{\mathrm{min}}\{{\rm ord}_{z}f^{*}D,n\}\log\frac{r}{|z|}+\mathop{\mathrm{min}}\{{\rm ord}_{0}f^{*}D,n\}\log r,$
-$N_{f}(D,r)$ is the counting function with respect to $D$ given by $N_{f}^{(\infty)}(D,r)$, and $T_{D,f}(r)$ is the (Nevanlinna) height function relative to the divisor $D$ (referring to [25, Section 12]).
-Paul Vojta formulated the following general abc conjecture (see [25, Conjecture 15.2] and [25, Conjecture 23.4]), which extends the above Griffiths’ conjecture.
+$N_{f}(D,r)$ is the counting function with respect to $D$ given by $N_{f}^{(\infty)}(D,r)$, and $T_{D,f}(r)$ is the (Nevanlinna) height function relative to the divisor $D$ (referring to [27, Section 12]).
+Paul Vojta formulated the following general abc conjecture (see [27, Conjecture 15.2] and [27, Conjecture 23.4]), which extends the above Griffiths’ conjecture.
 It is also viewed as the quantitive version of Campana’s orbifold conjecture which will be stated later in the introduction.
-Vojta’s abc conejcture.
+Vojta’s abc conjecture.
+Let $X$ be a smooth complex projective variety, $D$ a normal crossing divisor on $X$, and $K_{X}$ a canonical divisor on $X$.
 For any $\epsilon>0$, there exists a proper Zariski-closed subset $Z$ of $X$ such that for any entire curve $f:\mathbb{C}\to X$ whose image is not contained in $Z$, the inequality
 $T_{K_{X}+D,f}(r)\leq_{\rm exc}N_{f}^{(1)}(D,r)+\epsilon T_{A,f}(r)$
-holds, where $A$ is an ample divisor on $X$.
-Note that Vojta’s abc conjecture strengthens Griffiths’ conjecture in two key ways: (i) the exceptional set is independent of the map $f$, and (ii) the counting function is truncated at level one.
+holds for any $\epsilon>0$, where $A$ is an ample divisor on $X$.
+Note that Vojta’s abc conjecture strengthens Griffiths’ conjecture in two key ways: (i) the exceptional set $Z$ above is independent of the map $f$, and (ii) the counting function is truncated at level one.
 Proving these two features is quite challenging, even if the aforementioned version of Griffiths’ conjecture could be established.
 In this paper, we first prove Vojta’s abc conjecture for entire curves in ${\mathbb{P}}^{n}({\mathbb{C}})$, assuming the map is highly ramified over the coordinate hyperplanes.
-The precise statement of our main result is as follows.
+The precise statement of our main result is stated below.
 Additionally, we highlight that the exceptional set obtained in the theorem can be determined explicitly-this is a challenging issue in the study of Nevanlinna theory and complex hyperbolicity.
@@
 Here, the evaluation of $G$ is taken at $\mathbf{f}=(f_{0},\ldots,f_{n})$ in a reduced form of $f$, meaning that $f_{0},\ldots,f_{n}$ are entire functions without common zeros.
-Note that the case when $n=2$ is due to Ji Guo and the second named author ([9, Theorem 1.4]).
-The new idea in this paper is motivated by the recent paper [2] by considering the parabolic Riemann surface $\mathcal{Y}_{f}:=\mathbb{C}\setminus{f^{-1}(D)}$, reducing the problem to the omitting case.
-To achieve this, we need to extend the results in [9] and [7] to the parabolic setting.
+We note that this theorem corresponds to a case of Vojta’s abc conjecture for entire curves with $D=H_{0}+\cdots+H_{n}+[G=0]$, so that $K_{\mathbb{P}^{n}+D}$ is linearly equivalent to $[G=0]$.
+The case $n=2$ was previously established by Ji Guo and the second named author ([9, Theorem 1.4]).
+The new idea in this paper is motivated by the recent paper [2], where the authors considered the parabolic Riemann surface $\mathcal{Y}_{f}:=\mathbb{C}\setminus{f^{-1}(D)}$, reducing the problem to the omitting case.
+To achieve this in our setting, we need to extend the results in [7] and [9] to the parabolic setting.
 We state some consequences of our Main Theorem, starting with the introduction of Campana’s orbifolds.
@@
   1. (i)
-    The Campana’s conjecture is a consequence of the above Vojta’s strong abc conecutre.
+    The Campana’s conjecture is a consequence of the above Vojta’s abc conjecutre.
   2. (ii)
-    Let $D:={\rm Supp}(\Delta)=D_{1}+\cdots+D_{q}$.
+    Given an orbifold divisor $\Delta$ on $X$, let $D:={\rm Supp}(\Delta)=D_{1}+\cdots+D_{q}$.
 Recall that the pair $(X,D)$ is said to be of log-general type if the divisor $K_{X}+D$ is big.
-It’s clear that $(X,D)$ is of log-general type if $(X,\Delta)$ is of general type, i.e.
+It is clear that $(X,D)$ is of log-general type if $(X,\Delta)$ is of general type, i.e.
 $K_{X}+\Delta$ is big.
 On the other hand, the condition that $(X,D)$ is of log-general type implies that $(X,\Delta)$ is of general type if $(X,\Delta)$ has a sufficiently large multiplicity along each $D_{i}$, $1\leq i\leq q$.
-((((See [12, Corollary 2.2.24].))))
+((See [11, Corollary 2.2.24].))
 The Main Theorem leads to the following two theorems concerning Campana’s orbifold conjecture; however, we will not include the proofs, as they can be adapted from those in [9] for the case $n=2$.
@@
 Theorem 1.
-Let $\Delta_{0}$ be an orbifold divisor of ${\mathbb{P}}^{n}({\mathbb{C}})$ and let $H_{0},H_{1},\dots,H_{n}$ be the coordinate hyperplanes in ${\mathbb{P}}^{n}({\mathbb{C}})$, such that $\Delta_{0}$ and $H_{0},H_{1},\dots,H_{n}$ are in general position.
+Let $\Delta_{0}$ be an orbifold divisor on ${\mathbb{P}}^{n}({\mathbb{C}})$ and let $H_{0},H_{1},\dots,H_{n}$ be the coordinate hyperplanes in ${\mathbb{P}}^{n}({\mathbb{C}})$, such that $\Delta_{0}$ and $H_{0},H_{1},\dots,H_{n}$ are in general position.
 Let $m_{i}\in(1,\infty)\cap{\mathbb{Q}}$, $0\leq i\leq n$, and $\Delta=\Delta_{0}+(1-{1\over m_{0}})H_{0}+(1-{1\over m_{1}})H_{1}+\cdots+(1-{1\over m_{n}})H_{n}$.
@@
 Then there exist a proper Zariski closed subset $W$ of ${\mathbb{P}}^{n}({\mathbb{C}})$ and an effectively computable positive integer $\ell$ such that the image of any orbifold entire curve $f:{\mathbb{C}}\rightarrow({\mathbb{P}}^{n}({\mathbb{C}}),\Delta)$ with $\mathop{\mathrm{min}}\{m_{0},m_{1},\dots,m_{n}\}\geq\ell$ must be contained in $W$.
-The next case involves $n+1$ components, not all lines, in $\mathbb{P}^{n}({\mathbb{C}})$.
+The next case involves $n+1$ components, not all are hyperplanes, in $\mathbb{P}^{n}({\mathbb{C}})$.
 Theorem 2.
-Let $F_{i}$, $11\leq i\leq n+$ 1, be homogeneous irreducible polynomials of positive degrees in $\mathbb{C}[x_{0},\ldots,x_{n}]$.
-Assume that $D_{i}:=[F_{i}=0]\subset\mathbb{P}^{n}(\mathbb{C})$, $11\leq i\leq n+$ 1, intersect transversally.
+Let $F_{i}$, $1\leq i\leq n+1$, be homogeneous irreducible polynomials of positive degrees in $\mathbb{C}[x_{0},\ldots,x_{n}]$.
+Assume that $D_{i}:=[F_{i}=0]\subset\mathbb{P}^{n}(\mathbb{C})$, $1\leq i\leq n+1$, intersect transversally.
 Let $\Delta=(1-{1\over m_{1}})D_{1}+\cdots+(1-{1\over m_{n+1}})D_{n+1}$ be an orbifold divisor of $\mathbb{P}^{n}({\mathbb{C}})$.
@@
 Then, there exist an effectively computable positive integer $\ell$ and a proper Zariski closed subset $W$ of $\mathbb{P}^{n}({\mathbb{C}})$ such that the image of any orbifold entire curve $f:{\mathbb{C}}\rightarrow({\mathbb{P}}^{n}({\mathbb{C}}),\Delta)$ with $\mathop{\mathrm{min}}\{m_{1},\dots,m_{n+1}\}\geq\ell$ must be contained in $W$.
-Note that $\mathbb{P}^{n}({\mathbb{C}})$ is a toric variety when we identify ${\mathbb{G}}_{m}^{n}$ with ${\mathbb{P}}^{n}({\mathbb{C}})\backslash\cup_{i=0}^{n}H_{i}$, where $H_{0},\dots,H_{n}$ are the $n+1$ coordinate hyperplanes.
-A toric variety $X$ is defined as a variety over ${\mathbb{C}}$ that contains ${\mathbb{G}}_{m}^{n}$ as a dense open subvariety, with the action of ${\mathbb{G}}_{m}^{n}$ on itself extends to an algebraic action of ${\mathbb{G}}_{m}^{n}$ on $X$.
-We refer to [5] for more details on toric varieties.
-Using the arguments in [13] and [6], we can extend our Main Theorem to projective toric varieties.
-Recall from [13] that an “admissible pair” is a couple $(X,V)$ where $V$ is a nonsingular variety embedded in a nonsingular projective variety $X$ in such a way that $D_{0}=X\setminus V$ is a normal crossings divisor.
-A Campana orbifold $(X,\Delta)$ is said to be associated to the admissible pair $(X,V)$ if $D_{0}$ is the support of $\Delta$.
+[…]
 Theorem 3.
@@
 Let $A$ be a big divisor on $X$.
-For every $\epsilon>0$ there is a positive integral number $\ell$ and a proper Zariski-closed subset $Z$ of $X$ such that, for any Campana orbifold $(X,\Delta)$ associated to the admissible pair $(X,\mathbb{G}_{m}^{n})$ with multiplicities at least $\ell$ along the boundary $D_{0}$, the following two inequalities hold:
+Then, for every $\epsilon>0$, there is a positive integer $\ell$ and a proper Zariski-closed subset $Z$ of $X$ such that, for any Campana orbifold $(X,\Delta)$ associated to the admissible pair $(X,\mathbb{G}_{m}^{n})$ with multiplicities at least $\ell$ along the boundary $D_{0}$, the following two inequalities hold:
   1. (i)
@@
 Assume that $\pi(R)$ and $D_{0}$ are in general position on $X$.
-Then there exists a positive integer $\ell$ such that if the orbifold pair $(Y,\Delta)$ is of general type with ${\rm Supp}(\Delta)={\rm Supp}(H)$ and multiplicity at least $\ell$ for each component, then there exists a proper Zariski closed subset $W$ of $Y$ such that the image of any non-constant orbifold entire curve $f:{\mathbb{C}}\rightarrow(Y,\Delta)$ must be contained in $W$.
+Then there exists a positive integer $\ell$ such that, if the orbifold pair $(Y,\Delta)$ is of general type with ${\rm Supp}(\Delta)={\rm Supp}(H)$ and multiplicity at least $\ell$ for each component, then there exists a proper Zariski closed subset $W$ of $Y$ such that the image of any non-constant orbifold entire curve $f:{\mathbb{C}}\rightarrow(Y,\Delta)$ must be contained in $W$.
+The paper is organized as follows.
 In Section 2, we prove a version of the Main Theorem for rational maps and explain the motivation for using parabolic Riemann surfaces.
@@
 In [7], a version of Vojta’s abc conjecture for algebraic tori over algebraic function fields is established.
-As we will see, the algebraic reduction method presented in this article is a key technical component in proving the Main Theorem.
+As we will see, the algebraic reduction method presented in [7] is a key technical component in proving our Main Theorem.
 The core idea behind the proof is also inspired by the function field approach.
@@
 2.1. The Rational Case
-The statements and arguments apply verbatim to algebraic function fields.
+The statements and arguments apply verbatim to algebraic function fields (see also [6, Theorem 1.13]).
 Here, let $K=\mathbb{C}(t)$ be the function field of $\mathbb{P}^{1}({\mathbb{C}})$.
@@
 We use the notation $N_{S}(f)$ (respectively $N_{S}^{(1)}(f)$) to count the number of zeros of $f$ with multiplicity (respectively, without multiplicity) outside of $S$.
-When $S$ is empty, we simply write $N(f)$ (respectively $N^{(1)}(f)$) to count the number of zeros with multiplicity (respectively, without multiplicity) on $\mathbb{P}^{1}(\mathbb{C})$.
+When $S$ is empty, we simply write $N(f)$, which equals $\deg f$, to count the number of zeros with multiplicity on $\mathbb{P}^{1}(\mathbb{C})$; $N^{(1)}(f)$ denotes the same count without multiplicity.
 Recall a simplified version of the main technical theorem in [7].
@@
 Here $c_{0}$ can be effectively bounded from above in terms of $\epsilon$, $n$, and the degree of $G$.
-Moreover, the exceptional set $Z$ can be expressed as the zero locus of a finite set $\Sigma\subset\mathbb{C}[x_{1},\ldots,x_{n}]$ with the following properties: (Z1) $\Sigma$ depends on $\epsilon$ and $G$ and can be determined explicitly and (Z2) $|\Sigma|$ and the degree of each polynomial in $\Sigma$ can be effectively bounded from above in terms of $\epsilon$, $n$, and the degree of $G$.
+Moreover, the exceptional set $Z$ can be expressed as the zero locus of a finite set $\Sigma\subset\mathbb{C}[x_{1},\ldots,x_{n}]$ with the following properties: (Z1) $\Sigma$ depends on $\epsilon$ and $G$ but independent of $S$ and can be determined explicitly and (Z2) $|\Sigma|$ and the degree of each polynomial in $\Sigma$ can be effectively bounded from above in terms of $\epsilon$, $n$, and the degree of $G$.
+The construction of the set of polynomials $\Sigma$ follows from the specialization lemmas [7, Lemma 15 and 16].
+In particular, it is evident that $\Sigma$ is independent of the choice of $S$.
+Note that $\mathbf{u}:=(u_{1},\ldots,u_{n})\in({\mathcal{O}}_{S}^{*})^{n}\setminus Z$ means that $F(\mathbf{u})\neq 0$ (a nontrivial function and hence not identically zero) for every $F\in\Sigma$.
+This formulation is closely analogous to the number field case.
+Alternatively, one may view $\mathbf{u}$ as a morphism: $\mathbb{P}^{1}(\mathbb{C})\setminus S\to\mathbb{C}^{n}$; then the condition $F(\mathbf{u})\neq 0$ for every $F\in\Sigma$ is equivalent to saying that the image $\mathbf{u}(\mathbb{P}^{1}(\mathbb{C})\setminus S)$ is not contained in $Z$.
 We will demonstrate how to apply the above theorem to show the following version of Main Theorem for rational case.
@@
     $N^{(1)}(G(\mathbf{f}))\geq(\deg G-\epsilon)\deg\mathbf{f}$,
-for any non-constant entire rational curve $\mathbf{f}=(f_{0},\ldots,f_{n})$, where $f_{0},\ldots,f_{n}\in\mathbb{C}[z]$ are polynomials without common zeros, provided that $\mathbf{f}(\mathbb{C})\not\subset Z$ and ${\rm ord}_{\mathbf{p}}(f_{i})\geq\ell$ if ${\rm ord}_{\mathbf{p}}(f_{i})>0$ for $0\leq i\leq n$, and for all $\mathbf{p}\in\mathbb{P}^{1}({\mathbb{C}})$.
+for any non-constant entire rational curve $\mathbf{f}=(f_{0},\ldots,f_{n})$, where $f_{0},\ldots,f_{n}\in\mathbb{C}[z]$ are polynomials without common zeros, provided that $\mathbf{f}(\mathbb{C})\not\subset Z$ and that the following condition is satisfied: for each $0\leq i\leq n$ and $\mathbf{p}\in\mathbb{P}^{1}({\mathbb{C}})$, if ${\rm ord}_{\mathbf{p}}(f_{i})>0$, then ${\rm ord}_{\mathbf{p}}(f_{i})\geq\ell$.
 Here, $\deg\mathbf{f}=\max\{\deg f_{0},\ldots,\deg f_{n}\}$.
@@
 where $\tilde{c}=\frac{1}{2}\binom{n+\deg G}{n}(\binom{n+\deg G}{n}+1)$.
+The proof of Theorem 6 (i) can be outline as the follows.
+Given a rational map $\mathbf{f}=(f_{0},\ldots,f_{n})$, where $f_{0},\ldots,f_{n}\in\mathbb{C}[z]$ are polynomials without common zeros, we define a set $S_{\mathbf{f}}$ consisting of the zeros of all $f_{i}$.
+With this choice, each coordinate function $f_{i}$ becomes an $S_{\mathbf{f}}$-unit, so that Theorem 5 (ii) can be applied to estimate the multiplicities of the zeros of $G(\mathbf{f})$ outside $S_{\mathbf{f}}$.
+We then use Proposition 7 to estimate the zeros of $G(\mathbf{f})$ inside $S_{\mathbf{f}}$.
+Crucially, the constants and exceptional sets appearing in Theorem 5 and Proposition 7 do not depend on $S$, allowing this method to be used for all rational maps.
 Proof of Theorem 6.
 Denote by $\ell$ a sufficiently large integer to be determined later.
-Let $\mathbf{f}=(f_{0},\ldots,f_{n})$, where $f_{0},\ldots,f_{n}\in\mathbb{C}[z]$ are polynomials without common zeros such that ${\rm ord}_{\mathbf{p}}(f_{i})\geq\ell$ if ${\rm ord}_{\mathbf{p}}(f_{i})>0$ for $0\leq i\leq n$ and $\mathbf{p}\in{\mathbb{P}}^{1}(\mathbb{C})$.
+Let $\mathbf{f}=(f_{0},\ldots,f_{n})$, where $f_{0},\ldots,f_{n}\in\mathbb{C}[z]$ are polynomials without common zeros.
+By the assumption, for each $0\leq i\leq n$ and $\mathbf{p}\in\mathbb{P}^{1}({\mathbb{C}})$, if ${\rm ord}_{\mathbf{p}}(f_{i})>0$, then ${\rm ord}_{\mathbf{p}}(f_{i})\geq\ell$.
 Let
@@
 and $f_{i}\in\mathcal{O}_{\tilde{S}_{\mathbf{f}}}^{*}$ for $0\leq i\leq n$.
+Let $\widetilde{G}(x_{1},\ldots,x_{n}):=G(1,x_{1},\ldots,x_{n})$ and $u_{i}=\frac{f_{i}}{f_{0}}\in\mathcal{O}_{\tilde{S}_{\mathbf{f}}}^{*}$, $1\leq i\leq n$.
+Then $\widetilde{G}(u_{1},\ldots,u_{n}):=f_{0}^{-\deg G}G({\mathbf{f}}).$ An elementary argument shows that
+(2.3) $\max_{1\leq i\leq n}\{\deg u_{i}\}\leq\max_{0\leq i\leq n}\{\deg f_{i}\}\leq\sum_{j=1}^{n}\deg u_{j}.$
 Let $\epsilon>0$ be given.
-We now apply Theorem 5 to find a positive integer $c_{0}$ and a proper Zariski closed subset $Z_{0}$ independent of $\tilde{S}_{\mathbf{f}}$ of $\mathbb{A}^{n}$ such that we have either the inequality
+We now apply Theorem 5 for $\widetilde{G}\in\mathbb{C}[x_{1},\ldots,x_{n}]$ and $\tilde{S}_{\mathbf{f}}$ to find a positive integer $c_{0}$ and a proper Zariski closed subset $Z_{0}$ of $\mathbb{A}^{n}$ independent of $\tilde{S}_{\mathbf{f}}$ such that we have either the inequality
 (2.4) $\max_{0\leq i\leq n}\{\deg f_{i}\}(=\deg\mathbf{f})\leq c_{0}\max\{1,|\tilde{S}_{\mathbf{f}}|-2\}$
 holds or the following two inequalities hold:
-(2.4) $N_{\tilde{S}_{\mathbf{f}}}(G_{0}(u_{1},\ldots,u_{n}))-N_{\tilde{S}_{\mathbf{f}}}^{(1)}(G_{0}(u_{1},\ldots,u_{n}))\leq\epsilon\deg\mathbf{f},$
-(2.5) $N_{\tilde{S}_{\mathbf{f}}}^{(1)}(G_{0}(u_{1},\ldots,u_{n}))\geq(\deg G-\epsilon)\cdot\deg\mathbf{f},$
+(2.5) $N_{\tilde{S}_{\mathbf{f}}}(G({\mathbf{f}}))-N_{\tilde{S}_{\mathbf{f}}}^{(1)}(G({\mathbf{f}}))=N_{\tilde{S}_{\mathbf{f}}}(\widetilde{G}(u_{1},\ldots,u_{n}))-N_{\tilde{S}_{\mathbf{f}}}^{(1)}(\widetilde{G}(u_{1},\ldots,u_{n}))\leq\epsilon\deg\mathbf{f},$
+(2.6) $N_{\tilde{S}_{\mathbf{f}}}^{(1)}(G({\mathbf{f}}))=N_{\tilde{S}_{\mathbf{f}}}^{(1)}(\widetilde{G}(u_{1},\ldots,u_{n}))\geq(\deg G-\epsilon)\cdot\deg\mathbf{f},$
 if $\mathbf{f}(\mathbb{C})\not\subset\tilde{Z}_{0}$, where $\tilde{Z}_{0}$ is the projective closure of $Z_{0}$ in $\mathbb{P}^{n}({\mathbb{C}})$.
-Here $G_{0}(x_{1},\ldots,x_{n}):=G(1,x_{1},\ldots,x_{n})$ and $u_{i}=\frac{f_{i}}{f_{0}}\in\mathcal{O}_{\tilde{S}_{\mathbf{f}}}^{*}$, $1\leq i\leq n$.
-Since the constant $c_{0}$ and the exceptional set $Z$ in Theorem 5 do not depend on $S$, it is important to note that $c_{0}$ and $Z_{0}$ do not depend on $\mathbf{f}$ as well.
+Since the constant $c_{0}$ and the exceptional set $Z$ in Theorem 5 do not depend on $S$, it is important to note that $c_{0}$ and $\tilde{Z}_{0}$ do not depend on $\mathbf{f}$ as well.
 We first show that (2.4) does not hold if $\ell>2c_{0}n$.
@@
 $\max_{0\leq i\leq n}\{\deg f_{i}\}\leq\frac{(n+1)c_{0}}{\ell}\max_{0\leq i\leq n}\{\deg f_{i}\}+c_{0}<\frac{1}{2}\max_{0\leq i\leq n}\{\deg f_{i}\}+c_{0}.$
-Therefore, $\max_{0\leq i\leq n}\{\deg f_{i}\}<2c_{0}.$ This is not possible since not all the polynomials $f_{i}$ are constant, the assumption that ${\rm ord}_{\mathbf{p}}(f_{i})\geq\ell$ if ${\rm ord}_{\mathbf{p}}(f_{i})>0$ for $0\leq i\leq n$ and $\mathbf{p}\in{\mathbb{P}}^{1}(\mathbb{C})$ implies that $2nc_{0}<\ell<\max_{0\leq i\leq n}\{\deg f_{i}\}$.
+Therefore, $\max_{0\leq i\leq n}\{\deg f_{i}\}<2c_{0}.$ This is not possible since the $f_{i}$ are not all constant, and by our assumption that every zero has multiplicity at least $\ell$, we have $2nc_{0}<\ell<\max_{0\leq i\leq n}\{\deg f_{i}\}$.
 We continue to let $\ell>2c_{0}n$.
 As we have shown that (2.4) does not hold, it follows that (2.5) and (2.6) both hold if $\mathbf{f}(\mathbb{C})\not\subset\tilde{Z}_{0}$.
-Since $N^{(1)}(G_{0}(u_{1},\ldots,u_{n}))\geq N_{\tilde{S}_{\mathbf{f}}}^{(1)}(G_{0}(u_{1},\ldots,u_{n}))$, it follows trivially from (2.5) that
-(2.6) $N^{(1)}(G(\mathbf{f}))\geq N^{(1)}(G_{0}(u_{1},\ldots,u_{n}))\geq(\deg G-\epsilon)\cdot\deg\mathbf{f}.$
+Then it follows trivially from (2.6) that
+(2.7) $N^{(1)}(G(\mathbf{f}))\geq N_{\tilde{S}_{\mathbf{f}}}^{(1)}(G({\mathbf{f}}))\geq(\deg G-\epsilon)\cdot\deg\mathbf{f}.$
 This shows (ii) holds.
@@
 Assume that $\mathbf{f}(\mathbb{C})\not\subset W$.
-Since $G_{0}(0,\dots,0)=G(1,0,\ldots,0)\neq 0$, we now apply Propositon 7 to get
-$\sum_{\mathbf{p}\in\tilde{S}_{\mathbf{f}}}{\rm ord}^{+}_{\mathbf{p}}(G_{0}(u_{1},\ldots,u_{n}))\leq\tilde{c}\max\{1,|\tilde{S}_{\mathbf{f}}|-2\},$
+Since $\widetilde{G}(0,\dots,0)=G(1,0,\ldots,0)\neq 0$, we now apply Propositon 7 to get
+$\sum_{\mathbf{p}\in\tilde{S}_{\mathbf{f}}}{\rm ord}^{+}_{\mathbf{p}}(\widetilde{G}(u_{1},\ldots,u_{n}))\leq\tilde{c}\max\{1,|\tilde{S}_{\mathbf{f}}|-2\},$
 where $\tilde{c}=\frac{1}{2}\binom{n+degG}{n}(\binom{n+\deg G}{n}+1)$. Together with (2.2) and that $\ell\leq\max_{0\leq i\leq n}\{\deg f_{i}\}$, it yields
-$m_{\tilde{S}_{\mathbf{f}}}(G_{0}(u_{1},\ldots,u_{n})):=\sum_{\mathbf{p}\in\tilde{S}_{\mathbf{f}}}{\rm ord}^{+}_{\mathbf{p}}(G_{0}(u_{1},\ldots,u_{n}))\leq\frac{(n+1)\tilde{c}}{\ell}\max_{0\leq i\leq n}\{\deg f_{i}\}\leq\epsilon\deg\mathbf{f}$
+$m_{\tilde{S}_{\mathbf{f}}}(\widetilde{G}(u_{1},\ldots,u_{n})):=\sum_{\mathbf{p}\in\tilde{S}_{\mathbf{f}}}{\rm ord}^{+}_{\mathbf{p}}(\widetilde{G}(u_{1},\ldots,u_{n}))\leq\frac{(n+1)\tilde{c}}{\ell}\max_{0\leq i\leq n}\{\deg f_{i}\}\leq\epsilon\deg\mathbf{f}$
 provided $\ell>(n+1)\tilde{c}_{0}\epsilon^{-1}.$ Combining with (2.5), we have
-(2.7) $N(G_{0}(u_{1},\ldots,u_{n}))-N^{(1)}(G_{0}(u_{1},\ldots,u_{n}))\leq 2\epsilon\deg\mathbf{f}.$
+(2.8) $N(\widetilde{G}(u_{1},\ldots,u_{n}))-N^{(1)}(\widetilde{G}(u_{1},\ldots,u_{n}))\leq 2\epsilon\deg\mathbf{f}.$
 We now repeat the above arguments for $(\frac{f_{0}}{f_{i}},\cdots,\frac{f_{n}}{f_{i}})$, $0\leq i\leq n$.
@@
 2.2. Motivation for the proof of the Main Theorem
-As in Section 2.1, the starting point of the proof of Theorem 6 is as the follows: For a given rational map, we define a set $S$ (depending on the map) that contains the zeros and poles of each coordinate function of the map.
-In this way, all the coordinate functions become $S$-units, and so we can apply Theorem 5.
-Importantly, the constants and the exceptional set obtained in Theorem 5 do not depend on $S$, allowing this method to be used for all rational maps.
-To apply this view point to the analytic case, we consider $f=(f_{0},\ldots,f_{n})$, where $f_{i}$, $0\leq i\leq n$, are entire functions without common zeros.
-Let $\mathcal{E}_{f}$ denote the set of all the zeros of $f_{i}$, $0\leq i\leq n$, and define $\mathcal{Y}_{f}:=\mathbb{C}\setminus\mathcal{E}_{f}$.
-Then ${\mathcal{Y}_{f}}$ is a parabolic open Riemann surface.
-On ${\mathcal{Y}_{f}}$, the restriction $\tilde{f}_{i}:=f_{i}|_{\mathcal{Y}_{f}}$ becomes a unit on ${\mathcal{Y}}$ (i.e.
-it has neither zeros nor poles).
-This perspective was also adopted in [2], where they laid important foundational work.
-For our purposes, we will need to develop an analogue of Theorem 5 over parabolic open Riemann surfaces.
+In Section 2.1, the key to proving Theorem 6 is the use of the set $S$ (depending on the map) that contains all zeros and poles of each coordinate function of the map.
+The crucial observation is that all coordinate functions then become $S$-units, allowing us to apply Theorem 5.
+Importantly, the constants and the exceptional set obtained in Theorem 5 are independent of $S$, which makes the argument work uniformly for all rational maps.
+To extend this viewpoint to the analytic setting, we let $\mathcal{E}_{f}$ denote the set of all zeros of $f_{i}$, $0\leq i\leq n$, and define
+$\mathcal{Y}_{f}:=\mathbb{C}\setminus\mathcal{E}_{f},$
+where $f=(f_{0},\ldots,f_{n})$ is a holomorphic map such that the entire functions $f_{i}$, $0\leq i\leq n$, have no common zeros.
+Then $\mathcal{Y}_{f}$ is a parabolic open Riemann surface equipped with a smooth exhaustion function $\sigma$ for which the smooth $(1,1)$-form $dd^{c}\!\log\sigma$ has finite total mass on $\mathcal{Y}_{f}$.
+Although $\sigma$ may not be a parabolic exhaustion for ${\mathcal{Y}}_{f}$, Chen et al.
+(see [2]) observed that, after appropriate modifications, the key ingredients of Nevanlinna theory still admit analogues in this setting.
 3. Nevanlinna theory and the GCD theorem on parabolic Riemann surfaces
-3.1. Parabolic Riemann surfaces
+3.1. Parabolic Riemann surfaces and Jensen’s formula
 […]
-1. For a hypersurface $D=\{Q=0\}$ in $\mathbb{P}^{n}({\mathbb{C}})$, where $Q$ is a homogeneous polynomial of degree $d$, the proximity function is defined as
-[…]
-where $\lambda_{D}$ is the Weil-function on ${\mathbb{P}}^{n}({\mathbb{C}})$ given by
-$\lambda_{D}({\bf x})=\log{\|{\bf x}\|^{d}\over|Q({\bf x})|},~{}~{}{\bf x}\in\mathbb{P}^{n}({\mathbb{C}}).$
-2. The characteristic (order) function is defined by
-$T_{f,{\mathcal{Y}}}(r)=\int_{1}^{r}{dt\over t}\int_{B(t)}f^{*}\omega_{FS}.$
-3. The counting function and the $k$-truncated counting function are defined by
-$N_{f,{\mathcal{Y}}}(D,r):=\int_{1}^{r}\sum_{z\in B(t)}{\rm ord}_{z}f^{*}D{dt\over t},~{}~{}~{}~{}~{}~{}~{}\mbox{and}~{}~{}N_{f,{\mathcal{Y}}}^{(k)}(D,r):=\int_{1}^{r}\sum_{z\in B(t)}\mathop{\mathrm{min}}\{k,{\rm ord}_{z}f^{*}D\}{dt\over t}.$
-4. The counting function extends naturally to the sub-schemes, as well as the notion $N_{\text{gcd},{\mathcal{Y}}}(f,g)(r)$, where $f,g$ are meromorphic functions on $\mathcal{Y}$. See [14].
-Parabolic First Main Theorem.
-Let $D$ be a hypersurface in ${\mathbb{P}}^{n}({\mathbb{C}})$ and $f:{\mathcal{Y}}\rightarrow\mathbb{P}^{n}({\mathbb{C}})$ be a holomorphic map with $f({\mathcal{Y}})\not\subset D$.
-Then
-$T_{f,{\mathcal{Y}}}(r)=m_{f,{\mathcal{Y}}}(D,r)+N_{f,{\mathcal{Y}}}(D,r)+O(1).$
-It is important to note that, when we write $O(g(x))$ for some function $g$, we mean that $Cg(x)$ where $C$ is a constant independent of ${\mathcal{Y}}$ throughout the paper.
-Fix a nowhere vanishing global holomorphic vector field $\xi\in\Gamma({\mathcal{Y}},T_{\mathcal{Y}})$ (we denote it by ${\partial\over\partial z}$) on ${\mathcal{Y}}$.
-Such a vector field exists because ${\mathcal{Y}}$ is non-compact and consequently its holomorphic tangent bundle $T_{\mathcal{Y}}$ is holomorphically trivial.
-From Lemma 4.4 in [10],
-$\int_{S_{r}}\log\left|d\sigma\left({\partial\over\partial z}\right)\right|^{2}d\mu_{r}=-\int_{1}^{r}\chi_{\sigma}(t)\frac{dt}{t}+2\varsigma\log r+O(1),$
-where $\chi_{\sigma}(r)$ is the Euler characteristic of $B(r)$, and
-[…]
-which is independent of $r$ for $r$ being sufficiently large. Denote by
-$\mathfrak{X}^{+}_{\sigma}(r):=\int_{S(r)}\log^{+}\left|d\sigma\left({\partial\over\partial z}\right)\right|^{2}d\mu_{r}.$
-[…]
-(3.1) $\hat{\tau}:=\log\hat{\sigma}:=\log^{+}|z|+\sum_{j\geq 1}r_{j}\log^{+}{r_{j}\over|z-a_{j}|}.$
-Note that $\hat{\tau}$ takes values in $[0,+\infty)$, is continuous on ${\mathcal{Y}}$ and is smooth (indeed harmonic) outside the circle $S(0,1):=\{|z|=1\}$ and the disjoint circles $S(a_{j},r_{j}):=\{|z-a_{j}|=r_{j}\}$.
-By the Poincare-Lelong formula, it is clear that
-$dd^{c}\hat{\tau}={1\over 2}\nu(0,1)+{r_{j}\over 2}\sum_{j=1}^{\infty}\left(\nu(a_{j},r_{j})-\delta_{a_{j}}\right)$
-is a distribution of order 00 and locally of finite mass, where $\nu(a_{j},r_{j})$ is the Haar measure on the circle $S(a_{j},r_{j})$.
-The smoothing process was done in [2] as follows: as in the Appendix in [2], let
+Let us first briefly review the construction of $\sigma$ given in Lemma 3.1 and 3.2 of [2].
+Let $\{\mathcal{Y}$$\mathbb{C}$$\backslash\mathcal{E}$, where ${\mathcal{E}}:=\{a_{j}\}_{j=1}^{\infty}$ is a discrete countable set of points in $\mathbb{C}$}.
+Up to translation and dilation on $\mathbb{C}$}, we may assume that $|a_{j}|>2$ for all $j\geq 1$.
+We also assume that the radii satisfy the following conditions: (i) ${\mathbb{D}}(a_{j},2r_{j})$ for all $j$ and the disc ${\mathbb{D}}_{2}$ are disjoint, and (ii) the sum $\sum_{j\geq 1}r_{j}<+\infty$.
+The exhaustion function $\sigma$ on $\mathcal{Y}$} is obtained by smoothing the piecewise smooth exhaustion function $\hat{\sigma}$ given by
+(3.1) $\log\hat{\sigma}:=\log^{+}|z|+\sum_{j\geq 1}r_{j}\log^{+}{r_{j}\over|z-a_{j}|}.$
+Note that $\hat{\tau}:=\log\hat{\sigma}$ takes values in $0,+\infty)$, is continuous on $\mathcal{Y}$} and is smooth (indeed harmonic) outside the circle $S(0,1):=\{|z|=1\}$ and the disjoint circles $S(a_{j},r_{j}):=\{|z-a_{j}|=r_{j}\}$.
+By the Poincaré-Lelong formula (see, e.g., [17]), it is clear that
+$dd^{c}\hat{\tau}={1\over 2}\nu(0,1)+{1\over 2}\sum_{j=1}^{\infty}r_{j}\left(\nu(a_{j},r_{j})-\delta_{a_{j}}\right)$
+is a distribution of order 0 and locally of finite mass, where $\nu(a_{j},r_{j})$ is the Haar measure on the circle $S(a_{j},r_{j})$.
+The smoothing process goes as follows (see [2, Appendix]): Let
 $h(r)=\begin{cases}0&\text{ for $r\leq{3\over 4}$}\\ {1\over 4\pi}{1\over 1+e^{{r-1\over(r-1)^{2}-1/16}}}&\text{for ${3\over 4}<r<{5\over 4}$}\\ {1\over 4\pi}&\text{for $r\geq{5\over 4}$}\end{cases}$
 and
-$H(r):=4\pi\int_{0}^{r}h(s-c){ds\over s},$
-where the constant $c\in(0,1/2)$ is chosen such that $H$ agrees with $\log^{+}r$ outside $[{1\over 2},{3\over 2}]$.
+(3.2) $H(r):=4\pi\int_{0}^{r}h(s-c){ds\over s},$
+where the constant $c\in(0,1/2)$ is chosen such that $H$ agrees with $\log^{+}r$ outside ${1\over 2},{3\over 2}]$.
 Note that $H$ is an absolute function independent of $a_{j}$.
@@
 Define
-$\tau:=\log\sigma:=H(z)+\sum_{j=1}^{\infty}r_{j}H\left({r_{j}\over|z-a_{j}|}\right).$
-Then $\sigma\geq\hat{\sigma}$ and the difference $\sigma-\hat{\sigma}$ is supported on
-$\text{Supp}(\sigma-\hat{\sigma})\subset U:=\left(A\left(0,{1\over 2},{3\over 2}\right)\backslash{\mathcal{E}}\right)\bigcup_{j=1}^{\infty}A\left(a_{j},{1\over 2}r_{j},{3\over 2}r_{j}\right),$
+(3.3) $\tau:=H(|z|)+\sum_{j=1}^{\infty}r_{j}H\left({r_{j}\over|z-a_{j}|}\right),~~~~\sigma=\exp(\tau).$
+Then, as stated in [2, Lemma 3.1], $\sigma\geq\hat{\sigma}$ and the difference $\sigma-\hat{\sigma}$ is supported on
+(3.4) $\text{Supp}(\sigma-\hat{\sigma})\subset U:=\left(A\left(0,{1\over 2},{3\over 2}\right)\backslash{\mathcal{E}}\right)\bigcup\left(\bigcup_{j=1}^{\infty}A\left(a_{j},{1\over 2}r_{j},{3\over 2}r_{j}\right)\right),$
 where $A(a_{j},{1\over 2}r_{j},{3\over 2}r_{j}):=\{z\in{\mathcal{Y}}:{1\over 2}r_{j}\leq|z-a_{j}|\leq{3\over 2}r_{j}\}$ are pairwise disjoint annuli.
 For $z\not\in\cup_{j=1}^{\infty}\overline{{\mathbb{D}}(a_{j},{3\over 2}r_{j})}$, $\log^{+}\left|{r_{j}\over z-a_{j}}\right|=H\left(\left|{r_{j}\over z-a_{j}}\right|\right)=0$ for each $j$.
-Thus $\tau(z)=H(|z|)$ and $\hat{\tau}(z)=\log^{+}|z|$, which are eqaual when $\tau(z)\geq\log{3\over 2}$.
-Let $B_{r}^{\sigma}:=\{z\in{\mathcal{Y}}~{}|~{}\sigma(z)<r\}$ be the $\sigma$-ball of radius $r$.
-Then the above discussion implies that
-(3.2) $B_{t}^{\sigma}\subset{\mathbb{D}}_{t}~{}~{}~{}~{}(t\geq 1)~{}~{}~{}\mbox{and}~{}~{}~{}~{}~{}{\mathbb{D}}_{t}\backslash\left(\bigcup_{j=1}^{\infty}{\mathbb{D}}(a_{j},{3\over 2}r_{j})\right)\subset B_{t}^{\sigma}~{}~{}~{}~{}(t\geq{3\over 2}).$
-We now show that the smooth 2-form $dd^{c}\sigma$ defines an order 0 distribution of finite mass on ${\mathcal{Y}}$ (see Lemma 3.2 in [2]) so that Parabolic Logarithmic Derivative Lemma (Lemma 9) holds for $\sigma$.
-We use the formula that, in polar coordinates $z=a+re^{i\theta}$ for some $a\in{\mathbb{C}}$, for a smooth function $\phi$, one has
+Thus $\tau(z)=H(|z|)$ and $\hat{\tau}(z)=\log^{+}|z|$, which are equal when $\tau(z)\geq\log{3\over 2}$.
+Hence $\sigma=\hat{\sigma}$ when $\sigma\geq{3\over 2}$.
+We say that such $\sigma$ is associated to the discs $\{{\mathbb{D}}(a_{j},2r_{j})\}$ by assuming that $\{r_{j}\}$ satisfy (i) and (ii) above.
+Note that $\sigma$ is not necessarily a parabolic exhaustion function because $dd^{c}\log\sigma$ may not be compactly supported.
+Denote by $B^{\sigma}_{r}:=\{y\in{\mathcal{Y}}:\sigma(y)<r\}$ and by $S^{\sigma}_{r}:=\{y\in{\mathcal{Y}}:\sigma(y)=r\}$.
+Denote by $\mu_{r}$ the measure induced by the differential $d^{c}\log\sigma|_{S^{\sigma}_{r}}$ and write $d\mu_{r}=d^{c}\log\sigma|_{S^{\sigma}_{r}}$.
+Note that, from the above discussion (see also [2]),
+(3.5) $B_{t}^{\sigma}\subset{\mathbb{D}}_{t}~~~~\mbox{for}~t\geq 1,~~~\mbox{and}~~~~~{\mathbb{D}}_{t}\backslash\left(\bigcup_{j=1}^{\infty}{\mathbb{D}}(a_{j},{3\over 2}r_{j})\right)\subset B_{t}^{\sigma}~~~~\mbox{for}~t\geq{3\over 2}.$
+Lemma 8 (Lemma 3.2 in [2]).
+The smooth 2-form $dd^{c}\log\sigma$ defines an order 0 distribution of finite mass on $\mathcal{Y}$}.
+Proof.
+We use the formula: in polar coordinates $z=a+re^{i\theta}$, for some $a\in{\mathbb{C}}$, and for a smooth function $\phi$, one has
 $d^{c}\phi={1\over 4\pi}\left(r{\partial\phi\over\partial r}d\theta-{1\over r}{\partial\phi\over\partial\theta}dr\right).$
-The smooth 2-form $dd^{c}\sigma$ is supported on $U$ since $\log\sigma$ is harmonic elsewhere.
-On the annulus $A(0,{1\over 2},{3\over 2})$, in polar coordinates $z=re^{i\theta}$ one has (see the Appendix in [2])
-$dd^{c}H(|z|)=d\left({r\over 4\pi}{\partial H(r)\over\partial r}d\theta\right)=d(h(r-c)d\theta)=O(1)dr\wedge d\theta=O(1)rdr\wedge d\theta$
-since ${1\over 2}\leq r\leq{3\over 2}$ is bounded.
-Thus $dd^{c}H(|z|)$ is of finite mass on $A(0,{1\over 2},{3\over 2})\backslash{\mathcal{E}}$.
-On the annulus $A(a_{j},{1\over 2}r_{j},{3\over 2}r_{j})$,
-$dd^{c}\left(r_{j}H\left({r_{j}\over|z-a_{j}|}\right)\right)$ $=d\left({r\over 4\pi}{\partial r_{j}H\left({r_{j}\over r}\right)\over\partial r}d\theta\right)$
-$=d\left(-r_{j}h\left({r_{j}\over r}-c\right)d\theta\right)=O(r_{j})dr\wedge d\theta=O(1)rdr\wedge d\theta.$
-Since $\sum_{j=1}^{\infty}r^{2}_{j}\leq\sum_{j=1}^{\infty}r_{j}<+\infty$, the support $U$ is of finite Lebesgue measure. We conclude that $dd^{c}\log\sigma$ is of finite mass.
-Using the above computations with ${\mathcal{E}}:=f^{-1}(D)$, we recall the following important result due to Z. Chen, D.T. Huynh, R. Sun and S.Y. Xie [2].
+The smooth 2-form $dd^{c}\log\sigma$ is supported on $U$, where $U$ is given in (3.4).
+On the annulus $A(0,{1\over 2},{3\over 2})$, in polar coordinate $z=re^{i\theta}$, by (3.2) one has (see [2, Appendix])
+(3.6) $dd^{c}\tau=dd^{c}H(|z|)=d\left({r\over 4\pi}{\partial H(r)\over\partial r}d\theta\right)=d(h(r-c)d\theta)=h^{\prime}\left(r-c\right)dr\wedge d\theta=O(1)dr\wedge d\theta.$
+Thus $dd^{c}\tau$ has finite mass on $A(0,{1\over 2},{3\over 2})$.
+On the annulus $A(a_{j},{1\over 2}r_{j},{3\over 2}r_{j})$, since ${\mathbb{D}}(a_{j},2r_{j})$ for all $j$ and the disc ${\mathbb{D}}_{2}$ are disjoint, we have similarly that $\tau=r_{j}H\left({r_{j}\over|z-a_{j}|}\right)$ on $A(a_{j},{1\over 2}r_{j},{3\over 2}r_{j})$.
+Therefore,
+$dd^{c}\tau$ $=dd^{c}\left(r_{j}H\left({r_{j}\over|z-a_{j}|}\right)\right)$
+(3.7) $=d\left({r\over 4\pi}{\partial r_{j}H\left({r_{j}\over r}\right)\over\partial r}d\theta\right)=\left({r_{j}\over r}\right)^{2}h^{\prime}\left({r_{j}\over r}-c\right)dr\wedge d\theta=O(1)dr\wedge d\theta.$
+Since $\sum_{j=1}^{\infty}r_{j}<+\infty$, the support $U$ is of finite Lebesgue measure, this proves the lemma. ∎
+We now derive the parabolic Jensen’s formula in our setting.
+For any function $\upsilon:{\mathcal{Y}}\rightarrow(-\infty,+\infty)$ such that $dd^{c}\upsilon$ is of order 0, the parabolic Jensen’s formula (see [16, Remark 3.2, p.
+22]) is stated as
+(3.8) $\int_{1}^{r}{dt\over t}\int_{B_{t}^{\sigma}}dd^{c}\upsilon=\int_{S_{r}^{\sigma}}\upsilon d\mu_{r}-\int_{B_{r}^{\sigma}}\upsilon dd^{c}\tau\quad(r>1).$
+Let $g$ be meromorphic on $\mathbb{C}$}. In (3.8), by letting $\upsilon:=\log|g|$, and also using the Poincaré-Lelong formula, we get the following lemma.
+Lemma 9 (Parabolic Jensen’s formula).
+Let $g$ be a meromorphic function on $\mathbb{C}$}. Then
+$\int_{S_{r}^{\sigma}}\log|g|d\mu_{r}=N_{g,\sigma}(0,r)-N_{g,\sigma}(\infty,r)+c_{g,\sigma}(r),$
+where $N_{g,\sigma}(0,r):=\int_{1}^{r}\sum_{z\in B_{t}^{\sigma}}\max\{0,{\rm ord}_{z}g\}{dt\over t},$ $N_{g,\sigma}(\infty,r):=N_{\frac{1}{g},\sigma}(0,r)$, and
+(3.9) $c_{g,\sigma}(r):=\int_{B_{r}^{\sigma}}\log|g|dd^{c}\tau.$
+Denote by
+(3.10) $c^{+}_{g,\sigma}(r):=\int_{B_{r}^{\sigma}}\log^{+}|g|dd^{c}\tau.$
+Lemma 10.
+Let $g$ be a meromorphic function on $\mathbb{C}$}.
+Let $\mathcal{Y}$$\subset{\mathbb{C}}$ be given with $\sigma$ being associated to the discs $\{{\mathbb{D}}(a_{j},2r_{j})\}$.
+Then we can choose $r_{j}$ small enough such that there exists a positive constant $C_{g}$ depending on $g$ with $c^{+}_{g,\sigma}(r)\leq C_{g}$ for any $r>0$.
+Proof.
+By using (3.5) for $t\geq 2$, and notice that $dd^{c}\tau$ is supported on $U$ where $U$ is given in (3.4), we have
+[…]
+(3.12) $=\int_{A(0,{1\over 2},{3\over 2})}\log^{+}|g|dd^{c}\tau+\int_{\cup_{j\geq 1}A(a_{j},{1\over 2}r_{j},{3\over 2}r_{j})\cap{\mathbb{D}}_{t}}\log^{+}|g|dd^{c}\tau.$
+Notice that $\log^{+}|g|$ is locally integrable around the zeros or poles of $g$ and by (3.6), we get
+$\int_{A(0,{1\over 2},{3\over 2})}\log^{+}|g|dd^{c}\tau<M_{1},$
+where $M_{1}$ is a positive constant depending on $g$.
+We now estimate the integration over $A(a_{j},{1\over 2}r_{j},{3\over 2}r_{j})$.
+If $g(a_{j})\not$$\infty$.
+Then $\log^{+}|g|\leq C_{1}$ for some $C_{1}>0$ on ${\mathbb{D}}(a_{j},{3\over 2}r_{j})$.
+Then, by (3.1),
+$\int_{A(a_{j},{1\over 2}r_{j},{3\over 2}r_{j})}\log^{+}|g|dd^{c}\tau\leq C_{1}\cdot\frac{9\pi}{2}C_{h}r_{j}<{1\over 2^{j}},$
+by further shrinking $r_{j}$, where $C_{h}:=\max_{\frac{1}{2}\leq s\leq\frac{3}{2}}|h^{\prime}(s-c)|.$ If $g(a_{j})=\infty$.
+Then, locally, $g(z)=(z-a_{j})^{k_{j}}g_{j}(z)$ with $g_{j}(a_{j})\not=0$.
+Thus $\log^{+}|g|\leq C_{2}\log^{+}{1\over|z-a_{j}|}$ for some $C_{2}>0$ on ${\mathbb{D}}(a_{j},{3\over 2}r_{j})$.
+We have similarly
+$\int_{A(a_{j},{1\over 2}r_{j},{3\over 2}r_{j})}\log^{+}|g|dd^{c}\tau\leq\frac{9\pi}{2}C_{2}C_{h}\int_{{1\over 2}r_{j}}^{{3\over 2}r_{j}}\log rdr\leq\frac{9\pi}{2}C_{2}C_{h}r_{j}^{\delta},$
+for some $1<\delta<2$. Again, we can choose $r_{j}$ sufficiently small such that
+$\int_{A(a_{j},{1\over 2}r_{j},{3\over 2}r_{j})}\log^{+}|g|dd^{c}\tau<{1\over 2^{j}}.$
+Therefore, with this choice of the radii $r_{j}$, we obtain from (3.11) that $c^{+}_{g,\sigma}(r)\leq M_{1}+\sum_{j=1}^{\infty}{1\over 2^{j}}=M_{1}+1$. ∎
+Note that the shrinking process of $r_{j}$ in the above lemma depends on the given function $g$.
+So we must ensure in our arguments that such shrinking occurs only finitely many times.
+Throughout the rest of the paper, we will keep shrinking the radii $r_{j}$ whenever necessary.
+The $\sigma$ should be understood as being adjusted accordingly, although we will continue to denote it by the same symbol.
+3.2. The characteristic functions in the parabolic setting
+We continue to let $\{\mathcal{Y}$$\mathbb{C}$$\backslash\mathcal{E}$ with $\sigma$ being the exhaustion associated to the discs $\{{\mathbb{D}}(a_{j},2r_{j})\}$, constructed in Section 3.1.
+Let $f:{\mathbb{C}}\rightarrow{\mathbb{P}}^{n}({\mathbb{C}})$ be a holomorphic map.
+We briefly recall some standard notations in Nevanlinna theory in the parabolic context (for details, see [2], [16]).
+Let $D=\{Q=0\}$ be a hypersurface in $\mathbb{P}^{n}({\mathbb{C}})$, where $Q\in\mathbb{C}[x_{0},\ldots,x_{n}]$ is a homogeneous polynomial of degree $d$.
+  1. 1.
+    The proximity function of $f$ with respect to $D$ is defined as
+    $m_{f,\sigma}(D,r)=\int_{S_{r}^{\sigma}}\lambda_{D}(f)d\mu_{r},$
+    where $\lambda_{D}$ is the Weil-function on ${\mathbb{P}}^{n}({\mathbb{C}})$ given by
+    $\lambda_{D}({\bf x})=\log{\|{\bf x}\|^{d}\over|Q({\bf x})|},~~{\bf x}=[x_{0}:\cdots:x_{n}]\in\mathbb{P}^{n}({\mathbb{C}}),~~~\|{\bf x}\|:=\max\{|x_{0}|,\dots,|x_{n}|\}.$
+  2. 2.
+    The Nevanlinna characteristic (height) function is defined by
+    $T_{f,\sigma}(r):=\int_{S_{r}^{\sigma}}\log\|{\bf f}\|d\mu_{r},~~~~r>1,$
+    where ${\bf f}=(f_{0},\ldots,f_{n})$ is a reduced representation of $f$. The Ahlfors-Shimizu characteristic (height) function is defined by
+    $T^{A}_{f,\sigma}(r)=\int_{1}^{r}{dt\over t}\int_{B_{t}^{\sigma}}f^{*}\omega_{FS},~~~~r>1,$
+    where $\omega_{FS}$ denotes the Fubini-Study form.
+Note that $T_{f,\sigma}(r)$ depends on the choices of the representation $\bf f$} of $f$, while $T^{A}_{f,\sigma}(r)$ does not depend on the representation.
+Therefore, when we write $T_{f,\sigma}(r)$, we always assume that we have fixed a reduced a representation ${\bf f}=(f_{0},\dots,f_{n})$ of $f$.
+  3. 3.
+    The counting function and the $k$-truncated counting function are defined by
+    $N_{f,\sigma}(D,r):=\int_{1}^{r}\sum_{z\in B_{t}^{\sigma}}{\rm ord}_{z}f^{*}D{dt\over t},~~~~~~~\mbox{and}~~N_{f,\sigma}^{(k)}(D,r):=\int_{1}^{r}\sum_{z\in B_{t}^{\sigma}}\mathop{\mathrm{min}}\{k,{\rm ord}_{z}f^{*}D\}{dt\over t}.$
+  4. 4.
+    The counting function extends naturally to the sub-schemes, as well as the notion $N_{\text{gcd},\sigma}(f,g)(r)$, where $f,g$ are meromorphic functions on $\mathcal{Y}$. See [13].
+We also use the standard notation $T_{f}(r),m_{f}(D,r),n_{f}(D,t)$ and $N_{f}(D,r)$ of the Nevanlinna functions associated to a holomorphic map $f:{\mathbb{C}}\rightarrow{\mathbb{P}}^{n}({\mathbb{C}})$.
+In this case, the Ahlfors-Shimizu and the classical Nevanlinna characteristic function agree up to an additive constant (see [14, Theorem 1.1.19]), i.e., $T_{f}(r)-T^{A}_{f}(r)=O(1).$
+Theorem 11 (The parabolic First Main Theorem).
+Let $\mathcal{Y}$$\subset{\mathbb{C}}$ be as above.
+Let $D=\{Q=0\}$ be a hypersurface in $\mathbb{P}^{n}({\mathbb{C}})$, where $Q\in\mathbb{C}[x_{0},\ldots,x_{n}]$ is a homogeneous polynomial of degree $d$.
+Let $f=(f_{0},\ldots,f_{n}):{\mathbb{C}}\rightarrow{\mathbb{P}}^{n}({\mathbb{C}})$ be a holomorphic curve, where $f_{0},\ldots,f_{n}$ are entire functions without common zero.
+Suppose that ${f}({\mathcal{Y}})\not\subset\mbox{Supp}(D)$.
+Then
+$dT_{f,\sigma}(r)=m_{f,\sigma}(D,r)+N_{f,\sigma}(D,r)+c_{Q({\bf f}),\sigma}(r).$
+Proof.
+Following the definition of Nevanlinna characteristic function, we have
+$dT_{f,\sigma}(r)=\int_{S_{r}^{\sigma}}\log\|{\bf f}\|^{d}d\mu_{r}=\int_{S_{r}^{\sigma}}\log{\|{\bf f}\|^{d}\over|Q({\bf f})|}d\mu_{r}+\int_{S_{r}^{\sigma}}\log{|Q({\bf f})|}d\mu_{r}.$
+Hence the assertion follows from Lemma 9 together with the definition of the proximity function. ∎
+In the case of a meromorphic function $g$ on $\mathcal{Y}$}, write $g=g_{0}/g_{1}$, where $g_{0}$ and $g_{1}$ are entire functions without common zeros.
+We regard the function $g$ as the map $g:=[g_{0}:g_{1}]:{\mathcal{Y}}\rightarrow{\mathbb{P}}^{1}({\mathbb{C}})$.
+By taking $Q(z_{0},z_{1})=z_{1}$ (so $D=\{z_{1}=0\}=\infty$), we get
+$m_{g,\sigma}(r):=m_{g,\sigma}(\infty,r)=\int_{S_{r}^{\sigma}}\log{\max\{|g_{0}|,|g_{1}|\}\over|g_{1}|}d\mu_{r}=\int_{S_{r}^{\sigma}}\log^{+}|g|d\mu_{r}+O(1).$
+The above $m_{g,\sigma}(r)$ is well-defined, i.e., it is independent of the choice of $g_{0}$ and $g_{1}$.
+However, the Nevanlinna height function $T_{g,\sigma}(r)=\int_{S_{r}^{\sigma}}\log\max\{|g_{0}|,|g_{1}|\}d\mu_{r}$ is not well-defined.
+Instead we introduce a new height function that is well-defined:
+(3.13) $T^{M}_{g,\sigma}(r)=m_{g,\sigma}(\infty,r)+N_{g,\sigma}(\infty,r).$
+From the parabolic Jenen’s formula above,
+(3.14) $T^{M}_{1/g,\sigma}(r)=T^{M}_{g,\sigma}(r)+c_{1/g,\sigma}(r).$
+We now compare $T^{M}_{g,\sigma}(r)$ with $T^{A}_{g,\sigma}(r)$. From parabolic Jensen’s formula,
+$T^{A}_{g,\sigma}(r)$ $=$ $\int_{1}^{r}{dt\over t}\int_{B_{t}^{\sigma}}dd^{c}\log\max\{|g_{0}|,|g_{1}|\}=\int_{S_{r}^{\sigma}}\log\max\{|g_{0}|,|g_{1}|\}d\mu_{r}$
+$-\int_{B_{r}^{\sigma}}\log\max\{|g_{0}|,|g_{1}|\}dd^{c}\tau$
+$=$ $\int_{S_{r}^{\sigma}}\log^{+}|g|d\mu_{r}+\int_{S_{r}^{\sigma}}\log|g_{1}|d\mu_{r}-\int_{B_{r}^{\sigma}}\log\max\{|g_{0}|,|g_{1}|\}dd^{c}\tau$
+$=$ $T^{M}_{g,\sigma}(r)+{1\over 2}\int_{B_{r}^{\sigma}}\log\left(1/(1+|g|^{2})\right)dd^{c}\tau=T^{M}_{g,\sigma}(r)+{1\over 2}c_{1/(1+|g|^{2}),\sigma}(r).$
+Similarly to the proof of Lemma 10, by choosing $r_{j}$ small enough, there exists a constant $C>0$ with $c^{+}_{1/(1+|g|^{2}),\sigma}(r)\leq C$. Hence
+(3.15) $|T^{M}_{g,\sigma}(r)-T^{A}_{g,\sigma}(r)|\leq C.$
+It is important to note that these radii $r_{j}$ depend on the function $g$.
 Proposition 12 ([2]).
-Let $f:{\mathbb{C}}\rightarrow{\mathbb{P}}^{n}({\mathbb{C}})$ be a holomorphic map.
-Let $D$ be a hypersurface in ${\mathbb{P}}^{n}({\mathbb{C}})$.
-Let ${\mathcal{Y}}_{f}:={\mathbb{C}}\backslash f^{-1}(D)$ with the smooth exhaustion function $\tau:=\log\sigma$ given above.
+Let $f=(f_{0},\ldots,f_{n}):{\mathbb{C}}\rightarrow{\mathbb{P}}^{n}({\mathbb{C}})$ be a holomorphic curve, where $f_{0},\ldots,f_{n}$ are entire functions without common zero.
+Let $\mathcal{Y}$$\subset{\mathbb{C}}$ be given with $\sigma$ being associated to the discs $\{{\mathbb{D}}(a_{j},2r_{j})\}$.
+Then we can choose $r_{j}$ small enough such that, for all $r\geq 1$, the following hold:
+  1. (i)
+    $O(1)\leq T_{f}(r)-T^{A}_{f,\sigma}(r)\leq\log r+O(1).$
+  2. (ii)
+    $O(1)\leq T_{f}(r)-T_{f,\sigma}(r)\leq\log r+O(1).$
+Proof.
+From (3.5), for $t\geq{3\over 2}$, $B_{t}^{\sigma}\subset{\mathbb{D}}_{t}$ and
+${\mathbb{D}}_{t}\backslash\left(\bigcup_{j=1}^{\infty}D(a_{j},{3\over 2}r_{j})\right)\subset B_{t}^{\sigma}.$
+Thus
+$0\leq T_{f}^{A}(r)-T^{A}_{f,\sigma}(r)\leq\int_{{3\over 2}}^{r}{dt\over t}\int_{\left(\bigcup_{j=1}^{\infty}{\mathbb{D}}(a_{j},{3\over 2}r_{j})\right)\cap{\mathbb{D}}_{t}}f^{*}\omega_{FS}+O(1)$
+$\leq\int_{1}^{r}{dt\over t}\int_{\bigcup_{j=1}^{\infty}{\mathbb{D}}(a_{j},{3\over 2}r_{j})}f^{*}\omega_{FS}+O(1)$
+$\leq\int_{1}^{r}{dt\over t}\left(\sum_{j=1}^{\infty}\int_{{\mathbb{D}}(a_{j},{3\over 2}r_{j})}f^{*}\omega_{FS}\right)+O(1),~~~~(r\geq 1).$
+We choose $r_{j}>0$ sufficiently small so that
+(3.16) $\int_{{\mathbb{D}}(a_{j},{3\over 2}r_{j})}f^{*}\omega_{FS}<2^{-j}.$
+Hence the above estimate yields
+$0\leq T_{f}^{A}(r)-T^{A}_{f,\sigma}(r)\leq\log r+O(1),~~~~(r\geq 1).$
+By using the fact that $T_{f}(r)=T_{f}^{A}(r)+O(1)$, this completes the proof of (i).
+To prove (ii), by using (3.8) with $\upsilon:=\log\|{\bf f}\|$, we get
+$T^{A}_{f,\sigma}(r)=\int_{1}^{r}{dt\over t}\int_{B_{t}^{\sigma}}dd^{c}\log\|{\bf f}\|=\int_{S_{r}^{\sigma}}\log\|{\bf f}\|d\mu_{r}-\int_{B_{r}^{\sigma}}\log\|{\bf f}\|dd^{c}\tau=T_{f,\sigma}(r)-\int_{B_{r}^{\sigma}}\log\|{\bf f}\|dd^{c}\tau.$
+Thus
+$|T^{A}_{f,\sigma}(r)-T_{f,\sigma}(r)|\leq\int_{B_{r}^{\sigma}}\log^{+}\|{\bf f}\|dd^{c}\tau+\int_{B_{r}^{\sigma}}\log^{+}(1/\|{\bf f}\|)dd^{c}\tau.$
+Since $f_{0},\ldots,f_{n}$ have no common zero, we can argue similarly to the proof of Lemma 10 to choose $r_{j}$ small enough so that there exists a constant $C>0$ with
+$\int_{B_{r}^{\sigma}}\log^{+}\|{\bf f}\|dd^{c}\tau+\int_{B_{r}^{\sigma}}\log^{+}\frac{1}{\|{\bf f}\|}dd^{c}\tau\leq C.$
+for any $r>0$. By combining above with (i), we prove (ii). ∎
+Again, these radii $r_{j}$ in Proposition 12 depends on the map $f$.
+3.3. The parabolic Logarithmic Derivative Lemma and H. Cartan’s theorem
+To establish the parabolic Logarithmic Derivative Lemma, we introduce the weighted Euler characteristic function (see [16, Definition 3.4]):
+$\mathfrak{X}^{+}_{\sigma}(r):=\int_{S^{\sigma}_{r}}\log^{+}\left|d\sigma\left({\partial\over\partial z}\right)\right|^{2}d\mu_{r}.$
+Using the properties that the disks ${\mathbb{D}}(a_{j},2r_{j})$ are disjoint and that $\sum_{j\geq 1}r_{j}<+\infty$, we see from [16, Example (2), pp. 32 and 33] (see also [2, (2.10)]) that
+$\mathfrak{X}^{+}_{\sigma}(r)=\mathfrak{X}_{\sigma}(r)+O(\log r)=N_{\mathcal{E}}(r)+O(\log r),\quad r>1,$
+where $N_{\mathcal{E}}(r):=\int_{1}^{r}n_{\mathcal{E}}(t){dt\over t}$, and $n_{\mathcal{E}}(t)$ counts the number of points in ${\mathcal{E}}\cap\mathbb{D}_{t}$ for $t\geq 1$.
+In particular, if ${\mathcal{E}}:=f^{-1}(D)$ for a holomorphic map $f:{\mathbb{C}}\rightarrow{\mathbb{P}}^{n}({\mathbb{C}})$ and a divisor $D$ on ${\mathbb{P}}^{n}({\mathbb{C}})$ with $f({\mathbb{C}})\not\subset\mbox{Supp}D$, it follows (see also [2, (4.2)]) that
+(3.17) $\mathfrak{X}^{+}_{\sigma}(r)\leq N_{f}^{(1)}(D,r)+O(\log r),\quad r>1.$
+The above estimate is crucial to our paper.
+In [2], the authors consider the case when $N_{f}(D,r)=o(T_{f}(r))$, and therefore $\mathfrak{X}^{+}_{\sigma}(r)$ is a small function with respect to $T_{f}(r)$.
+In our case, we treat the situation where $N_{f}^{(1)}(D,r)\leq{1\over\ell}T_{f}(r)$ for a sufficiently large $\ell$.
+In this setting, $\mathfrak{X}^{+}_{\sigma}(r)$ should be regarded as having the same order of growth as $T_{f}(r)$.
+Lemma 13 (Parabolic Logarithmic Derivative Lemma).
+Let $g$ be a meromorphic function on $\mathbb{C}$. Let $\mathcal{Y}$$\subset{\mathbb{C}}$ be as above. For any $\delta>0$, one has
+  1. (i)
+    $m_{g^{\prime}/g,\sigma}(r)\leq_{\rm exc}{1\over 2}((1+\delta)^{2}\log T_{g}(r)+\delta\log r+\mathfrak{X}^{+}_{\sigma}(r)).$
+  2. (ii)
+    $m_{g^{(l)}/g,\sigma}(r)\leq_{\rm exc}{l\over 2}\left((1+\delta)^{2}\log T_{g}(r)+\delta\log r+\mathfrak{X}^{+}_{\sigma}(r)\right).$
+Proof.
+(i) We follow the the approach in [16, Theorem 3.8, pp. 26-28]. Consider the auxiliary form
+$\Omega:={1\over|w|^{2}(1+\log^{2}|w|)}{\sqrt{-1}\over 2\pi}dw\wedge d\bar{w},$
+introduced in [16, (15), p. 27], which has finite volume on $\mathbb{C}$}. By the change of variable formula, we have
+$\int_{1}^{r}{dt\over t}\int_{B_{t}^{\sigma}}g^{*}\Omega=\int_{w\in{\mathbb{C}}}N_{g,\sigma}(w,r)\Omega.$
+Since $B_{t}^{\sigma}\subset\mathbb{D}_{t}$ for any $t\geq 1$, and by the classical First Main Theorem,
+$N_{g,\sigma}(w,r)\leq N_{g}(w,r)\leq T_{g}(r)+O(1),~~~\mbox{for}~r>1.$
+Thus
+$\int_{1}^{r}{dt\over t}\int_{B_{t}^{\sigma}}g^{*}\Omega=\int_{w\in{\mathbb{C}}}N_{g,\sigma}(w,r)\Omega\leq\int_{\mathbb{C}}T_{g}(r)\Omega+\int_{\mathbb{C}}C\Omega.$
+Consequently, we obtain
+$\int_{1}^{r}{dt\over t}\int_{B_{t}^{\sigma}}g^{*}\Omega\leq C_{0}T_{g}(r)$
+for some constant $C_{0}>0$. The rest of the proof follows from [16, Theorem 3.8, p. 28]. This finishes the proof of (i).
+To prove (ii), notice that
+${g^{(l)}\over g}={g^{(l)}\over g^{(l-1)}}{g^{(l-1)}\over g^{(l-2)}}\cdots{g^{\prime}\over g}.$
+Therefore, by (i), we obtain
+$m_{g^{(l)}/g,\sigma}(r)\leq\sum_{j=1}^{l}m_{g^{(j)}/g^{(j-1)},\sigma}(r)\leq_{\rm exc}{1\over 2}\left((1+\delta)^{2}\sum_{j=1}^{l}\log T_{g^{(j)}}(r)+\delta l\log r+l\mathfrak{X}^{+}_{\sigma}(r)\right).$
+By the classical First Main Theorem,
+$T_{g^{(j)}}(r)=m_{g^{(j)}}(r)+N_{g^{(j)}}(r)\leq m_{g^{(j)}/g^{(j-1)}}(r)+m_{g^{(j-1)}}(r)+2N_{g^{(j-1)}}(r)\leq_{\rm exc}3T_{g^{(j-1)}}(r)+O(1).$
+Hence
+$T_{g^{(j)}}(r)\leq_{\rm exc}3^{j}T_{g}(r)+O(1).$
+Therefore, we get
+$m_{g^{(l)}/g,\sigma}(r)\leq_{\rm exc}{l\over 2}\left((1+\delta)^{2}\log T_{g}(r)+\delta\log r+\mathfrak{X}^{+}_{\sigma}(r)\right).$
+In particular, by the definition, the above lemma gives the following:
+Lemma 14.
+If $u$ is an entire function such that it has no zeros on $\mathcal{Y}$ ((i.e. $u$ is a unit on $\mathcal{Y}$)), then, for any $\delta>0$,
+$T^{M}_{u^{\prime}/u,\sigma}(r)\leq_{\rm exc}{1\over 2}((1+\delta)^{2}\log T_{u}(r)+\delta\log r+\mathfrak{X}^{+}_{\sigma}(r)).$
+Using the parabolic Logarithmic Derivative Lemma and the parabolic Jensen’s formula, we can easily derive the following parabolic version of H. Cartan’s theorem.
+Theorem 15.
+Let $\{\mathcal{Y}$$\mathbb{C}$$\backslash\mathcal{E}$ with the exhaustion function $\sigma$ constructed above.
+Let $f=(f_{0},\dots,f_{n}):{\mathbb{C}}\rightarrow{\mathbb{P}}^{n}({\mathbb{C}})$ be a holomorphic map whose image is not contained in any proper linear subspace of ${\mathbb{P}}^{n}({\mathbb{C}})$, where $f_{0},\dots,f_{n}$ are entire functions without common zeros.
+Let $H_{1},\dots,H_{q}$ be hyperplanes on $\mathbb{P}^{n}(\mathbb{C})$.
+Then, for $\delta>0$,
+$\int_{S_{r}^{\sigma}}\max_{J}\sum_{j\in J}\lambda_{H_{j}}(f)d\mu_{r}+N_{W,\sigma}(0,r)$
+$\leq_{\rm exc}(n+1)T_{f,\sigma}(r)+O(\log T_{f}(r))+\frac{n(n+1)}{2}(\mathfrak{X}^{+}_{\sigma}(r)+\delta\log r)+c_{\frac{1}{W},\sigma}(r),$
+where $W:=W(f_{0},\dots,f_{n})$ is the Wronskian of $f$ and the maximum is taken over all subsets $J\subset\{1,2,\dots,q\}$ such that $H_{j},j\in J$, are in general position.
+Proof.
+We follow the proof in [17, Theorem A.5.1.3].
+Let $H_{j}=\{{\bf w}\in{\mathbb{P}}^{n}({\mathbb{C}})~|~<{{\bf w},\bf a}_{j}>=0\}$ for $j=1,\dots,q$.
+Let $T$ be the set of all injective maps $\mu:\{0,1,\dots,n\}\rightarrow\{1,\dots,q\}$ such that ${\bf a}_{\mu(0)},\dots,{\bf a}_{\mu(n)}$ are linearly independent.
 Then
-  1. (i)
-    $0\leq T_{f}(r)-T_{\widetilde{f},\mathcal{Y}_{f}}(r)\leq\log r+O(1),$
-    for $r\geq 1$, where $\widetilde{f}$ the restriction of $f$ on $\mathcal{Y}_{f}$.
-  2. (ii)
-    For $r>1$,
-    $\mathfrak{X}^{+}_{\sigma}(r)\leq N^{(1)}_{f}(r,D)+C\log r,$
-    where $C$ is a constant independent of $f$ and $D$.
-We outline a proof here.
+(3.18) $\int_{S_{r}^{\sigma}}\max_{J}\sum_{j\in J}\lambda_{H_{j}}(f)d\mu_{r}$ $\leq$ $\int_{S_{r}^{\sigma}}\max_{\mu\in T}\sum_{j=0}^{n}\log~\left({\|{\bf f}\|\|{\bf a}_{\mu(j)}\|\over|<{\bf f},{\bf a}>|}\right)d\mu_{r}$
+$\leq$ $\int_{S_{r}^{\sigma}}\max_{\mu\in T}\log^{+}{|W(<{\bf f},{\bf a}_{\mu(0)}>,\dots,<{\bf f},{\bf a}_{\mu(n)}>)|\over\prod_{j=0}^{n}|<{\bf f},{\bf a}_{\mu(j)}>|}d\mu_{r}$
+$+$ $(n+1)T_{f,\sigma}(r)+\int_{S_{r}^{\sigma}}\log(1/|W(f_{0},\dots,f_{n})|)d\mu_{r}+O(1).$
+By applying the logarithmic derivative lemma (Lemma 13) and the parabolic Jensen formula (Lemma 9), the theorem follows. ∎
+The following result can be derived from the proof of [19, Theorem 2.1] and Lemma 13.
+Theorem 16.
+Let $f_{0},\dots,f_{n},f_{n+1}$ be entire functions without common zeros.
+Let $\{\mathcal{Y}$$\mathbb{C}$$\backslash\mathcal{E}$ with the exhaustion function $\sigma$ constructed above.
+Assume that $f_{0}+\dots+f_{n}+f_{n+1}=0.$ If $\sum_{i\in I}f_{i}\neq 0$ for any proper subset $I\subset\{0,\dots,n+1\}$, then we have
+$T^{M}_{f_{i}/f_{j},\sigma}(r)\leq_{\rm exc}\sum_{l=0}^{n+1}N_{f_{l},\sigma}^{(n)}(0,r)+\frac{n(n+1)}{2}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+c_{\frac{1}{D_{0}},\sigma}(r)+\sum_{\begin{subarray}{c}l=0\\ l\neq j\end{subarray}}^{n+1}c_{f_{l},\sigma}(r)+O\left(\max_{1\leq l\leq n}\log T_{\frac{f_{l}}{f_{0}}}(r)\right),$
+where $D_{0}$ is an entire function appeared in the proof of [19, Theorem 2.1] or [17, Theorem A5.2.6]), which is determined in terms of $f_{0},\dots,f_{n}$ only.
 Proof.
-Let ${\mathcal{E}}:=f^{-1}(D)$ and write :=${\mathcal{E}}:=\{a_{j}\}_{j=1}^{\infty}$.
-Denote by $B_{r}^{\sigma}$ and $S_{r}^{\sigma}$ the $\sigma$-ball of radius $r$ and its boundary.
-From (3.2), $B_{t}^{\sigma}\subset{\mathbb{D}}_{t}$ ($t\geq 1$) and
-${\mathbb{D}}_{t}\backslash\left(\bigcup_{j=1}^{\infty}D(a_{j},{3\over 2}r_{j})\right)\subset B_{t}^{\sigma}~{}~{}~{}~{}~{}~{}~{}~{}(t\geq{3\over 2}).$
+We first sketch the proof by assuming that $f_{0},\dots,f_{n}$ are linearly independent. By applying Theorem 15 to $f=(f_{0},\dots,f_{n+1})$, we have
+$\sum_{j=1}^{n+2}m_{f,\sigma}(H_{j},r)+N_{W,\sigma}(r)\leq(n+1)T_{f,\sigma}(r)+\frac{n(n+1)}{2}(\mathfrak{X}^{+}_{\sigma}(r)+\delta\log r)+c_{\frac{1}{W},\sigma}(r)+O(\log T_{f}(r)),$
+where $H_{i}=\{x_{i-1}=0\},i=1,\dots,n+1$ and $H_{n+2}=\{x_{0}+\cdots+x_{n}=0\}$. By Theorem 11, for $1\leq j\leq n+2$,
+$T_{f,\sigma}(r)=m_{f,\sigma}(H_{j},r)+N_{f,\sigma}(H_{j},r)+c_{f_{j-1},\sigma}(r).$
+Thus we get,
+$T_{f,\sigma}(r)\leq\sum_{j=0}^{n+1}N_{f_{j},\sigma}(0,r)-N_{W,\sigma}(r)+\frac{n(n+1)}{2}(\mathfrak{X}^{+}_{\sigma}(r)+\delta\log r)+c_{\frac{1}{W},\sigma}(r)+\sum_{j=0}^{n+1}c_{f_{j},\sigma}(r)+O(\log T_{f}(r)).$
+It is standard to verify that
+$\sum_{j=0}^{n+1}N_{f_{j},\sigma}(0,r)-N_{W,\sigma}(r)\leq\sum_{j=0}^{n+1}N^{(n)}_{f_{j},\sigma}(0,r).$
+The proof can be completed by applying the following estimate, valid for any for any $0\leq i\leq n$ and $0\leq j\leq n+1$:
+$T_{f,\sigma}(r)$ $=\int_{S_{r}^{\sigma}}\log\max_{0\leq l\leq n+1}|f_{l}(z)|d\mu_{r}=\int_{S_{r}^{\sigma}}\log\max_{0\leq l\leq n+1}(|f_{l}(z)|/|f_{j}(z)|)d\mu_{r}+\int_{S_{r}^{\sigma}}\log|f_{j}(z)|d\mu_{r}$
+$\geq\int_{S_{r}^{\sigma}}\log^{+}\left|\frac{f_{i}(z)}{f_{j}(z)}\right|d\mu_{r}+N_{f_{j},\sigma}(0,r)+c_{f_{j},\sigma}(r)$
+(3.19) $\geq m_{\frac{f_{i}}{f_{j}},\sigma}(r)+N_{\frac{f_{i}}{f_{j}},\sigma}(\infty,r)+c_{f_{j},\sigma}(r)=T^{M}_{f_{i}/f_{j},\sigma}(r)+c_{f_{j},\sigma}(r).$
 Thus
-$0\leq T_{f}(r)-T_{\widetilde{f},\mathcal{Y}_{f}}(r)\leq\int_{{3\over 2}}^{r}{dt\over t}\int_{\left(\bigcup_{j=1}^{\infty}{\mathbb{D}}(a_{j},{3\over 2}r_{j})\right)\cap{\mathbb{D}}_{t}}f^{*}\omega_{FS}+O(1)$
-$\leq\int_{1}^{r}{dt\over t}\int_{\bigcup_{j=1}^{\infty}{\mathbb{D}}(a_{j},{3\over 2}r_{j})}f^{*}\omega_{FS}+O(1)$
-$=\int_{1}^{r}{dt\over t}\left(\sum_{j=1}^{\infty}\int_{{\mathbb{D}}(a_{j},{3\over 2}r_{j})}f^{*}\omega_{FS}\right)+O(1),~{}~{}~{}~{}(r\geq 1).$
-Recall in (3.1) that the radius $r_{j}>0$ can be chosen arbitrarily small.
-For our purpose, for each $j\geq 1$, we choose $r_{j}>0$ sufficiently small so that $\int_{{\mathbb{D}}(a_{j},{3\over 2}r_{j})}f^{*}\omega_{FS}<2^{-j}$.
-Hence the above estimate yields
-$0\leq T_{f}(r)-T_{\widetilde{f},\mathcal{Y}_{f}}(r)\leq\log r+O(1),~{}~{}~{}~{}(r\geq 1).$
-This proves (i).
-To prove (ii), using the above computations and following the argument in [16, Proposition 3.3, pp. 32-33] (see also [2, Line 3 on page 12]), the weighted Euler characteristic satisfies
-$\mathfrak{X}^{+}_{\sigma}(r)=\mathfrak{X}_{\sigma}(r)=\int_{1}^{r}\#\{j:|a_{j}|<t\}{dt\over t}+O(\log r).$
-Note that #$\#\{{\mathbb{D}}_{t}\cap{\mathcal{E}}\}$ is exactly $n_{f}^{(1)}(D,t)$, which is finite. Since $B^{\sigma}_{r}\subset{\mathbb{D}}_{r}$ for $r\geq 1$,
-$\mathfrak{X}^{+}_{\sigma}(r)\leq N^{(1)}_{f}(r,D)+C\log r.$
-This completes the proof of (ii). ∎
-3.2. Nevanlinna theory on parabolic Riemann surfaces
-Let ${\mathcal{Y}}$ be a parabolic open Riemann surface. The following is the starting point of the parabolic Nevanlinna theory.
-Lemma 9 (Parabolic Logarithmic Derivative Lemma, Theorem 3.7 in [16] and Proposition 5.1 in [10]).
-Let $f$ be a meromorphic function on ${\mathcal{Y}}$. For any $\delta>0$, one has
-$m_{f^{\prime}/f,{\mathcal{Y}}}(r)\leq_{\rm exc}{1\over 2}((1+\delta)^{2}\log T_{f,{\mathcal{Y}}}(r)+\delta\log r+\mathfrak{X}^{+}_{\sigma}(r))+O(1).$
-In particular, it gives the following
-Lemma 10.
-If $u$ is a nowhere vanishing holomorphic function on $\mathcal{Y}$ ((((i.e. $u$ is a unit)))), then, for any $\delta>0$, one has
-$T_{u^{\prime}/u,{\mathcal{Y}}}(r)\leq_{\rm exc}{1\over 2}((1+\delta)^{2}\log T_{u,{\mathcal{Y}}}(r)+\delta\log r+\mathfrak{X}^{+}_{\sigma}(r))+O(1).$
-We recall the following theorem which is an extension of H. Cartan’s theorem. It was from Theorem 6.3 in [10], although the statement is slightly different.
-Theorem 11 (Theorem 6.3 in [10]).
-Let $f=(f_{0},\dots,f_{n}):{\mathcal{Y}}\rightarrow{\mathbb{P}}^{n}({\mathbb{C}})$ be a holomorphic map whose image is not contained in any proper linear subspaces of ${\mathbb{P}}^{n}({\mathbb{C}})$, where $f_{0},\dots,f_{n}$ are holomorphic functions on ${\mathcal{Y}}$ without common zeros.
-Let $H_{1},\dots,H_{q}$ be hyperplanes on $\mathbb{P}^{n}(\mathbb{C})$.
-Then, for $\epsilon>0$,
-$\int_{S(r)}\max_{J}\sum_{k\in J}$ $\lambda_{H_{k}}(f)d\mu_{r}+N_{W,\mathcal{Y}}(0,r)$
-$\leq_{\rm exc}(n+1+\epsilon)T_{f,{\mathcal{Y}}}(r)+\left(\frac{n(n+1)}{2}+2\epsilon\right)(\mathfrak{X}^{+}_{\sigma}(r)+\delta\log r)+O(1),$
-where $W$ is the Wronskian of $f$, and the maximum is taken over all subsets $J$ such that $H_{j},j\in J$, are in general position.
-The following theorem can be derived from the proof of [18, Theorem 2.1] and Lemma 9.
-Theorem 12.
-Let $f_{0},\dots,f_{n}$ be entire functions on $\mathcal{Y}$ without common zeros.
-Assume that $f_{n+1}$ is the holomorphic function on ${\mathcal{Y}}$ such that $f_{0}+\dots+f_{n}+f_{n+1}=0$.
-If $\sum_{i\in I}f_{i}\neq 0$ for any proper subset $I\subset\{0,\dots,n+1\}$, then, for any $\epsilon>0$, we have
-$(1-\epsilon)T_{f_{j}/f_{i},{\mathcal{Y}}}(r)\ \leq_{\operatorname{exc}}\sum_{i=0}^{n+1}N_{f_{i},{\mathcal{Y}}}^{(n)}(0,r)+\frac{n(n+1)}{2}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)$
-for any pair $0\leq i,j\leq n$.
-3.3. The moving GCD theorem for parabolic Riemann surfaces
+$T^{M}_{f_{i}/f_{j},\sigma}(r)\leq\sum_{j=0}^{n+1}N^{(n)}_{f_{j},\sigma}(0,r)+\frac{n(n+1)}{2}(\mathfrak{X}^{+}_{\sigma}(r)+\delta\log r)+c_{\frac{1}{W},\sigma}(r)+\sum_{\begin{subarray}{c}l=0\\ l\neq j\end{subarray}}^{n+1}c_{f_{l},\sigma}(r)+O\left(\max_{1\leq l\leq n}\log T_{\frac{f_{l}}{f_{0}}}(r)\right).$
+Passing from the condition that $f_{0},\dots,f_{n}$ are linearly independent to the situation in which no proper subsum of $f_{0}+\dots+f_{n+1}$ vanishes involves only algebraic arguments.
+In particular, the resulting statements are obtained from the present ones by changing $W$ to $D_{0}$, the modified Wronskian that appears in the proof of [19, Theorem 2.1].
+∎
+3.4. The moving GCD theorem in the parabolic setting
 One of the key components in [9] is the extension of the GCD theorem in [13] to the case where the coefficients of two coprime polynomials are allowed to be meromorphic functions of the same growth.
-In this subsection, we aim to extend this line of argument to the setting of parabolic Riemann surfaces.
+In this subsection, we aim to extend this line of argument to the parabolic setting.
 Before we proceed, let us briefly outline the development of the GCD theorem.
@@
 More recently, a further generalization was achieved in [10], which applied Vojta’s results on integral points on semiabelian varieties for the number field case, as well as [15] for the complex case.
-To proceed, we will first derive a version of the second main theorem with moving targets of the same growth, which will serve our purpose.
-Let $\mathcal{Y}$ be a parabolic open Riemann surface.
-Let $a_{0},\ldots,a_{n}$ be meromorphic functions, not all zeros, on $\mathcal{Y}$, and let $L:=a_{0}X_{0}+\dots+a_{n}X_{n}$.
+To proceed, we will first derive a version of the second main theorem with moving targets of the same growth, which will serve our purpose. See [18] or [17, Chapter 6].
+Let $a_{0},\ldots,a_{n}$ be meromorphic functions, not all zeros, on $\mathbb{C}$}, and let $L:=a_{0}X_{0}+\dots+a_{n}X_{n}$.
 Then $L$ defines a hyperplane $H$ in $\mathbb{P}^{n}$, over the field generated by $a_{0},\ldots,a_{n}$ over $\mathbb{C}$.
-We note that $H(z)$ is the hyperplane determined by the linear form $L(z)=a_{0}(z)X_{0}+\dots+a_{n}(z)X_{n}$ for $z\in\mathcal{Y}$ where $a_{0},\cdots,a_{n}$ are entire and there is no common zero of $a_{0},\cdots,a_{n}$ at $z$.
+We note that $H(z)$ is the hyperplane determined by the linear form $L(z)$.
 The definition of the Weil function, proximity function and counting function can be easily extended to moving hyperplanes.
@@
 (3.20) $\lambda_{H(z)}(P)=-\log\frac{|a_{0}(z)x_{0}+\cdots+a_{n}(z)x_{n}|}{\max\{|x_{0}|,\dots,|x_{n}|\}\max\{|a_{0}(z)|,\dots,|a_{n}(z)|\}},$
-for $z\in\mathcal{Y}$ which is not a common zero of $a_{0},\cdots,a_{n}$, or a pole of any $a_{k}$, $0\leq k\leq n$. The First Main Theorem for a moving hyperplane $H$ can be stated as
-(3.4) $T_{\mathbf{f},\mathcal{Y}}(r)=N_{\mathbf{f},\mathcal{Y}}(H,r)+m_{\mathbf{f},\mathcal{Y}}(H,r)+T_{\mathbf{a},\mathcal{Y}}(r)+O(1),$
-where ${\mathbf{a}}:=[a_{0}:\cdots:a_{n}]$.
-Without loss of generality, we will normalize the linear forms $L_{j}$, $1\leq j\leq q$, such that for each $1\leq j\leq q$, there exists $0\leq j^{\prime}\leq n$ with $a_{ij^{\prime}}=1$.
+for $z\in\mathcal{Y}$ which is not a common zero of $a_{0},\cdots,a_{n}$, or a pole of any $a_{k}$, $0\leq k\leq n$.
+We now consider a family of linear forms $L_{j}:=\sum_{i=0}^{n}a_{ij}X_{j}$, $1\leq j\leq q$, as the aformentioned linear forms.
+Without loss of generality, we normalize the linear forms $L_{j}$, $1\leq j\leq q$, such that for each $1\leq j\leq q$, there exists $0\leq j^{\prime}\leq n$ with $a_{jj^{\prime}}=1$.
 Let $t$ be a positive integer and let $V(t)$ be the complex vector space spanned by the elements
@@
 (3.21) $\liminf_{t\to\infty}\dim V(t+1)/\dim V(t)=1.$
+See [24, Lemma 6] for a proof.
 Definition 17.
 Let $V$ be a $\mathbb{C}$-vector space spanned by finitely many meromorphic functions.
-We say that a holomorphic map ${\mathbf{f}}=(f_{0},\dots,f_{n}):\mathcal{Y}\to\mathbb{P}^{n}({\mathbb{C}})$ is linearly non-degenerate over $V$ if whenever we have a linear combination $\sum_{i=1}^{m}a_{i}f_{i}=0$ with $a_{i}\in V$, then $a_{i}=0$ for each $i$.
-Otherwise we say that ${\mathbf{f}}$ is linearly degenerate over $V$.
+We say that a holomorphic map $f=(f_{0},\dots,f_{n}):{\mathbb{C}}\to\mathbb{P}^{n}({\mathbb{C}})$ is linearly independent over $V$ if whenever we have a linear combination $\sum_{i=0}^{n}a_{i}f_{i}=0$ with $a_{i}\in V$, then $a_{i}=0$ for each $i$.
+Otherwise we say that $f$ is linearly dependent over $V$.
 The following formulation of the second main theorem with moving targets follows from the strategy in [26] and the proof of [17, Theorem A6.2.4], by adding the Wronskian term and computing the error term explicitly when applying the second main theorem (Theorem 15).
 Theorem 18.
-[…]
-(3.6) $\int_{S(r)}\max_{J}\sum_{k\in J}$ $\lambda_{H_{k}(z)}({\mathbf{f}}(z))d\mu_{r}+\frac{1}{u}N_{W}(0,r)$
-(3.7) $\leq_{\operatorname{exc}}\left(\frac{w}{u}(n+1)+\epsilon\right)T_{\mathbf{f},\mathcal{Y}}(r)+\frac{w}{u}(n+1)(t+2)\max_{1\leq j\leq q}T_{L_{j},{\mathcal{Y}}}(r)$
-(3.8) $+{1\over 2}(n+1)^{2}{w^{2}\over u}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(1),$
-where $T_{L_{j},{\mathcal{Y}}}(r)=T_{[a_{j0}:\cdots:a_{jn}],{\mathcal{Y}}}(r)$, $w=\dim V(t+1)$, $u=\dim V(t)$, and the maximum is taken over all subsets $J$ of $\{1,\dots,q\}$ such that $H_{j}(z)$, $j\in J$, are in general position.
-Using the above theorem, and through the same argument from the proofs of [14, Theorem 5.7], [8, Theorem 26], and [9, Theorem 4.4], we can get the following GCD theorem (note we now consider the case that $u_{1},\ldots,u_{n}$ are units on $\mathcal{Y}$).
-We first fix some notations: Let $K$ be a subfield of the field of meromorphic functions on $\mathcal{Y}$.
+Let $f=(f_{0},\ldots,f_{n}):{\mathbb{C}}\to\mathbb{P}^{n}(\mathbb{C})$ be a holomorphic map where $f_{0},f_{1},\ldots,f_{n}$ are entire functions without common zeros.
+Let $H_{j}$, $1\leq j\leq q$, be arbitrary (moving) hyperplanes given by $L_{j}:=a_{j0}X_{0}+\dots+a_{jn}X_{n}$ where $a_{j0},\cdots,a_{jn}$ are meromorphic functions, not all zeros on $\mathbb{C}$}.
+Let $t$ be a positive integer, $1=b_{1},b_{2},\cdots,b_{u}$ be a basis of $V(t)$ and $b_{1},\cdots,b_{w}$ a basis of $V(t+1)$.
+Denote by $W$ the Wronskian of $\{hb_{m}f_{k}\,|\,1\leq m\leq w,\,0\leq k\leq n\}$, where $h$ is a meromorphic function such that $hb_{1},\ldots,hb_{w}$ are entire functions without common zeros.
+Let $\mathcal{Y}\subset{\mathbb{C}}$ be as above.
+If $f$ is linearly non-degenerate over $V(t+1)$, then, we have the following inequality
+(3.22) $\int_{S^{\sigma}_{r}}\max_{J}\sum_{k\in J}$ $\lambda_{H_{k}(z)}(f(z))d\mu_{r}+\frac{1}{u}N_{W,\sigma}(0,r)$
+(3.23) $\leq_{\operatorname{exc}}\left(\frac{w}{u}(n+1)\right)T_{f,\sigma}(r)+\frac{w}{u}(n+1)(t+2)\sum_{j=1}^{q}T_{L_{j},\sigma}(r)$
+(3.24) $+\frac{w}{2u}(n+1)(nw+w-1)(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+\frac{1}{u}c_{{1\over W},\sigma}(r)+O(\log T_{f}(r)),$
+where $T_{L_{j},\sigma}(r)=\sum_{i=0}^{n}T^{M}_{a_{ji},\sigma}(r)$, $w=\dim V(t+1)$, $u=\dim V(t)$, and the maximum is taken over all subsets $J$ of $\{1,\dots,q\}$ such that $H_{j}(z)$, $j\in J$, are in general position.
+Remark.
+Let $\epsilon$ be a sufficiently small positive number.
+By (3.21), we find $t$ such that $\frac{w}{u}\leq 1+\epsilon$.
+An explicit upper bound for $t$ can be found in the proof of [3, Proposition 5.3] in term of $\epsilon$ and the number of generator for $V(1)$, which is less than $qn$.
+More precisely, if $\epsilon<\frac{1}{16}$, then we can find
+(3.25) $t\leq\frac{3qn}{\epsilon}\log\frac{1}{\epsilon}$
+such that $\frac{w}{u}\leq 1+\epsilon$. Furthermore, for such $t$, we have
+(3.26) $u:=\dim V(t)\leq\left(\frac{1}{\epsilon}\right)^{qn},~~~\mbox{and consequently},~~~~w:=\dim V(t+1)\leq(1+\epsilon)\left(\frac{1}{\epsilon}\right)^{qn}.$
+Using the above theorem, and through the same argument from the proofs of [13, Theorem 5.7], [8, Theorem 26], and [9, Theorem 4.4], we can get the following GCD theorem (note we now consider the case that $u_{1},\ldots,u_{n}$ are units on $\mathcal{Y}$).
+We first fix some notation.
+Let $K$ be a subfield of the field of meromorphic functions on $\mathbb{C}$}.
 Let $F,G$ be coprime polynomials in $n$ variables of the same degree $d>0$ over $K$.
-Assume that one of the coefficients in each expansion of $F$ and $G$ is 1.
+Assume that one of the coefficients in the expansions of $F$ and $G$ is equal to 1.
 For every positive integer $t$, we denote by $V_{F,G}(t)$ the (finite-dimensional) $\mathbb{C}$-vector space spanned by $\prod_{\alpha}\alpha^{n_{\alpha}}$, where $\alpha$ runs through all non-zero coefficients of $F$ and $G$, $n_{\alpha}\geq 0$ and $\sum n_{\alpha}=t$; we also put $d_{t}:=\dim V_{F,G}(t)$.
 For every integer $m\geq d$, we let $M=M_{m,n,d}=2\binom{m+n-d}{n}-\binom{m+n-2d}{n}$.
-We now define the characteristic function of a polynomial over $K$.
 Let $Q=\sum_{{\mathbf{i}}\in I_{Q}}a_{\mathbf{i}}\cdot{\mathbf{x}}^{\mathbf{i}}\in K[x_{1},\ldots,x_{n}]$, where ${\mathbf{i}}=(i_{1},\ldots,i_{n})$, ${\mathbf{x}}^{\mathbf{i}}=x_{1}^{i_{1}}\ldots x_{n}^{i_{n}}$, and $a_{\mathbf{i}}\neq 0$ if ${\mathbf{i}}\in I_{Q}$.
-We let
-(3.9) $\|Q\|_{z}:=\max_{{\mathbf{i}}\in I_{Q}}\{|a_{\mathbf{i}}(z)|\},$
-and define the characteristic function of $Q$ as
-(3.10) $T_{Q,\mathcal{Y}}(r):=T_{[\cdots:a_{\mathbf{i}}:\cdots],\mathcal{Y}}(r),$
-where $a_{\mathbf{i}}$ is taken for every ${\mathbf{i}}\in I_{Q}$.
+We define the parabolic characteristic function of $Q$ as
+(3.27) $T_{Q,\sigma}(r):=\sum_{{\mathbf{i}}\in I_{Q}}T^{M}_{a_{\mathbf{i}},\sigma}(r).$
 Theorem 19 (GCD theorem on parabolic Riemann surfaces evaluating at units).
 Let $n\geq 2$ be a positive integer.
-Let $u_{1},\ldots,u_{n}$ be units on $\mathcal{Y}$, i.e.
-holomorphic functions on $\mathcal{Y}$ without zeros and let ${\bf u}=(u_{1},\ldots,u_{n})$.
+Let $K$ be a subfield of the field of meromorphic functions on $\mathbb{C}$}.
+Let $\mathcal{Y}\subset{\mathbb{C}}$ be as above.
+Let ${\bf u}=(u_{1},\ldots,u_{n})$, where $u_{1},\ldots,u_{n}$ are meromorphic functions on $\mathbb{C}$} such that each $u_{i}|_{\mathcal{Y}},1\leq i\leq n,$ is a unit (i.e.
+without zeros and poles on $\mathcal{Y}$}).
 Let $F,G$ be coprime polynomials in $n$ variables of the same degree $d>0$ over $K$.
-Assume that one of the coefficients in each expansion of $F$ and $G$ is 1.
-Then there exists a positive constant $c$ and a positive integer $b$ such that if the set $\{u_{1}^{i_{1}}\cdots u_{n}^{i_{n}}:i_{1}+\cdots+i_{n}\leq m\}$ is linearly independent over $V_{F,G}(Mb+1)$, then, for any $\delta>0$, we have
-$N_{\rm gcd,{\mathcal{Y}}}(F({\bf u}),G({\bf u}),r)$ $\leq_{\rm exc}\left(\frac{M^{\prime}}{M}+\frac{w}{u}-1+\frac{\epsilon}{M}\right)mn\max_{1\leq i\leq n}T_{u_{i},{\mathcal{Y}}}(r)$
-$+c^{\prime}(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r))+c^{\prime\prime}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(1),$
-where $w=d_{Mb}$, $u=d_{M(b-1)}$, $c^{\prime}=\frac{w}{u}M(b+2)+1$, and $c^{\prime\prime}=M\frac{w^{2}}{u}$, and $M^{\prime}=\binom{m+n}{n}-M$.
+Assume that at least one of the coefficients in the expansions of $F$ and $G$ is equal to 1.
+Let $m\geq d$ and $b$ be a positive integer.
+If the set $\{u_{1}^{i_{1}}\cdots u_{n}^{i_{n}}:i_{1}+\cdots+i_{n}\leq m\}$ is linearly independent over $V_{F,G}(Mb+1)$, then we have
+$N_{\rm gcd,\sigma}(F({\bf u}),G({\bf u}),r)$ $\leq_{\rm exc}\left(\frac{M^{\prime}}{M}+\frac{w}{u}-1\right)mn\max_{1\leq i\leq n}T^{M}_{u_{i},\sigma}(r)+O(\max_{1\leq i\leq n}\log T_{u_{i}}(r))$
+$+c^{\prime}(T_{F,\sigma}(r)+T_{G,\sigma}(r))+c^{\prime\prime}(\mathfrak{X}^{+}_{\sigma}(r)+\log r+c_{1/h,\sigma}(r)+c_{1/W,\sigma}(r)),$
+where $h$ is meromorphic function such that $F({\bf u})/h$ and $G({\bf u})/h$ are entire and have no common zero, and $W$ is the Wronskian which arises when applying Theorem 18, $w=d_{Mb}$, $u=d_{M(b-1)}$, $c^{\prime}=\frac{w}{u}M(b+3)+2,$ and $c^{\prime\prime}=M{w^{2}\over u}$, and $M^{\prime}=\binom{m+n}{n}-M$.
 We outline a proof here.
 Proof.
+For a subset $T\subset K[\mathbf{x}]$, where $\mathbf{x}:=(x_{1},\ldots,x_{n})$, we denote
+$T_{m}=\{f\in T\,:\deg f\leq m\}.$
 As in the proof of [9, Theorem 4.3], we assume that the ideal $(F,G)$ is proper in $K[x]$.
@@
 ${\bf x}^{\bf i}+\sum_{{\bf j}\in I_{z}}c_{z,{\bf j}}{\bf x}^{\bf j}\in(F,G)_{m}$
-for some choice $c_{z,{\bf j}}\in K$. Then, for each such ${\bf i}$, there is a linear form over $K$ such that
+for some choice $c_{z,{\bf j}}\in K$. Then, for each such $\bf i$}, there is a linear form $\tilde{L}_{z,{\bf i}}$ over $K$ such that
 $\tilde{L}_{z,{\bf i}}(\phi_{1},\dots,\phi_{M})={\bf x}^{\bf i}+\sum_{{\bf j}\in I_{z}}c_{z,{\bf j}}{\bf x}^{\bf j}.$
-We note that since there are correspondingly only finitely many choices of a monomial basis of $V_{m}$, there are only finitely many choices of $c_{z,{\bf j}}$, even as $z$ ranges over all of $\mathbb{C}$.
+We note that since there are correspondingly only finitely many choices of a monomial basis of $V_{m}$, there are only finitely many choices of $c_{z,{\bf j}}$, even as $z$ ranges over all of $\mathcal{Y}$}.
 Following the explicit construction in the proof of [9, Theorem 4.3], we can find a nonzero element $c_{z}\in V_{F,G}(M)$ such that the normalized linear form $L_{z,{\bf i}}:=c_{z}\tilde{L}_{z,{\bf i}}$ satisfies
@@
 and
-(3.11) $T_{L_{z,{\bf i}},{\mathcal{Y}}}(r)\leq M(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r)),$
-by [9, (4.12)]. Moreover, the explicit construction also demonstrates that there are only a finitely many choice for $c_{z}$, even though $z$ runs through all of $\mathbb{C}$
+(3.28) $T_{L_{z,{\bf i}},\sigma}(r)\leq M(T_{F,\sigma}(r)+T_{G,\sigma}(r)),$
+by [9, (4.12)]. Moreover, the explicit construction also demonstrates that there are only a finitely many choice for $c_{z}$, even though $z$ runs through all of $\mathcal{Y}$}.
 Following the steps in the proof of Theorem 4.3 in [9], we then obtain
 (3.29) $\log|L_{z,{\bf i}}(\Phi({\bf u}))|-\log\|L_{z,{\bf i}}\|_{z}\leq\log|{\bf u}(z)^{\bf i}|+\log\|F\|_{z}+\log\|G\|_{z}+O(1).$
-Consequently, we have
-(3.13) $\int_{S(r)}\sum_{|{\bf i}|=m,{\bf i}\not\in I_{z}}(\log|L_{z,{\bf i}}(\Phi({\bf u})|-\log\|L_{z,{\bf i}}\|)d\mu_{r}\leq M^{\prime}mnT_{{\bf u},\mathcal{Y}}(r)+M(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r))+O(1).$
-Also, from (3.3) in the proof of [14, Theorem 5.7], we have
-(3.14) $\int_{S(r)}\log\|\Phi({\bf u})\|d\mu_{r}=T_{\Phi({\bf u}),{\mathcal{Y}}}(r)+N_{\rm gcd,{\mathcal{Y}}}(F({\bf u}),G({\bf u}),r)+O(1).$
-Since the set $\{u_{1}^{i_{1}}\cdots u_{n}^{i_{n}}:i_{1}+\cdots+i_{n}\leq m\}$ is linearly independent over $V_{F,G}(Mb+1)$, we have that $\Phi({\bf u}):{\mathbb{C}}\rightarrow{\mathbb{P}}^{M-1}$ is linearly nondegenerate over $V_{F,G}(Mb)=V(b)$.
-We note that, when computing Weil functions, we may assume that the coefficients of $L_{z,{\bf i}}$ are entire functions without common zeros.
+Since the functions $u_{i}$ are units on $\mathcal{Y}$, and that
+(3.30) $\log^{+}|{\bf u}(z)^{\bf i}|\leq\sum_{j=1}^{n}i_{j}\log^{+}|u_{j}(z)|\leq m\sum_{j=1}^{n}\log^{+}|u_{j}(z)|,$
+we have
+(3.31) $T^{M}_{{\bf u}^{\bf i},\sigma}(r)\leq m\sum_{j=1}^{n}T^{M}_{u_{j},\sigma}(r)\leq mn\max_{1\leq i\leq n}T^{M}_{u_{i},\sigma}(r).$
+Consequently, by (3.29), we obtain
+(3.32) $\int_{S^{\sigma}_{r}}\sum_{|{\bf i}|=m,{\bf i}\not\in I_{z}}(\log|L_{z,{\bf i}}(\Phi({\bf u})|-\log\|L_{z,{\bf i}}\|_{z})d\mu_{r}\leq M^{\prime}mn\max_{1\leq i\leq n}T^{M}_{u_{i},\sigma}(r)+M(T_{F,\sigma}(r)+T_{G,\sigma}(r))+O(1).$
+Note that the map $\phi_{1}({\bf u}),\dots,\phi_{M}({\bf u})$ may not be a reduced presentation of $\Phi({\bf u})$.
+Let $h$ be a meromorphic function such that $F({\bf u})/h$ and $G({\bf u})/h$ are entire and have no common zero.
+Let $\psi_{i}:=\phi_{i}({\bf u})/h$ for $i=1,\dots,M$.
+Then $\Psi:=(\psi_{1},\dots,\psi_{M})$ is a reduced representation of $\Phi({\bf u})$.
+Then
+(3.33) $\int_{S^{\sigma}_{r}}\log\|\Phi({\bf u})\|d\mu_{r}=\int_{S^{\sigma}_{r}}\log\|\Psi\|d\mu_{r}+\int_{S^{\sigma}_{r}}\log|h|d\mu_{r}=T_{\Psi,\sigma}(r)+\int_{S^{\sigma}_{r}}\log|h|d\mu_{r}.$
+On the other hand, since $u_{i}$ are units on $\mathcal{Y}$, the poles of $h$ only arises from the poles of coefficients of $F$ and $G$. More precisely,
+(3.34) $N_{h,\sigma}(\infty,r)\leq T_{F,\sigma}(r)+T_{G,\sigma}(r)$
+by applying the inequality of the first main theorem in Theorem 11. It is also clear that this choice of $h$ gives
+(3.35) $N_{h,\sigma}(0,r)=N_{\rm gcd,\sigma}(F({\bf u}),G({\bf u}),r).$
+By Lemma 9, (3.34) and (3.35), we have
+(3.36) $\int_{S^{\sigma}_{r}}\log|h|d\mu_{r}$ $=N_{h,\sigma}(0,r)-N_{h,\sigma}(\infty,r)+c_{h,\sigma}(r)$
+(3.37) $\geq N_{\rm gcd,\sigma}(F({\bf u}),G({\bf u}),r)-\left(T_{F,\sigma}(r)+T_{G,\sigma}(r)\right)+c_{h,\sigma}(r).$
+Then we can derive from (3.33) that
+(3.38) $\int_{S^{\sigma}_{r}}\log\|\Phi({\bf u})\|d\mu_{r}\geq T_{\Psi,\sigma}(r)+N_{\rm gcd,\sigma}(F({\bf u}),G({\bf u}),r)+c_{h,\sigma}(r)-\left(T_{F,\sigma}(r)+T_{G,\sigma}(r)\right).$
+Since the set $\{u_{1}^{i_{1}}\cdots u_{n}^{i_{n}}:i_{1}+\cdots+i_{n}\leq m\}$ is linearly independent over $V_{F,G}(Mb+1)$, it follows that $\Phi({\bf u}):{\mathbb{C}}\rightarrow{\mathbb{P}}^{M-1}$ is linearly independent over $V_{F,G}(Mb)=V(b)$.
+We note that, when computing Weil functions, we have the freedom to assume that the coefficients of $L_{z,{\bf i}}$ are entire functions without common zeros, by multiplying a suitable meromorphic function.
 By this construction, for each $z\in{\mathcal{Y}}$, the hyperplanes defined by evaluating the linear forms $L_{z,{\bf i}}$ at $z$ with $|{\bf i}|=m$ and ${\bf i}\not\in I_{z}$ are in general position.
-Additionally, we observe that (3.11) and (3.13) remain valid under this choice of linear forms.
-Applying Theorem 14 to $\Phi({\bf u})$ and the set of the linear forms $\{L_{z,{\bf i}}:|{\bf i}|=m,{\bf i}\not\in I_{z}\}$, we have the following inequality:
-(3.15) $\int_{S(r)}\sum_{|{\bf i}|=m,{\bf i}\not\in I_{z}}\lambda_{L_{z,{\bf i}}}(\Phi({\bf u}))d\mu_{r}+\frac{1}{u}N_{W}(0,r)$
-(3.16) $\leq_{\operatorname{exc}}\left(\frac{w}{u}M+\epsilon\right)T_{\Phi({\bf u}),\mathcal{Y}}(r)+\frac{w}{u}M^{2}(b+2)(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r))$
-(3.17) $+M^{2}{w^{2}\over u}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(1),$
+For the computation (3.28) and (3.32), however, we will use the original $L_{z,{\bf i}}$.
+By applying Theorem 18 to $\Phi({\bf u})$ and the set of the linear forms $\{L_{z,{\bf i}}:|{\bf i}|=m,{\bf i}\not\in I_{z}\}$, we have the following inequality:
+(3.39) $\int_{S^{\sigma}_{r}}\sum_{|{\bf i}|=m,{\bf i}\not\in I_{z}}\lambda_{L_{z,{\bf i}}}(\Phi({\bf u}))d\mu_{r}+\frac{1}{u}N_{W,\sigma}(0,r)$
+(3.40) $\leq_{\operatorname{exc}}\left(\frac{w}{u}M+\epsilon\right)T_{\Psi,\sigma}(r)+\frac{w}{u}M^{2}(b+2)(T_{F,\sigma}(r)+T_{G,\sigma}(r))$
+(3.41) $+M^{2}{w^{2}\over u}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+{1\over u}c_{1/W,\sigma}(r)+O(\log\max_{1\leq j\leq n}T_{u_{j}}(r)),$
 where
-$\lambda_{L_{z,{\bf i}}}(\Phi({\bf u}))=-\log|L_{z,{\bf i}}(\Phi({\bf u})|+\log\|\Phi({\bf u})\|+\log\|L_{z,{\bf i}}\|+O(1).$
-By combining (3.15) with (3.13), (3.14), we get
-$MN_{\rm gcd,{\mathcal{Y}}}(F({\bf u}),G({\bf u}),r)+\frac{1}{u}N_{W}(0,r)$
-$\leq_{\operatorname{exc}}\left(\frac{w}{u}M-M+\epsilon\right)T_{\Phi({\bf u}),\mathcal{Y}}(r)+M^{\prime}mnT_{{\bf u},\mathcal{Y}}(r)$
-$+\left(\frac{w}{u}M^{2}(b+2)+M\right)(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r))+M^{2}{w^{2}\over u}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(1),$
-which proves the theorem. ∎
-Remark.
-Let $\epsilon$ be a sufficiently small positive number.
-By (3.5), we find $t$ such that $\frac{w}{u}\leq 1+\epsilon$.
-An explicit upper bound for $t$ can be found in the proof of [3, Proposition 5.3] in term of $\epsilon$ and the number of generator for $V(1)$, which is less than $qn$.
-More precisely, if $\epsilon<\frac{1}{16}$, then we can find
-(3.18) $t\leq\frac{3qn}{\epsilon}\log\frac{1}{\epsilon}$
-such that $\frac{w}{u}\leq 1+\epsilon$. Furthermore, for such $t$, we have
-(3.19) $u:=\dim V(t)\leq\left(\frac{1}{\epsilon}\right)^{qn}.$
-Consequently,
-(3.20) $w:=\dim V(t+1)\leq(1+\epsilon)\left(\frac{1}{\epsilon}\right)^{qn}.$
-Therefore, the constants appeared in (3.6) depend only on $\epsilon$ and the number of hyperplanes.
+$\lambda_{L_{z,{\bf i}}}(\Phi({\bf u}))=-\log|L_{z,{\bf i}}(\Phi({\bf u}))|+\log\|\Phi({\bf u})\|+\log\|L_{z,{\bf i}}\|+O(1),$
+and $W$ is the wronskian of $\psi_{1},\dots,\psi_{M}$. By combining (3.39) with (3.32), (3.38) and use (3.36), we get
+$MN_{\rm gcd,\sigma}(F({\bf u}),G({\bf u}),r)+\frac{1}{u}N_{W,\sigma}(0,r)$
+$\leq_{\operatorname{exc}}\left(\frac{w}{u}M-M+\epsilon\right)T_{\Psi,\sigma}(r)+M^{\prime}mn\max_{1\leq i\leq n}T^{M}_{u_{i},\sigma}(r)$
+$+M\left(\frac{w}{u}M(b+2)+1\right)(T_{F,\sigma}(r)+T_{G,\sigma}(r))$
+$+M^{2}{w^{2}\over u}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+{1\over u}c_{{1\over W},\sigma}(r)+Mc_{{1\over h},\sigma}(r)+O(1).$
+Finally, we need to estimate $T_{\Psi,\sigma}(r)$ to complete the proof.
+Recall that $\psi_{i}=\phi(\mathbf{u})/h$ and $\phi_{i}$ is of the form $F{\bf x}^{\bf i}$, or $G{\bf x}^{\bf i}$, which is a combination of monomial of degree $\leq m$ with coefficients coming from $F$ or $G$.
+Therefore, by (3.30)
+$\log\|\psi_{i}\|\leq m\sum_{j=1}^{n}\log^{+}|u_{j}|+\log\|F\|+\log\|G\|+\log\frac{1}{h}.$
+Combining this with Lemma 9 and (3.34), we obtain
+$T_{\Psi,\sigma}(r)\leq mn\max_{1\leq i\leq n}T^{M}_{u_{i},\sigma}(r)+2\left(T_{F,\sigma}(r)+T_{G,\sigma}(r)\right)+c_{\frac{1}{h},\sigma}(r).$
 Theorem 20 (The key theorem).
-Let $n\geq 2$ be a positive integer.
-Let $K$ be a subfield of the field of meromorphic functions on $\mathcal{Y}$.
+Let $\mathcal{Y}$$\subset{\mathbb{C}}$ be given as above with $\sigma$ being associated to the discs $\{{\mathbb{D}}(a_{j},2r_{j})\}$.
+Let $u_{1},\dots,u_{n}$, with $n\geq 2$, be meromorphic functions on $\mathbb{C}$} such that they are units on $\mathcal{Y}$}.
+Let $K$ be a subfield of the field of meromorphic functions on $\mathbb{C}$}.
 Let $F$ and $G$ be coprime polynomials in $K[x_{1},\ldots,x_{n}]$.
-Then, for any $\epsilon>0$, there exist a positive integer $m$ and positive reals $c_{i}$, $1\leq i\leq 4$, all depending only on $\epsilon$, $n$, and the degree of $F$ and $G$ such that for all $n$-tuple $(u_{1},\ldots,u_{n})$ of units on $\mathcal{Y}$ satisfying
-(3.21) $\max_{1\leq i\leq n}T_{u_{i},\mathcal{Y}}(r)\geq c_{1}(\mathfrak{X}^{+}_{\sigma}(r)+\log r),$
-we have either
-(3.22) $T_{u_{1}^{m_{1}}\cdots u_{n}^{m_{n}},\mathcal{Y}}(r)\leq_{\rm exc}c_{2}(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r))+c_{3}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)$
+Assume that at least one of the coefficients in the expansions of $F$ and $G$ is equal to 1.
+Let $\epsilon>0$.
+Then, by shrinking $r_{j}$ small enough (depending on $\epsilon$, $u_{i},1\leq i\leq n$, and on $F$ and $G$), we can choose a positive integer $m$ and positive reals $c_{i}$, $0\leq i\leq 5$, all depending only on $\epsilon$, $n$, and the degree of $F$ and $G$ such that, either
+(3.42) $T^{M}_{u_{1}^{m_{1}}\cdots u_{n}^{m_{n}},\sigma}(r)\leq_{\rm exc}c_{1}\max_{1\leq i\leq n}\{\log T_{u_{i}}(r)\}+c_{2}(T_{F,\sigma}(r)+T_{G,\sigma}(r))+c_{3}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)$
 for some non-trivial tuple of integers $(m_{1},\ldots,m_{n})$ with $|m_{1}|+\cdots+|m_{n}|\leq 2m$, or
-(3.23) $N_{\rm gcd,\mathcal{Y}}(F(u_{1},\ldots,u_{n}),G(u_{1},\ldots,u_{n}),r)\leq_{\rm exc}\epsilon\max_{1\leq i\leq n}T_{u_{i},\mathcal{Y}}(r)+c_{4}(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r)).$
+(3.43) $N_{\rm gcd,\sigma}(F(u_{1},\ldots,u_{n}),G(u_{1},\ldots,u_{n}),r)$ $\leq_{\rm exc}\epsilon\max_{1\leq i\leq n}\{T_{u_{i}}(r)\}+c_{4}(T_{F,\sigma}(r)+T_{G,\sigma}(r))$
+$+c_{5}(\mathfrak{X}^{+}_{\sigma}(r)+\log r).$
 Furthermore, if $\epsilon>0$ is sufficiently small, then we may choose $m=C_{1}\epsilon^{-1}$, where $C_{1}$ is a positive real independent of $\epsilon$.
@@
 We note that $c_{i},1\leq i\leq 5,$ will be determined during the course of the proof.
-As in the proof of [9, Theorem 4.1], we will assume that one of the coefficients in each expansion of $F$ and $G$ is 1.
 Recall that $M:=M_{m}:=2\binom{m+n-d}{n}-\binom{m+n-2d}{n}$ and $M^{\prime}:=M^{\prime}_{m}:=\binom{m+n}{n}-M=O(m^{n-2})$.
-Let $\epsilon>0$.
-To establish (3.22) or (3.23), we can assume that $\epsilon$ is sufficiently small.
 Based on the above estimates, we can choose a real $C_{1}\geq 1$ independent of $\epsilon$ and the $u_{i},1\leq i\leq n,$ such that $m=C_{1}\epsilon^{-1}\geq 2d$, and
@@
 where $w:=\dim_{{\bf k}}V_{F,G}(Mb)$ and $u:=\dim_{{\bf k}}V_{F,G}(Mb-M)$.
-Let ${\bf u}=(u_{1},\ldots,u_{n})$. Suppose that the set $\{u_{1}^{i_{1}}\cdots u_{n}^{i_{n}}:i_{1}+\cdots+i_{n}\leq m\}$ is linearly independent over $V_{F,G}(Mb+1)$. Then, by Theorem 15, we have
-(3.26) $N_{\rm gcd,{\mathcal{Y}}}(F({\bf u}),G({\bf u}),r)$ $\leq_{\rm exc}\left(\frac{M^{\prime}}{M}+\frac{w}{u}-1+\frac{\epsilon}{M}\right)mn\max_{1\leq i\leq n}T_{u_{i},\mathcal{Y}}(r)$
-(3.27) $+c^{\prime}(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r))+c^{\prime\prime}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(1).$
-Let $c_{1}=4c^{\prime\prime}\epsilon^{-1}$. Then (3.21) reads
-$c^{\prime\prime}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(1)\leq\frac{\epsilon}{3}\max_{1\leq i\leq n}T_{u_{i},\mathcal{Y}}(r).$
-Together with (3.24) and (3.25), we derive from (3.26) and the above inequality that
-$N_{\rm gcd,{\mathcal{Y}}}(F({\bf u}),G({\bf u}),r)\leq_{\rm exc}\epsilon\max_{1\leq i\leq n}T_{u_{i},\mathcal{Y}}(r)+c^{\prime}(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r)).$
-Finally, if the set $\{u_{1}^{i_{1}}\cdots u_{n}^{i_{n}}:i_{1}+\cdots+i_{n}\leq m\}$ is linearly dependent over $V_{F,G}(Mb+1)$, i.e. there is a non-trivial relation
-(3.28) $\sum_{\mathbf{i}}\alpha_{\mathbf{i}}\mathbf{u}^{\mathbf{i}}=0,$
-where the sum runs over those $\mathbf{i}\in\mathbb{Z}_{\geq 0}^{n}$ and $|\mathbf{i}|\leq m$, and $\alpha_{\mathbf{i}}\in V_{F,G}(Mb+1)$ for each $\mathbf{i}$ such that $\alpha_{\mathbf{i}_{0}}\neq 0$ for some $\mathbf{i}_{0}$.
-Then we have
-(3.29) $\sum_{\mathbf{i}\neq\mathbf{i}_{0}}\frac{\alpha_{\mathbf{i}}}{\alpha_{\mathbf{i}_{0}}}\mathbf{u}^{\mathbf{i}-\mathbf{i}_{0}}=-1.$
-Since $\alpha_{\mathbf{i}}\in V_{F,G}(Mb+1)$, we have
-(3.30) $T_{\frac{\alpha_{\mathbf{i}}}{\alpha_{\mathbf{i}_{0}}},\mathcal{Y}}(r)\leq T_{\alpha_{\mathbf{i}},\mathcal{Y}}(r)+T_{\alpha_{\mathbf{i_{0}}},\mathcal{Y}}(r)\leq 2(Mb+1)(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r)).$
-By Theorem 12 and (3.30), the relation (3.29) implies
-$T_{\frac{\alpha_{\mathbf{i}}}{\alpha_{\mathbf{i}_{0}}}\mathbf{u}^{\mathbf{i}-\mathbf{i}_{0}},\mathcal{Y}}(r)\leq_{\operatorname{exc}}C_{2}\cdot(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r))+c_{3}(\mathfrak{X}^{+}_{\sigma}(r)+\log r),$
-where $C_{2}=4\binom{n+m}{n}(Mb+1)$ and $c_{3}={\binom{n+m}{n}}\left({\binom{n+m}{n}}+1\right)$. Then
-(3.31) $T_{\mathbf{u}^{\mathbf{i}-\mathbf{i}_{0}},\mathcal{Y}}(r)$ $\leq T_{\frac{\alpha_{\mathbf{i}}}{\alpha_{\mathbf{i}_{0}}}\mathbf{u}^{\mathbf{i}-\mathbf{i}_{0}},\mathcal{Y}}(r)+T_{\frac{\alpha_{\mathbf{i}_{0}}}{\alpha_{\mathbf{i}}},\mathcal{Y}}(r)$
-(3.32) $\leq_{\rm exc}c_{2}(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r))+c_{3}(\mathfrak{X}^{+}_{\sigma}(r)+\log r),$
-where $c_{2}=C_{2}+2(Mb+1)$. ∎
-For the convenience of later application, we state the following result for $n=1$, which can be easily obtained from the theory of resultant (see [19, Proposition 10]).
+Suppose that the set $\{u_{1}^{i_{1}}\cdots u_{n}^{i_{n}}:i_{1}+\cdots+i_{n}\leq m\}$ is linearly independent over $V_{F,G}(Mb+1)$. Then, Theorem 19 implies that, by noticing (3.44) and (3.45),
+$N_{\rm gcd,\sigma}(F({\bf u}),G({\bf u}),r)$ $\leq_{\rm exc}\epsilon\max_{1\leq i\leq n}\{T^{M}_{u_{i},\sigma}(r)\}+{\tilde{c}}\max_{1\leq i\leq n}\{\log T_{u_{i}}(r)\}$
+$~+c_{4}(T_{F,\sigma}(r)+T_{G,\sigma}(r))+c_{5}\left(\mathfrak{X}^{+}_{\sigma}(r)+\log r+c_{{1\over h},\sigma}(r)+c_{{1\over W},\sigma}(r)\right),$
+where $\tilde{c}$} and $c_{i}$, $4\leq i\leq 5$, are computable positive reals.
+If the set $\{u_{1}^{i_{1}}\cdots u_{n}^{i_{n}}:i_{1}+\cdots+i_{n}\leq m\}$ is linearly dependent over $V_{F,G}(Mb+1)$, i.e. there is a non-trivial relation
+(3.46) $\sum_{\mathbf{i}\in I}\alpha_{\mathbf{i}}\mathbf{u}^{\mathbf{i}}=0,$
+where the sum runs over those $\mathbf{i}\in\mathbb{Z}_{\geq 0}^{n}$ and $|\mathbf{i}|\leq m$, $\alpha_{\mathbf{i}}\in V_{F,G}(Mb+1)$ for each $\mathbf{i}$, and $\alpha_{\mathbf{i}}\neq 0$ for $\mathbf{i}\in I$.
+Without loss of generality, we assume that no proper subsum of (3.46) vanishes.
+Then Theorem 16 implies that
+$T^{M}_{\frac{\alpha_{\mathbf{i}}}{\alpha_{\mathbf{i}_{0}}}\mathbf{u}^{\mathbf{i}-\mathbf{i}_{0}},\sigma}(r)$ $\leq_{\operatorname{exc}}\sum_{\mathbf{i}\in I}N_{\beta\alpha_{\mathbf{i}},\sigma}(0,r)+c_{3}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)$
+(3.47) $+c_{\frac{1}{D_{0}},\sigma}(r)+\sum_{\begin{subarray}{c}\mathbf{i}\in I\\ \mathbf{i}\neq\mathbf{i}_{0}\end{subarray}}c_{\beta\alpha_{\mathbf{i}}{\bf u}^{\bf i},\sigma}(r)+O(\max_{0\leq i\leq n}\{\log T_{u_{i}}(r)\})+O(\log(T_{F}(r)+\log F_{G}(r)),$
+where $c_{3}={\binom{n+m}{n}}\left({\binom{n+m}{n}}+1\right)$. We note that
+(3.48) $N_{\beta\alpha_{\mathbf{i}},\sigma}(0,r)\leq N_{\alpha_{\mathbf{i}},\sigma}(0,r)+N_{\beta,\sigma}(0,r)\leq 2(Mb+1)(T_{F,\sigma}(r)+T_{G,\sigma}(r))+c_{{1\over\alpha_{\mathbf{i}}},\sigma}(r),$
+since the zero of $\beta$ comes from the pole of $\alpha_{\mathbf{i}}$ and by (3.14). Also, since $\alpha_{\mathbf{i}}\in V_{F,G}(Mb+1)$,
+$T^{M}_{\frac{\alpha_{\mathbf{i}{0}}}{\alpha_{\mathbf{i}}},\sigma}(r)$ $\leq T^{M}_{1/\alpha_{\mathbf{i}},\sigma}(r)+T^{M}_{\alpha_{\mathbf{i_{0}}},\sigma}(r)\leq T^{M}_{\alpha_{\mathbf{i}},\sigma}(r)+T^{M}_{\alpha_{\mathbf{i_{0}}},\sigma}(r)+c_{{1\over\alpha_{\mathbf{i}}},\sigma}(r)$
+(3.49) $\leq 2(Mb+1)(T_{F,\sigma}(r)+T_{G,\sigma}(r))+c_{{1\over\alpha_{\mathbf{i}}},\sigma}(r).$
+Then, by combining (3.4), (3.48), (3.4),
+$T^{M}_{\mathbf{u}^{\mathbf{i}-\mathbf{i}_{0}},\sigma}(r)$ $\leq T^{M}_{\frac{\alpha_{\mathbf{i}}}{\alpha_{\mathbf{i}_{0}}}\mathbf{u}^{\mathbf{i}-\mathbf{i}_{0}},\sigma}(r)+T^{M}_{\frac{\alpha_{\mathbf{i}_{0}}}{\alpha_{\mathbf{i}}},\sigma}(r)\leq_{\rm exc}c_{2}(T_{F,\sigma}(r)+T_{G,\sigma}(r))+c_{3}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)$
+$+c_{\frac{1}{D_{0}},\sigma}(r)+2c_{{1\over\alpha_{\mathbf{i}}},\sigma}(r)+\sum_{\begin{subarray}{c}\mathbf{i}\in I\\ \mathbf{i}\neq\mathbf{i}_{0}\end{subarray}}c_{\beta\alpha_{\mathbf{i}}{\bf u}_{\mathbf{i}},\sigma}(r)+O(\max_{0\leq i\leq n}\{\log T_{u_{i}}(r)\}),$
+where $c_{2}=4(Mb+1)$.
+By Lemma 10, we can shrink $r_{j}$ small enough so that all of the quantities
+$c^{+}_{1/h,\sigma}(r),\quad c^{+}_{1/W,\sigma}(r),\quad c_{1/D_{0},\sigma}(r),\quad c_{1/\alpha_{\mathbf{i}},\sigma}(r)\ \ (\mathbf{i}\in I),\quad\sum_{\begin{subarray}{c}\mathbf{i}\in I\\ \mathbf{i}\neq\mathbf{i}_{0}\end{subarray}}c_{\beta\alpha_{\mathbf{i}}\mathbf{u}^{\mathbf{i}},\,\sigma}(r)$
+are bounded by $O(1)$, where the implied constant depends on $\mathbf{u}$, $F$, and $G$, but is independent of $r$.
+In addition, we shrink $r_{j}$ small enough such that (3.15) and the Proposition 12 for $u_{i},1\leq i\leq n$, hold, i.e.
+$T_{u_{i}}(r)$ and $T^{M}_{u_{i},\sigma}(r)$ differ only by $\log r+O(1)$.
+This proves the key theorem.
+∎
+For the convenience of later application, we state the following result for $n=1$, which can be easily obtained from the theory of resultant (see [20, Proposition 10]).
 Proposition 21.
-Let $K$ be a subfield of the field of meromorphic functions on $\mathcal{Y}$.
-Let $F,G\in K[x_{0},x_{1}]$ be nonconstant coprime homogeneous polynomials.
-Then there exists a positive constant $c$ such that for any holomorphic functions $g_{0}$ and $g_{1}$ on ${\mathcal{Y}}$ with no common zeros, the following holds:
-$N_{\rm gcd,\mathcal{Y}}(F(g_{0},g_{1}),G(g_{0},g_{1}),r)\leq c(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r)).$
+Let $K$ be a subfield of the field of meromorphic functions on $\mathbb{C}$}.
+Let $F,G\in K[x]$ be nonconstant coprime polynomials.
+Let $\mathcal{Y}$$\subset{\mathbb{C}}$ be given as above with $\sigma$ being associated to the discs $\{{\mathbb{D}}(a_{j},2r_{j})\}$.
+Then there exists a positive constant $c$ such that for any meromorphic function $g$ on $\mathcal{Y}$}, the following holds:
+$N_{\rm gcd,\sigma}(F(g),G(g),r)\leq c(T_{F,\sigma}(r)+T_{G,\sigma}(r)).$
 4. The abc theorem for parabolic Riemann surfaces
-We use the above key theorem (Theorem 16) in the following manner: Let $\mathbf{u}=(u_{1},\ldots,u_{n})$, where $u_{i}$, $1\leq i\leq n$, are holomorphic functions on ${\mathcal{Y}}$ without zeros (i.e.
-$u_{i}$ are the units on ${\mathcal{Y}}$).
-Let $\mathbb{K}_{\mathbf{u}}$ be the subfield of the meromorphic function field such that, for any element $a\in\mathbb{K}_{\mathbf{u}}$, there exists a nonnegative real constant $c_{a}$ (which may depend on $a$) satisfying
-$T_{a,{\mathcal{Y}}}(r)\leq_{\rm exc}c_{a}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(\max_{1\leq i\leq n}\{\log T_{u_{i},{\mathcal{Y}}}(r)\}).$
+Let $\mathcal{Y}$$\subset{\mathbb{C}}$ be given as above with $\sigma$ being associated to the discs $\{{\mathbb{D}}(a_{j},2r_{j})\}$.
+Let ${\bf u}=(u_{1},\ldots,u_{n})$ where $u_{1},\ldots,u_{n}$ are meromorphic functions on $\mathbb{C}$} such that each restriction $u_{i}|_{\mathcal{Y}},1\leq i\leq n,$ is a unit.
+Let $\mathbb{K}_{\mathbf{u}}$ be the set of meromorphic function $a$ on $\mathbb{C}$} for which there exists a nonnegative constant $c_{a}$ (depending on $a$) with
+(4.1) $T^{M}_{a,\sigma}(r)\leq_{\rm exc}c_{a}\left(\mathfrak{X}^{+}_{\sigma}(r)+\log r+\max_{1\leq i\leq n}\{\log T_{u_{i}}(r)\}\right),$
+after shrinking radii $r_{j}$ suffciently small (depending on $a$).
+By (3.14) and Lemma 10, it is easy to see that $\mathbb{K}_{\mathbf{u}}$ is a subfield of the meromorphic function field.
 Notably, in our application, the constant $c_{a}$ can be effectively computed.
-We will use Theorem 16 for polynomials in the ring $\mathbb{K}_{\mathbf{u}}[x_{1},\ldots,x_{n}]$.
-Furthermore, we observe that $u_{i}^{\prime}/u_{i}\in\mathbb{K}_{\mathbf{u}}$ for $1\leq i\leq n$, since
-(4.1) $T_{u_{i}^{\prime}/u_{i},{\mathcal{Y}}}(r)\leq_{\rm exc}{1\over 2}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+2\log T_{u_{i},{\mathcal{Y}}}(r)$
-by Lemma 10.
+We will use Theorem 20 for $K:=\mathbb{K}_{\mathbf{u}}$.
+Note that, by Lemma 13 and by (3.14), if $a\in\mathbb{K}_{\mathbf{u}}$, then $a^{\prime}\in\mathbb{K}_{\mathbf{u}}$, since
+$T^{M}_{a^{\prime},\sigma}(r)$ $\leq T^{M}_{\frac{a^{\prime}}{a},\sigma}(r)+T^{M}_{a,\sigma}(r)\leq m_{a^{\prime}/a,\sigma}(r)+N_{a,\sigma}(0,r)+N_{a,\sigma}(\infty,r)+T^{M}_{a,\sigma}(r)$
+$\leq m_{a^{\prime}/a,\sigma}(r)+T^{M}_{\frac{1}{a},\sigma}(r)+2T^{M}_{a,\sigma}(r).$
 Let $\mathbf{x}:=(x_{1},\ldots,x_{n})$.
@@
 For a non-constant polynomial $F(\mathbf{x})=\sum_{\mathbf{i}}a_{\mathbf{i}}\mathbf{x}^{\mathbf{i}}\in\mathbb{K}_{\mathbf{u}}[\mathbf{x}]:=\mathbb{K}_{\mathbf{u}}[x_{1},\dots,x_{n}]$, we define
-(4.2) $D_{\mathbf{u}}(F)(\mathbf{x}):=\sum_{\mathbf{i}}\frac{(a_{\mathbf{i}}\mathbf{u}^{\mathbf{i}})^{\prime}}{\mathbf{u}^{\mathbf{i}}}\mathbf{x}^{\mathbf{i}}=\sum_{\mathbf{i}}\left(a_{\mathbf{i}}\cdot\sum_{j=1}^{n}i_{j}\frac{u_{j}^{\prime}}{u_{j}}\right)\mathbf{x}^{\mathbf{i}}\in\mathbb{K}_{\mathbf{u}}[\mathbf{x}].$
+(4.2) $D_{\mathbf{u}}(F)(\mathbf{x}):=\sum_{\mathbf{i}}\frac{(a_{\mathbf{i}}\mathbf{u}^{\mathbf{i}})^{\prime}}{\mathbf{u}^{\mathbf{i}}}\mathbf{x}^{\mathbf{i}}=\sum_{\mathbf{i}}\left(a_{\mathbf{i}}^{\prime}+a_{\mathbf{i}}\cdot\sum_{j=1}^{n}i_{j}\frac{u_{j}^{\prime}}{u_{j}}\right)\mathbf{x}^{\mathbf{i}}\in\mathbb{K}_{\mathbf{u}}[\mathbf{x}].$
 A direct computation shows that
@@
 holds for $F,G\in\mathbb{K}_{\mathbf{u}}[\mathbf{x}]$.
+By Lemma 14, we see that $u_{i}^{\prime}/u_{i}\in\mathbb{K}_{\mathbf{u}}$.
+In the following lemma and in Theorem 23, we will identify certain monomial products of bounded degree in the $u_{i}$ that satisfy (4.1), and hence are contained in $\mathbb{K}_{\mathbf{u}}$.
 The following lemma shares the same proof of Lemma 5.1 in [9].
@@
 Let $\mathbf{u}=(u_{1},\ldots,u_{n})$, where $u_{i}$, $1\leq i\leq n$, are the units on $\mathcal{Y}$}.
-Let $G$ be a non-constant polynomial in $\mathbb{K}_{\mathbf{u}}[\mathbf{x}]$ with no monomial factors and no repeated factors.
-Then $F:=D_{\mathbf{u}}(G)$ and $G$ are coprime in $\mathbb{K}_{\mathbf{u}}[\mathbf{x}]$ unless there exists a non-trivial tuple of integers $(m_{1},\ldots,m_{n})$ with $\sum_{i=1}^{n}|m_{i}|\leq 2\deg G$ such that $T_{u_{1}^{m_{1}}\cdots u_{n}^{m_{n}},{\mathcal{Y}}}\leq T_{G,{\mathcal{Y}}}(r).$
+Let $G$ be a non-constant polynomial in $\mathbb{K}_{\mathbf{u}}[\mathbf{x}]$ with no monomial factors and no repeated factors, and assume that at least one of its coefficients is equal to 1.
+Then $F:=D_{\mathbf{u}}(G)$ and $G$ are coprime in $\mathbb{K}_{\mathbf{u}}[\mathbf{x}]$ unless there exists a non-trivial tuple of integers $(m_{1},\ldots,m_{n})$ with $\sum_{i=1}^{n}|m_{i}|\leq 2\deg G$ such that $T^{M}_{u_{1}^{m_{1}}\cdots u_{n}^{m_{n}},\sigma}\leq T_{G,\sigma}(r).$
 Proof.
-The first step of the proof of [9, Lemma 5.1] shows that we may assume that $G$ is irreducible.
+Relation (4.4) implies that we may assume, as in the first step of the proof of [9, Lemma 5.1], that $G$ is irreducible.
 Write $G=\sum_{\mathbf{i}}b_{\mathbf{i}}\mathbf{x}^{\mathbf{i}}\in\mathbb{K}_{\mathbf{u}}[\mathbf{x}]$, which contains at least two distinct terms, since $G$ has no monomial factors.
 Then $D_{\mathbf{u}}(G)/G\in\mathbb{C}^{*}$ since $D_{\mathbf{u}}(G)$ is not zero and $\deg D_{\mathbf{u}}(G)=\deg G$.
-Comparing the coefficients of $G$ and $D_{\mathbf{u}}(G)$, for nonzero $b_{\mathbf{i}}$ and $b_{\mathbf{j}}$ in $\mathbb{K}_{\mathbf{u}}$ ($\mathbf{i}\neq\mathbf{j}$), we have from (4.2) that
-$\frac{(b_{\mathbf{i}}\mathbf{u}^{\mathbf{i}})^{\prime}}{b_{\mathbf{i}}\mathbf{u}^{\mathbf{i}}}=\frac{(b_{\mathbf{j}}\mathbf{u}^{\mathbf{j}})^{\prime}}{b_{\mathbf{j}}\mathbf{u}^{\mathbf{j}}},$
-where $\mathbf{i}=(i_{1},\ldots,i_{n})$, and $\mathbf{j}=(j_{1},\ldots,j_{n})$.
-It implies that $\frac{b_{\mathbf{i}}\mathbf{u}^{\mathbf{i}}}{b_{\mathbf{j}}\mathbf{u}^{\mathbf{j}}}\in\mathbb{C}$.
-Therefore,
-$T_{u_{1}^{i_{1}-j_{1}}\cdots u_{n}^{i_{n}-j_{n}},\mathcal{Y}}\leq T_{G,{\mathcal{Y}}}(r).$
+Comparing the coefficients of $G$ and $D_{\mathbf{u}}(G)$, for nonzero $b_{\mathbf{i}}$ and $b_{\mathbf{j}}=1$ in $\mathbb{K}_{\mathbf{u}}$ ($\mathbf{i}\neq\mathbf{j}$), we have from (4.2) that
+$\frac{(b_{\mathbf{i}}\mathbf{u}^{\mathbf{i}})^{\prime}}{b_{\mathbf{i}}\mathbf{u}^{\mathbf{i}}}=\frac{(\mathbf{u}^{\mathbf{j}})^{\prime}}{\mathbf{u}^{\mathbf{j}}},$
+where $\mathbf{i}=(i_{1},\ldots,i_{n})$, and $\mathbf{j}=(j_{1},\ldots,j_{n})$. It implies that $b_{\mathbf{i}}\frac{\mathbf{u}^{\mathbf{i}}}{\mathbf{u}^{\mathbf{j}}}\in\mathbb{C}$. Therefore,
+$T^{M}_{u_{1}^{j_{1}-i_{1}}\cdots u_{n}^{j_{n}-i_{n}},\sigma}\leq T_{G,\sigma}(r).$
 As an application of the key theorem (Theorem 20) and using the above lemma, we obtain the following theorem.
 Theorem 23.
-Let $\mathbf{u}=(u_{1},\ldots,u_{n})$, where $u_{i}$, $1\leq i\leq n$, are units on ${\mathcal{Y}}$.
+Let $\mathcal{Y}$$\subset{\mathbb{C}}$ be given as above with $\sigma$ being associated to the discs $\{{\mathbb{D}}(a_{j},2r_{j})\}$.
+Let ${\bf u}=(u_{1},\dots,u_{n})$, where each $u_{i}$, $1\leq i\leq n$, is a meromorphic function on $\mathbb{C}$} which is a unit on $\mathcal{Y}$}.
 Let $G$ be a non-constant polynomial in $\mathbb{K}_{\mathbf{u}}[\mathbf{x}]$ with no monomial factors and no repeated factors.
-Then, for any $\epsilon>0$, there exists a positive integer $m$ and a positive constants $\tilde{c}_{1}$ and $\tilde{c}_{2}$ depending only on $\epsilon$, $n$, and the degree and the coefficients of $G$, but independent of $\mathbf{u}$, such that the following holds: If
-(4.5) $\max_{1\leq i\leq n}T_{u_{i},\mathcal{Y}}(r)\geq\tilde{c}_{1}(\mathfrak{X}^{+}_{\sigma}(r)+\log r),$
-then we have either
-(4.6) $T_{u_{1}^{m_{1}}\cdots u_{n}^{m_{n}},\mathcal{Y}}(r)\leq_{\rm exc}\tilde{c}_{2}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(\max_{1\leq j\leq n}\{\log T_{u_{j},{\mathcal{Y}}}(r)\})$
-for a non-trivial $n$-tuple $(m_{1},\ldots,m_{n})$ of integers with $\sum_{i=0}^{n}|m_{i}|\leq 2m$, or
-$N_{G(\mathbf{u}),{\mathcal{Y}}}(0,r)-N^{(1)}_{G(\mathbf{u}),{\mathcal{Y}}}(0,r)\leq_{\rm exc}\epsilon\max_{1\leq j\leq n}\{T_{u_{j},{\mathcal{Y}}}(r)\}.$
+Assume that at least one of the coefficients in the expansions of $G$ is equal to 1.
+Then, by shrinking $r_{j}$ small enough (depending on $\epsilon$, $\bf u$} and $G$), we can find a positive integer $m$ and positive constants $\tilde{c}_{1}$ and $\tilde{c}_{2}$ depending only on $\epsilon$, $n$, and on the degree and the coefficients of $G$, but independent of $\mathbf{u}$, for which the following holds: Either there exists a non-trivial $n$-tuple $(m_{1},\ldots,m_{n})$ of integers with $\sum_{i=0}^{n}|m_{i}|\leq 2m$ such that
+(4.5) $T^{M}_{u_{1}^{m_{1}}\cdots u_{n}^{m_{n}},\sigma}(r)\leq_{\rm exc}\tilde{c}_{1}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(\max_{1\leq j\leq n}\{\log T_{u_{j}}(r)\}),$
+or else
+(4.6) $N_{G(\mathbf{u}),\sigma}(0,r)-N^{(1)}_{G(\mathbf{u}),\sigma}(0,r)\leq_{\rm exc}\epsilon\max_{1\leq j\leq n}\{T_{u_{j}}(r)\}+\tilde{c}_{2}(\mathfrak{X}^{+}_{\sigma}(r)+\log r).$
 Proof.
@@
 Consequently,
-(4.7) $N_{\gcd,{\mathcal{Y}}}(G(\mathbf{u}),D_{\mathbf{u}}(G)(\mathbf{u}),r)\geq N_{G(\mathbf{u}),{\mathcal{Y}}}(0,r)-N^{(1)}_{G(\mathbf{u}),{\mathcal{Y}}}(0,r).$
-Note that there exists a nonnegative real $c_{G}$ such that
-(4.8) $T_{G,{\mathcal{Y}}}(r)\leq_{\rm exc}c_{G}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(\max_{1\leq i\leq n}\{\log T_{u_{i},{\mathcal{Y}}}(r)\}),$
-since the coefficients of $G$ are in $\mathbb{K}_{\mathbf{u}}$. Furthermore,
-(4.9) $T_{D_{\mathbf{u}}(G),{\mathcal{Y}}}(r)$ $\leq 2T_{G,{\mathcal{Y}}}(r)+\deg G\sum_{i=1}^{n}T_{u_{i}^{\prime}/u_{i},{\mathcal{Y}}}(r)$
-(4.10) $\leq_{\rm exc}(2c_{G}+n\deg G)(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(\max_{1\leq i\leq n}\{\log T_{u_{i},{\mathcal{Y}}}(r)\}),$
-and
-(4.11) $T_{G,{\mathcal{Y}}}(r)+T_{D_{\mathbf{u}}(G),{\mathcal{Y}}}(r)\leq_{\rm exc}(3c_{G}+n\deg G)(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(\max_{1\leq i\leq n}\{\log T_{u_{i},{\mathcal{Y}}}(r)\})$
-by (4.8) and (4.9). By Lemma 18 and (4.8), $G$ and $D_{\mathbf{u}}(G)$ are either coprime or
-$T_{u_{1}^{m_{1}}\cdots u_{n}^{m_{n}},\mathcal{Y}}(r)\leq_{\rm exc}T_{G,{\mathcal{Y}}}(r)\leq_{\rm exc}c_{G}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(\max_{1\leq i\leq n}\{\log T_{u_{i},{\mathcal{Y}}}(r)\})$
-for a non-trivial $n$-tuple $(m_{1},\ldots,m_{n})$ of integers with $\sum_{i=0}^{n}|m_{i}|\leq 2\deg G$, i.e., either $G$ and $D_{\mathbf{u}}(G)$ are coprime or (4.6) holds.
+(4.7) $N_{\gcd,\sigma}(G(\mathbf{u}),D_{\mathbf{u}}(G)(\mathbf{u}),r)\geq N_{G(\mathbf{u}),\sigma}(0,r)-N^{(1)}_{G(\mathbf{u}),\sigma}(0,r).$
+Write $G=\sum_{{\bf i}}a_{{\bf i}}{\bf x}^{{\bf i}}$ with $a_{{\bf i}}\in\mathbb{K}_{\mathbf{u}}$.
+Then, from the definition of $\mathbb{K}_{{\bf u}}$, there exists a nonnegative real $c_{G}$ (depending on $G$) such that
+(4.8) $T_{G,\sigma}(r)\leq_{\rm exc}c_{G}\left(\mathfrak{X}^{+}_{\sigma}(r)+\log r+\max_{1\leq i\leq n}\{\log T_{u_{i}}(r)\}\right)$
+after shrinking radii $r_{j}$ suffciently small (depending on $G$). Furthermore, by Lemma 13, we have
+$T_{D_{\mathbf{u}}(G),\sigma}(r)$ $\leq 2T_{G,\sigma}(r)+\deg G\sum_{i=1}^{n}T^{M}_{u_{i}^{\prime}/u_{i},\sigma}(r)$
+(4.9) $\leq_{\rm exc}(2c_{G}+n\deg G)(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(\max_{1\leq i\leq n}\{\log T_{u_{i}}(r)\}).$
+By Lemma 22 and (4.8), either $G$ and $D_{\mathbf{u}}(G)$ are coprime or or there exists a non-trivial $n$-tuple $(m_{1},\ldots,m_{n})\in\mathbb{Z}^{n}$ with $\sum_{i=0}^{n}|m_{i}|\leq 2\deg G$ such that
+$T^{M}_{u_{1}^{m_{1}}\cdots u_{n}^{m_{n}},\sigma}(r)\leq_{\rm exc}T_{G,\sigma}(r)\leq_{\rm exc}c_{G}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(\max_{1\leq i\leq n}\{\log T_{u_{i}}(r)\}).$
 It remains to consider where $G$ and $D_{\mathbf{u}}(G)$ are coprime.
 Let $\epsilon>0$ be given, which may assume to be sufficiently small.
-We apply Theorem 16 and find $m$ and $c_{i}>0$, $1\leq i\leq 4$, such that if (4.5) holds with $\tilde{c}_{1}\geq c_{1}$ , then we have either
-(4.12) $T_{u_{1}^{m_{1}}\cdots u_{n}^{m_{n}},\mathcal{Y}}(r)\leq_{\rm exc}c_{2}(T_{G,{\mathcal{Y}}}(r)+T_{D_{\mathbf{u}}(G),{\mathcal{Y}}}(r))+c_{3}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)$
+By apply Theorem 20, after shrinking $r_{j}$ small enough (depending on $\epsilon$, $\bf u$} and $G$), we find $m$ and $c_{i}>0$, $1\leq i\leq 4$, such that either
+(4.10) $T^{M}_{u_{1}^{m_{1}}\cdots u_{n}^{m_{n}},\sigma}(r)\leq_{\rm exc}c_{1}(T_{G,\sigma}(r)+T_{D_{\mathbf{u}}(G),\sigma}(r))+c_{2}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)$
 for some non-trivial tuple of integers $(m_{1},\ldots,m_{n})$ with $|m_{1}|+\cdots+|m_{n}|\leq 2m$, or
-(4.13) $N_{\rm gcd,\mathcal{Y}}(G({\mathbf{u}}),D_{\mathbf{u}}(G)({\mathbf{u}}),r)\leq_{\rm exc}\frac{\epsilon}{2}\max_{1\leq i\leq n}T_{u_{i},\mathcal{Y}}(r)+c_{4}(T_{G,{\mathcal{Y}}}(r)+T_{D_{\mathbf{u}}(G),{\mathcal{Y}}}(r)).$
-We can conclude the proof by combining (4.11), (4.12), and (4.13), and letting
-(4.14) $\tilde{c}_{1}=\max\{c_{1},2c_{4}(3c_{G}+n\deg G)\epsilon^{-1}\},~{}~{}~{}\mbox{and}~{}~{}\tilde{c}_{2}=\max\{c_{G},c_{3}+3c_{2}c_{G}+nc_{2}\deg G\}.$
+$N_{\rm gcd,\sigma}(G({\mathbf{u}}),D_{\mathbf{u}}(G)({\mathbf{u}}),r)$
+(4.11) $\leq_{\rm exc}{\epsilon\over 2}\max_{1\leq i\leq n}T_{u_{i}}(r)+c_{3}(T_{G,\sigma}(r)+T_{D_{\mathbf{u}}(G),\sigma}(r))+c_{4}(\mathfrak{X}^{+}_{\sigma}(r)+\log r).$
+We can conclude the proof by combining (4.8), (4), (4.10), and (4). ∎
 The proof of [7, Theorem 4] can be adapted to the current situation, leading to the following result.
@@
 Let $G$ be a non-constant polynomial in $\mathbb{C}[x_{1},\ldots,x_{n}]$ with no monomial factors and no repeated factors.
-Then, for any $\epsilon>0$, there exist a positive real number $c_{0}$ and a proper Zariski closed subset $Z$ of $\mathbb{A}^{n}(\mathbb{C})$ such that for all $n$-tuple $\mathbf{u}=(u_{1},\ldots,u_{n})$ of units on ${\mathcal{Y}}$ with $\mathbf{u}(\mathcal{Y})\not\subset Z$ and
-(4.15) $\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}}(r)\}\geq c_{0}(\mathfrak{X}^{+}_{\sigma}(r)+\log r),$
-we have
-(4.16) $N_{G(\mathbf{u}),\mathcal{Y}}(0,r)-N^{(1)}_{G(\mathbf{u}),\mathcal{Y}}(0,r))\leq_{\rm exc}\epsilon\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}}(r)\}$
-if $G(\mathbf{u})\neq 0$.
+Then, for any $\epsilon>0$, there exist a positive real number $c_{0}$, a proper Zariski closed subset $Z$ of $\mathbb{A}^{n}(\mathbb{C})$ expressible as the zero locus of a finite set $\Sigma\subset\mathbb{C}[x_{1},\ldots,x_{n}]$ with the following property: Let ${\bf u}=(u_{1},\dots,u_{n})$, where each $u_{i}$, $1\leq i\leq n$, is meromorphic function on $\mathbb{C}$}, with $H({\mathbf{u}})\neq 0$ for all $H\in\Sigma$ and $G(\mathbf{u})\neq 0$.
+Let $\mathcal{Y}$$\subset{\mathbb{C}}$ be given as above with $\sigma$ being associated to the discs $\{{\mathbb{D}}(a_{j},2r_{j})\}$.
+Assume that $u_{1},\dots,u_{n}$ are units on $\mathcal{Y}$}.
+Then, by shrinking $r_{j}$ small enough (depending on $\bf u$} and $G$), we have
+(4.12) $N_{G(\mathbf{u}),\sigma}(0,r)-N^{(1)}_{G(\mathbf{u}),\sigma}(0,r))\leq_{\rm exc}\epsilon\max_{1\leq i\leq n}\{T_{u_{i}}(r)\}+c_{0}(\mathfrak{X}^{+}_{\sigma}(r)+\log r).$
 Here, $c_{0}$ can be effectively bounded from above in terms of $\epsilon$, $n$, and the degree of $G$.
-Moreover, the exceptional set $Z$ can be expressed as the zero locus of a finite set $\Sigma\subset\mathbb{C}[x_{1},\ldots,x_{n}]$ with the following properties: (Z1) $\Sigma$ depends on $\epsilon$ and $G$, but not $\mathcal{Y}$ and can be determined explicitly, (Z2) $|\Sigma|$ and the degree of each polynomial in $\Sigma$ can be effectively bounded from above in terms of $\epsilon$, $n$, and the degree of $G$.
+Moreover, the finite set $\Sigma\subset\mathbb{C}[x_{1},\ldots,x_{n}]$ satisfies: (Z1) $\Sigma$ depends on $\epsilon$ and $G$, but not on $\bf u$} or $\mathcal{Y}$, and can be determined explicitly, (Z2) $|\Sigma|$ and the degree of each polynomial in $\Sigma$ can be effectively bounded from above in terms of $\epsilon$, $n$, and the degree of $G$.
 We outline a proof here.
@@
 We follow the proof of [7, Theorem 4].
-Let $u_{i},1\leq i\leq n$, be units on ${\mathcal{Y}}$.
 We apply Theorem 23 through several reduction steps to derive a contradiction.
-Assume that (4.16) does not hold and let $c_{0}\geq\tilde{c}_{1}$, where $\tilde{c}_{1}$ is the constant from (4.5) in Theorem 19.
-Then, there exists a positive integer $M_{1}$, depending only on $\epsilon$, $n$, and the degree of $G$, but independent of $\mathbf{u}$, such that
-(4.17) $T_{u_{1}^{m_{1}}\cdots u_{n}^{m_{n}},\mathcal{Y}}(r)\leq_{\rm exc}\tilde{c}_{2}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O\left(\max_{1\leq j\leq n}\{\log T_{u_{j},{\mathcal{Y}}}(r)\}\right)$
-for a non-trivial $n$-tuple $(m_{1},\ldots,m_{n})$ of integers with $\sum|m_{i}|\leq M_{1}$. Thus, we have
-(4.18) $\lambda_{1}:={u_{1}}^{m_{1}}\cdots{u_{n}}^{m_{n}}\in\mathbb{K}_{\mathbf{u}}.$
+The constant $c_{0}$ will be determined in the course of the proof.
+We also remark that, since the coefficients of $G$ lie in $\mathbb{C}$, the assumption that at least one coefficient in the expansion of $G$ is equal to $1$ imposes no restriction.
+By Theorem 23, after shrinking $r_{j}$, we have either
+(4.13) $T^{M}_{u_{1}^{m_{1}}\cdots u_{n}^{m_{n}},\sigma}(r)\leq_{\rm exc}\tilde{c}_{1}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(\max_{1\leq j\leq n}\{\log T_{u_{j}}(r)\})$
+or
+(4.14) $N_{G(\mathbf{u}),\sigma}(0,r)-N^{(1)}_{G(\mathbf{u}),\sigma}(0,r)\leq_{\rm exc}\epsilon\max_{1\leq j\leq n}\{T_{u_{j}}(r)\}+\tilde{c}_{2}(\mathfrak{X}^{+}_{\sigma}(r)+\log r).$
+Let $c_{0}\geq\tilde{c}_{2}$.
+If (4.14) holds, then we are done.
+Assume it is not, then (4.13) holds for a non-trivial $n$-tuple $(m_{1},\ldots,m_{n})$ of integers with $\sum|m_{i}|\leq M_{1}$, where $M_{1}$, depending only on $\epsilon$, $n$, and the degree of $G$, but independent of $\mathbf{u}$.
+Thus, we have
+(4.15) $\lambda_{1}:={u_{1}}^{m_{1}}\cdots{u_{n}}^{m_{n}}\in\mathbb{K}_{\mathbf{u}}.\$
 We may assume $\gcd(m_{1},\ldots,m_{n})=1$.
@@
 Consider $G_{1}(\lambda_{1},X_{1,2},\ldots,X_{1,n})\in\mathbb{C}(\lambda_{1})[X_{1,2},\ldots,X_{1,n}]\subset{\mathbb{K}_{\mathbf{u}}}[X_{1,2},\ldots,X_{1,n}]$.
-For the particular change of variables in (4.22), (4.23), and (4.24) (that depends on the matrix $A_{1}$), we apply the Lemma 16 in [7] to find a nonconstant polynomial $H_{1}^{\prime}\in\mathbb{C}[x_{1},\ldots,x_{n}]$ such that $G_{1}(\lambda_{1},X_{1,2},\ldots,X_{1,n})$ has neither monomial nor repeated irreducible factors if $H_{1}^{\prime}(u_{1},\ldots,u_{n})\not\equiv 0$.
-It is important to note that the polynomial $H_{1}^{\prime}$ is not only independent of $\mathbf{u}$, but also independent of $\mathcal{Y}$.
-We now take $H_{1}$ to be the product of all such $H_{1}^{\prime}$ where $A_{1}$ ranges over the finitely many elements of $\operatorname{GL}_{n}(\mathbb{Z})$ with $\|A_{1}\|_{\infty}\leq M_{1}+n$.
+For the particular change of variables in (4.19), (4.20), and (4.21) (that depends on the matrix $A_{1}$), we apply the Lemma 16 in [7] to find a nonconstant polynomial $\widetilde{H}_{1}\in\mathbb{C}[x_{1},\ldots,x_{n}]$ such that $G_{1}(\lambda_{1},X_{1,2},\ldots,X_{1,n})$ has neither monomial nor repeated irreducible factors if $\widetilde{H}_{1}(u_{1},\ldots,u_{n})\not\equiv 0$.
+It is important to note that the polynomial $\widetilde{H}_{1}$ is not only independent of $\mathbf{u}$, but also independent of $\mathcal{Y}$.
+We now take $H_{1}$ to be the product of all such $\widetilde{H}_{1}$ where $A_{1}$ ranges over the finitely many elements of $\operatorname{GL}_{n}(\mathbb{Z})$ with $\|A_{1}\|_{\infty}\leq M_{1}+n$.
 From Lemma 16 in [7]  we know that $\deg H_{1}$ depends only on $\epsilon$, $n$ and $\deg G$.
 Since the $u_{i}$’s, $\lambda_{1}$, and $\beta_{1,j}$’s are the units on $\mathcal{Y}$, we have
-(4.25) $\begin{split}N_{G(\mathbf{u}),\mathcal{Y}}(0,r)-N_{G(\mathbf{u}),\mathcal{Y}}^{(1)}(0,r)=N_{G_{1}(\lambda_{1},\beta_{1,2},\ldots,\beta_{1,n}),\mathcal{Y}}(0,r)-N_{G_{1}(\lambda_{1},\beta_{1,2},\ldots,\beta_{1,n}),\mathcal{Y}}^{(1)}(0,r).\end{split}$
-On the other hand, by taking $\tilde{c}_{1}\geq 2\tilde{c}_{2}\epsilon^{-1}$, we deduce from (4.15) and (4.17) that
-(4.26) $T_{{u_{1}}^{m_{1}}\cdots{u_{n}}^{m_{n}},\mathcal{Y}}(r)\ \leq_{\rm exc}\epsilon\max_{1\leq j\leq n}\{T_{u_{j},{\mathcal{Y}}}(r)\}.$
-Then, by (4.22), (4.23) and (4.26), we have:
-$\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}}(r)\}={\rm O}\left(\max\{T_{\lambda_{1},\mathcal{Y}}(r),T_{\beta_{1,2},\mathcal{Y}}(r),\ldots,T_{\beta_{1,n},\mathcal{Y}}(r)\}\right)={\rm O}\left(\max_{2\leq i\leq n}\{T_{\beta_{1,i},\mathcal{Y}}(r)\}\right).$
-More precisely, we find positive real numbers $b_{1,1}$ and $b_{1,2}$ such that
-(4.27) $b_{1,1}\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}}(r)\}\leq\max_{2\leq i\leq n}\{T_{\beta_{1,i},\mathcal{Y}}(r)\},\quad\text{and }\quad\max_{2\leq i\leq n}\{T_{\beta_{1,i},\mathcal{Y}}(r)\}\leq b_{1,2}\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}}(r)\}.$
-In conclusion, at the end of this step we have that
-$\max_{2\leq i\leq n}\{T_{\beta_{1,i},\mathcal{Y}}(r)\}\geq b_{1,1}c_{0}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)$
-holds, and the inequality
-(4.28) $N_{G_{1}(\lambda_{1},\beta_{1,2},\ldots,\beta_{1,n})}(0,r)-N_{G_{1}(\lambda_{1},\beta_{1,2},\ldots,\beta_{1,n})}^{(1)}(0,r)$ $\leq_{\rm exc}\frac{\epsilon}{b_{1,2}}\max_{2\leq i\leq n}\{T_{\beta_{1,i},\mathcal{Y}}(r)\}$
+(4.22) $\begin{split}N_{G(\mathbf{u}),\sigma}(0,r)-N_{G(\mathbf{u}),\sigma}^{(1)}(0,r)=N_{G_{1}(\lambda_{1},\beta_{1,2},\ldots,\beta_{1,n}),\sigma}(0,r)-N_{G_{1}(\lambda_{1},\beta_{1,2},\ldots,\beta_{1,n}),\sigma}^{(1)}(0,r).\end{split}$
+By (4.19) and (4.20), we have:
+(4.23) $\max_{1\leq i\leq n}\{T_{u_{i}}(r)\}={\rm O}\left(\max\{T_{\lambda_{1}}(r),T_{\beta_{1,2}}(r),\ldots,T_{\beta_{1,n}}(r)\}\right).$
+By (4.16), there are only a finitely many choice of $\beta_{1,t}$, $2\leq t\leq n$. Hence, by Proposition 12, after possibly shrinking $r_{j}$ (depending on $u_{i}$ ), we may assume that
+$\bigl|T_{u_{i}}(r)-T^{M}_{u_{i},\sigma}(r)\bigr|,\quad\bigl|T_{\beta_{1,t}}(r)-T^{M}_{\beta_{1,t},\sigma}(r)\bigr|,\quad\text{and}\quad\bigl|T_{\lambda_{1}}(r)-T^{M}_{\lambda_{1},\sigma}(r)\bigr|$
+are all bounded by $\log r+O(1)$ for $1\leq i\leq n$ and $2\leq t\leq n$. Then by (4.23) and (4.13), we find positive real numbers $b_{1,1}$ and $b_{1,2}$ such that
+$b_{1,1}\max_{1\leq i\leq n}\{T_{u_{i}}(r)\}\leq_{\rm exc}\max_{2\leq t\leq n}\{T_{\beta_{1,t}}(r)\}+\tilde{c}_{1}(\mathfrak{X}^{+}_{\sigma}(r)+\log r),\quad\text{and}$
+$\max_{2\leq t\leq n}\{T_{\beta_{1,t}}(r)\}\leq b_{1,2}\left(\max_{1\leq i\leq n}\{T_{u_{i}}(r)\}+\log r\right).$
+In conclusion, at the end of this step we have that, after possibly shrinking $r_{j}$, the inequality
+(4.24) $N_{G_{1}(\lambda_{1},\beta_{1,2},\ldots,\beta_{1,n}),\sigma}(0,r)-N_{G_{1}(\lambda_{1},\beta_{1,2},\ldots,\beta_{1,n}),\sigma}^{(1)}(0,r)$ $\leq_{\rm exc}\frac{\epsilon}{b_{1,2}}\max_{2\leq i\leq n}\{T_{\beta_{1,i}}(r)\}+c_{0}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)$
 fails to hold under the assumption that $H_{1}(u_{1},\ldots,u_{n})\not\equiv 0$.
 The above is the first step.
-There are $(n-1)$ many steps in total as in the proof of [7, Theorem 4].
-Let $12\leq s\leq n-$ 1 and suppose that we have completed the step $s-1$.
+The next step is to apply Theorem 23 to $G_{1}$ and $\bf u$}.
+Note that, to do so, we need to further shrink $r_{j}$ according to $G_{1}$.
+There are (n1)(n-1) many steps in total as in the proof of [7, Theorem 4].
+Let $2\leq s\leq n-1$ and suppose that we have completed the step s1s-1.
 This includes the construction of $H_{s-1}\in{\mathbb{C}}[x_{1},\ldots,x_{n}]$ with degree depends on $\epsilon$, $n$ and $\deg G$ only.
-The last one is Step $(n-1)$ resulting in $H_{n-1}\in{\mathbb{C}}[x_{1},\ldots,x_{n}]$.
+The last one is Step (n1)(n-1) resulting in $H_{n-1}\in{\mathbb{C}}[x_{1},\ldots,x_{n}]$.
 We now define $H=H_{1}\cdots H_{n-1}$.
@@
 Moreover, the polynomial $H$ is not only independent of $\mathbf{u}$, but also independent of $\mathcal{Y}$.
-Suppose $H(u_{1},\ldots,u_{n})\not\equiv 0$. Assume we go through all the $(n-1)$ steps to get the polynomial
+Suppose $H(u_{1},\ldots,u_{n})\not\equiv 0$. Assume we go through all the (n1)(n-1) steps to get the polynomial
 $P(X_{n-1,n}):=G_{n-1}(\lambda_{1},\ldots,\lambda_{n-1},X_{n-1,n})\in{\mathbb{K}_{\mathbf{u}}}[X_{n-1,n}]$
 such that its degree can be bounded explicitly in terms of $M_{n-1},\ldots,M_{1}$, $n$, and $\deg(G)$.
-At the end of the step $(n-1)$, we have that $\beta:=\beta_{n-1,n}$, which is a (non-constant) unit on ${\mathcal{Y}}$, such that
-(4.29) $T_{\beta,\mathcal{Y}}(r)\geq b_{n-1,1}c_{0}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)$
-holds and the inequality
-(4.30) $N_{P(\beta)}(0,r)-N^{(1)}_{P(\beta)}(0,r)\leq_{\rm exc}\frac{\epsilon}{b_{n-1,2}}T_{\beta,\mathcal{Y}}(r)$
+At the end of the step (n1)(n-1), we have that $\beta:=\beta_{n-1,n}$, which is a (non-constant) unit on $\mathcal{Y}$}, such that the inequality, after further shrinking $r_{j}$ according to $G,G_{1},\dots,G_{n-1}$,
+(4.25) $N_{P(\beta),\sigma}(0,r)-N^{(1)}_{P(\beta),\sigma}(0,r)\leq_{\rm exc}\frac{\epsilon}{b_{n-1,2}}T_{\beta}(r)+c_{0}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)$
 fails to hold.
@@
 On the other hand, since $H_{n-1}(u_{1},\ldots,u_{n})\not\equiv 0$, the polynomial $P(X_{n-1,n})$ has neither monomial nor repeated irreducible factors, and thus, similar to (4.7), we have
-(4.31) $N_{\gcd,{\mathcal{Y}}}(P(\beta),D_{\beta}(P)(\beta),r)\geq N_{P(\beta)}(0,r)-N^{(1)}_{P(\beta)}(0,r).$
+(4.26) $N_{\gcd,\sigma}(P(\beta),D_{\beta}(P)(\beta),r)\geq N_{P(\beta),\sigma}(0,r)-N^{(1)}_{P(\beta),\sigma}(0,r).$
 By Proposition 21, there is a positive constant $C_{1}$ independent of $\beta$ such that
-(4.32) $N_{\gcd,{\mathcal{Y}}}(P(\beta),D_{\beta}(P)(\beta),r)\leq C_{1}(T_{P,{\mathcal{Y}}}(r)+T_{D_{\beta}(P),{\mathcal{Y}}}(r)).$
-Similar to the arguments for (4.9), we find a positive constant $C_{2}$ independent of $\beta$ such that
-(4.33) $T_{P,{\mathcal{Y}}}(r)+T_{D_{\beta}(P),{\mathcal{Y}}}(r)\leq_{\rm exc}C_{2}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(\log T_{\beta,\mathcal{Y}}(r)).$
-Then by (4.29), (4.31), (4.32), and (4.33), we have
-(4.34) $N_{P(\beta)}(0,r)-N^{(1)}_{P(\beta)}(0,r)\leq_{\rm exc}\frac{C_{1}C_{2}}{b_{n-1,1}c_{0}}T_{\beta,\mathcal{Y}}(r)+O(\log T_{\beta,\mathcal{Y}}(r)),$
-which contradicts (4.30) by taking $c_{0}\geq 2C_{1}C_{2}b_{n-1,2}b_{n-1,1}^{-1}\epsilon^{-1}$. ∎
+(4.27) $N_{\gcd,\sigma}(P(\beta),D_{\beta}(P)(\beta),r)\leq C_{1}(T_{P,\sigma}(r)+T_{D_{\beta}(P),\sigma}(r)).$
+Similar to the arguments for (4), after further shrinking $r_{j}$ according to $P$, we find a positive constant $C_{2}$ independent of $\beta$ such that
+(4.28) $T_{P,\sigma}(r)+T_{D_{\beta}(P),\sigma}(r)\leq_{\rm exc}C_{2}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(\log T_{\beta}(r)).$
+Hence,
+(4.29) $N_{P(\beta),\sigma}(0,r)-N^{(1)}_{P(\beta),\sigma}(0,r)\leq_{\rm exc}C_{1}C_{2}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(\log T_{\beta}(r)),$
+which contradicts that fact that (4.25) fails to hold by taking $c_{0}\geq C_{1}C_{2}$. ∎
 Theorem 25.
-Let $G=\sum_{\mathbf{i}\in I_{G}}\alpha_{\mathbf{i}}{\mathbf{x}}^{\mathbf{i}}\in$ be a non-constant polynomial in $\mathbb{C}[x_{1},\cdots,x_{n}]$.
+Let $\mathcal{Y}$$\subset{\mathbb{C}}$ be given as above with $\sigma$ being associated to the discs $\{{\mathbb{D}}(a_{j},2r_{j})\}$.
+Let $G=\sum_{\mathbf{i}\in I_{G}}\alpha_{\mathbf{i}}{\mathbf{x}}^{\mathbf{i}}$ be a non-constant polynomial in $\mathbb{C}[x_{1},\cdots,x_{n}]$.
 Assume that $G(0,\ldots,0)\neq 0$ and $\deg_{x_{i}}G=\deg G=d$ for $1\leq i\leq n$.
-Let $Z_{1}$ be the Zariski closed subset that is the union of hypersurfaces of $\mathbb{A}^{n}$ of the form $\sum_{\mathbf{i}\in J}\alpha_{\mathbf{i}}{\mathbf{x}}^{\mathbf{i}}=0$, where $J$ is a non-empty subset of $I_{G}$.
-Let $u_{1},\ldots,u_{n}$ be the units on $\mathcal{Y}$ such that the image of ${\bf u}:=(u_{1},\ldots,u_{n}):\mathcal{Y}\to(\mathbb{C}^{*})^{n}$ is not contained in $Z_{1}$.
-Then
-(4.35) $N_{G({\bf u}),\mathcal{Y}}(0,r)\geq_{\operatorname{exc}}(\deg G-\epsilon)\cdot\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}}(r)\}-c(\mathfrak{X}^{+}_{\sigma}(r)+\log r),$
+Let $f_{0},\dots,f_{n}$ be entire functions without common zero which are units on $\mathcal{Y}$.
+Let $u_{i}:=f_{i}/f_{0}$, $1\leq i\leq n$, and ${\bf u}=(u_{1},\dots,u_{n})$.
+Assume that $\sum_{\mathbf{i}\in J}\alpha_{\mathbf{i}}{\mathbf{u}}^{\mathbf{i}}\neq 0$ for any non-empty subset $J$ of $I_{G}$.
+Then, for every $\epsilon>0$, by shrinking $r_{j}$ small enough (depending on $\epsilon$, $\bf u$} and $G$), we have
+(4.30) $N_{G({\bf u}),\sigma}(0,r)\geq_{\operatorname{exc}}(\deg G-\epsilon)\cdot\max_{1\leq i\leq n}\{T_{u_{i}}(r)\}-c(\mathfrak{X}^{+}_{\sigma}(r)+\log r),$
 where $c=\frac{1}{2}\binom{n+d}{n}(\binom{n+d}{n}+1)$.
@@
 where $\alpha_{\mathbf{i}_{i}}\neq 0$ for $0\leq i\leq n$ and $I=\{\mathbf{i}_{0}:=(0,0,\ldots,0),\mathbf{i}_{1}:=(d,0,\ldots,0),\ldots,\mathbf{i}_{n}:=(0,\ldots,0,d)\}$.
-If ${\mathbf{u}}({\mathbb{C}})$ is not contained in $Z_{1}$, we may use Theorem 12 to show that
-$(d-\epsilon)\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}}(r)\}\leq N_{G(\mathbf{u}),\mathcal{Y}}(0,r)+c(\mathfrak{X}^{+}_{\sigma}(r)+\log r),$
-where $c=\frac{1}{2}\binom{n+d}{n}(\binom{n+d}{n}+1)$. ∎
+Since $\sum_{\mathbf{i}\in J}\alpha_{\mathbf{i}}{\mathbf{u}}^{\mathbf{i}}\neq 0$, for any non-empty subset $J$ of $I_{G}$, we may use Theorem 16 to show that
+$d\max_{1\leq i\leq n}\{T^{M}_{u_{i},\sigma}(r)\}\leq N_{G(\mathbf{u}),\sigma}(0,r)+c(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+c^{+}_{h_{\mathbf{u}},\sigma}(r)+O(\max_{1\leq i\leq n}\log T_{u_{i}}(r)),$
+where $c=\frac{1}{2}\binom{n+d}{n}(\binom{n+d}{n}+1)$, and $h_{\mathbf{u}}$ is a meromorphic function depending on $\bf u$}.
+By Lemma 10, by shrinking $r_{j}$ small enough if necessary, we can make $c^{+}_{h_{\mathbf{u}},\sigma}(r)\leq O(1)$.
+In addition, we can shrink $r_{j}$ small enough such that (3.15) and the Proposition 12 for $u_{i},1\leq i\leq n$, hold, i.e., $|T^{M}_{u_{i},\sigma}(r)-T_{u_{i}}(r)|\leq\log r+O(1)$.
+This proves the theorem.
+∎
 Theorem 26 (The abc Theorem on parabolic Riemann surfaces).
+Let $\mathcal{Y}$$\subset{\mathbb{C}}$ be given with $\sigma$ being associated to the discs $\{{\mathbb{D}}(a_{j},2r_{j})\}$.
 Let $G$ be a non-constant polynomial in $\mathbb{C}[x_{1},\ldots,x_{n}]$ with no monomial factors and no repeated factors.
-Then, for any $\epsilon>0$, there exist a positive real number $c_{0}$ and a proper Zariski closed subset $W$ of $\mathbb{A}^{n}(\mathbb{C})$ such that for all holomorphic maps $\mathbf{u}=(u_{1},\ldots,u_{n}):\mathcal{Y}\to(\mathbb{C}^{*})^{n}$ with $\mathbf{u}(\mathcal{Y})\not\subset W$ and satisfying
-(4.37) $\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}}(r)\}\geq c_{0}(\mathfrak{X}^{+}_{\sigma}(r)+\log r),$
-we have the following two statements hold
+Then, for any $\epsilon>0$, there exist a positive real number $c_{0}$, a proper Zariski closed subset $W$ of $\mathbb{A}^{n}(\mathbb{C})$ expressible as the zero locus of a finite set $\Sigma\subset\mathbb{C}[x_{1},\ldots,x_{n}]$ with the following property: Let $f_{0},\dots,f_{n}$ be entire functions without common zeros, which are units on $\mathcal{Y}$.
+For $1\leq i\leq n$, set $u_{i}:=f_{i}/f_{0}$ and write ${\bf u}=(u_{1},\dots,u_{n})$.
+Assume that $H({\bf u})\neq 0$ for all $H\in\Sigma$.
+Then, by shrinking $r_{j}$ small enough (depending on $\epsilon$, $\bf u$} and $G$), we have
   1. (a)
-    $N_{G(\mathbf{u}),\mathcal{Y}}(0,r)-N^{(1)}_{G(\mathbf{u}),\mathcal{Y}}(0,r))\leq_{\rm exc}\epsilon\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}}(r)\}$.
+    $N_{G(\mathbf{u}),\sigma}(0,r)-N^{(1)}_{G(\mathbf{u}),\sigma}(0,r))\leq_{\rm exc}\epsilon\max_{1\leq i\leq n}\{T_{u_{i}}(r)\}+c_{0}(\mathfrak{X}^{+}_{\sigma}(r)+\log r).$
   2. (b)
     If $G(0,\ldots,0)\neq 0$ and $\deg_{x_{i}}G=\deg G=d$ for $1\leq i\leq n$, then
-    $N^{(1)}_{G(\mathbf{u}),\mathcal{Y}}(0,r))\geq_{\rm exc}(\deg G-\epsilon)\cdot\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}}(r)\}.$
-Here, $c_{0}$ can be effectively bounded from above in terms of $\epsilon$, $n$, and the degree of $G$.
-Moreover, the exceptional set $W$ can be expressed as the zero locus of a finite set $\Sigma\subset K[x_{1},\ldots,x_{n}]$ with the following properties: (Z1) $\Sigma$ depends on $\epsilon$ and $G$ (not on $\mathcal{Y}$) and can be determined explicitly, (Z2) $|\Sigma|$ and the degree of each polynomial in $\Sigma$ can be effectively bounded from above in terms of $\epsilon$, $n$, and the degree of $G$.
+    $(\deg G-\epsilon)\cdot\max_{1\leq i\leq n}\{T_{u_{i}}(r)\}\leq_{\rm exc}N^{(1)}_{G(\mathbf{u}),\sigma}(0,r))+c_{1}(\mathfrak{X}^{+}_{\sigma}(r)+\log r).$
+Here, $c_{0}$ and $c_{1}$ can be effectively bounded from above in terms of $\epsilon$, $n$, and the degree of $G$.
+Moreover, the finite set $\Sigma\subset\mathbb{C}[x_{1},\ldots,x_{n}]$ satisfies: (Z1) $\Sigma$ depends on $\epsilon$ and $G$, but not on $\bf u$} or $\mathcal{Y}$, and can be determined explicitly, (Z2) $|\Sigma|$ and the degree of each polynomial in $\Sigma$ can be effectively bounded from above in terms of $\epsilon$, $n$, and the degree of $G$.
 Proof.
-The proof follows from Theorem 20 and Theorem 6.4.
-Let $W=Z\cup Z_{1}\cup[G=0]$, where $Z$ and $Z_{1}$ are the exceptional set from Theorem 20 and Theorem 6.4 for $\frac{\epsilon}{2}$ respectively.
-Let $\mathbf{u}=(u_{1},\ldots,u_{n}):\mathcal{Y}\to(\mathbb{C}^{*})^{n}$ with $\mathbf{u}(\mathcal{Y})\not\subset W$.
-Theorem 20 implies that
-$N_{G(\mathbf{u}),\mathcal{Y}}(0,r)-N^{(1)}_{G(\mathbf{u}),\mathcal{Y}}(0,r))\leq\frac{\epsilon}{2}\max_{1\leq i\leq n}\{T_{u_{i}}(r)\}.$
-Together with (4.35) in Theorem 6.4, we derive
-(4.38) $N^{(1)}_{G(\mathbf{u}),\mathcal{Y}}(0,r))\geq_{\operatorname{exc}}\left(\deg G-\frac{\epsilon}{2}\right)\max_{1\leq i\leq n}\{T_{u_{i}}(r)\}-c(\mathfrak{X}^{+}_{\sigma}(r)+\log r).$
-By replacing the constant $c_{0}$ in (4.37) by $c_{1}=\max\{c_{0},2c\epsilon^{-1}\}$, we may derive from (4.38) that
-$N^{(1)}_{G(\mathbf{u}),\mathcal{Y}}(0,r))\geq_{\operatorname{exc}}(\deg G-\epsilon)\cdot\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}}(r)\}.$
+The proof follows from Theorem 24 and Theorem 25.
+Let $W=Z\cup Z_{1}\cup[G=0]$, where $Z$ and $Z_{1}$ are the exceptional sets arising from Theorem 24 and Theorem 25 for $\frac{\epsilon}{2}$, respectively.
+We express the zero locus of $W$ as the vanishing set of a finite collection $\Sigma\subset\mathbb{C}[x_{1},\ldots,x_{n}]$, and assume that $H({\bf u})\neq 0$ for all $H\in\Sigma$.
+By Theorem 24, we obtain
+$N_{G(\mathbf{u}),\sigma}(0,r)-N^{(1)}_{G(\mathbf{u}),\sigma}(0,r))\leq\frac{\epsilon}{2}\max_{1\leq i\leq n}\{T_{u_{i}}(r)\}+c_{0}(\mathfrak{X}^{+}_{\sigma}(r)+\log r).$
+On the other hand, from (4.30) in Theorem 25,
+$N_{G({\bf u}),\sigma}(0,r)\geq_{\operatorname{exc}}(\deg G-\epsilon)\cdot\max_{1\leq i\leq n}\{T_{u_{i}}(r)\}-c(\mathfrak{X}^{+}_{\sigma}(r)+\log r).$
+Thus we derive
+(4.32) $N^{(1)}_{G(\mathbf{u}),\sigma}(0,r))\geq_{\operatorname{exc}}\left(\deg G-\frac{\epsilon}{2}\right)\max_{1\leq i\leq n}\{T_{u_{i}}(r)\}-c(\mathfrak{X}^{+}_{\sigma}(r)+\log r).$
+Let $c_{1}=c_{0}+c$, then
+$N^{(1)}_{G(\mathbf{u}),\sigma}(0,r))\geq_{\operatorname{exc}}(\deg G-\epsilon)\cdot\max_{1\leq i\leq n}\{T_{u_{i}}(r)\}+c_{1}(\mathfrak{X}^{+}_{\sigma}(r)+\log r).$
 5. Proof of the main theorem
@@
 $\log r=o(T_{f}(r)).$
-Let $D:=\sum_{i=0}^{n}H_{i}$, where $H_{i}:=[x_{i}=0]$, $0\leq i\leq n$, are the coordinate hyperplanes. We consider ${\mathcal{Y}_{f}}:={\mathbb{C}}\backslash f^{-1}(D)$. By (ii) of Proposition 8,
-(5.1) $\mathfrak{X}_{\sigma}^{+}(r)\leq\sum_{i=0}^{n}N_{f}^{(1)}(H_{i},r)+O(\log r).$
-Since ${\rm mult}_{z_{0}}(f^{*}H_{i})\geq\ell$ for all $i$ with $0\leq i\leq n$ and for every $z_{0}\in{\mathbb{C}}$, we have $N_{f}^{(1)}(H_{i},r)\leq{1\over\ell}T_{f}(r)$ for $i=0,1,\dots,n$. Thus
+Let $D:=\sum_{i=0}^{n}H_{i}$, where $H_{i}:=[x_{i}=0]$, $0\leq i\leq n$, are the coordinate hyperplanes.
+We consider ${\mathcal{Y}_{f}}:={\mathbb{C}}\backslash f^{-1}(D)$ with $\sigma$ being associated to the discs $\{{\mathbb{D}}(a_{j},2r_{j})\}$, as being constructed in Section 3.1.
+Furthermore, by further shrinking $r_{j}$, we may assume that
+(5.1) $|T_{u_{i}}(r)-T^{M}_{u_{i},\sigma}(r)|\leq\log r+O(1)$
+for $1\leq i\leq n$ by (3.15) and Proposition 12 (ii), and
+(5.2) $\mathfrak{X}_{\sigma}^{+}(r)\leq\sum_{i=0}^{n}N_{f}^{(1)}(H_{i},r)+O(\log r)$
+by (3.17).
+Since ${\rm mult}_{z_{0}}(f^{*}H_{i})\geq\ell$ for all $i$ with $0\leq i\leq n$ and for every $z_{0}\in{\mathbb{C}}$, we have $N_{f}^{(1)}(H_{i},r)\leq{1\over\ell}T_{f}(r)$ for $i=0,1,\dots,n$.
+Thus
 (5.3) $\mathfrak{X}_{\sigma}^{+}(r)\leq\frac{n+1}{\ell}T_{f}(r)+O(\log r),$
 where $\ell$ is a positive integer to be determined.
-Let $\tilde{f}_{i}$ be the restriction of $f_{i}$ to $\mathcal{Y}_{f}$ and let $u_{i}:=\frac{\tilde{f}_{i}}{\tilde{f}_{0}}$.
+Let $u_{i}:=\frac{f_{i}}{f_{0}}$.
 Then $u_{i},1\leq i\leq n$, are units on $\mathcal{Y}_{f}$.
-Denote by ${\tilde{u}}=(1,u_{1},\ldots,u_{n})$ and $\mathbf{u}=(u_{1},\ldots,u_{n})$.
-Then $\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}_{f}}(r)\}=T_{\tilde{u},\mathcal{Y}_{f}}(r)+O(1)=T_{{\tilde{f}},\mathcal{Y}_{f}}(r)+O(1),$ where $\tilde{f}$ is the restriction of $f$ to $\mathcal{Y}_{f}$.
-We verify that
-(5.3) $\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}_{f}}(r)\}\geq c_{0}(\mathfrak{X}_{\sigma}^{+}(r)+\log r)$
-holds, where $c_{0}$ is a sufficient large constant determined in Theorem 22. Indeed, if it is not true, then there exists a constant $c>0$ and a sequence $r_{n}\rightarrow+\infty$ such that
-$\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}_{f}}(r_{n})\}\leq c(\mathfrak{X}_{\sigma}^{+}(r_{n})+\log r_{n}).$
-Using (ii) of Proposition 8, (5.2) and the fact that $\log r=o(T_{f}(r))$, we have that
-$T_{\tilde{u},\mathcal{Y}_{f}}(r_{n})\leq{c(n+2)\over\ell}T_{\tilde{u},\mathcal{Y}_{f}}(r_{n})+o(T_{\tilde{u},\mathcal{Y}_{f}}(r_{n})),$
-which gives a contradiction when $\ell>c(n+2)$.
-Hence (5.3) holds.
-Therefore, by Theorem 22, there exist a proper Zariski closed subset $W$ of $\mathbb{A}^{n}$ and a positive constant $c$, which are independent of $f$, such that the following two statements hold
-(5.4) $N_{\tilde{G}({\mathbf{u}}),\mathcal{Y}_{f}}(0,r)-N^{(1)}_{\tilde{G}({\mathbf{u}}),\mathcal{Y}_{f}}(0,r))\leq_{\operatorname{exc}}\frac{\epsilon}{3}\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}_{f}}(r)\}.$
-(5.5) $N^{(1)}_{\tilde{G}({\mathbf{u}}),\mathcal{Y}_{f}}(0,r))\geq_{\operatorname{exc}}\left(\deg G-\frac{\epsilon}{2}\right)\cdot\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}_{f}}(r)\}.$
-Here, $\tilde{G}\in\mathbb{C}[x_{1},\ldots,x_{n}]$ such that $\tilde{G}(x_{1},\ldots,x_{n})=G(1,x_{1},\ldots,x_{n})$.
-Therefore, $\tilde{G}({\mathbf{u}})=G(\tilde{u})$.
-Since $\widetilde{f}$ is the restriction of $f$ to $\mathcal{Y}_{f}=\mathbb{C}\setminus{f^{-1}(D)}$ and $\tilde{f}_{i}$ has no zero on $\mathcal{Y}_{f}$, it is obvious that
-(5.6) $N^{(1)}_{\tilde{G}({\mathbf{u}}),\mathcal{Y}_{f}}(0,r)=N^{(1)}_{G(\widetilde{f}),\mathcal{Y}_{f}}(0,r)\leq N^{(1)}_{G(f)}(0,r).$
-Then it follows immediately from (5.5) that
-$N^{(1)}_{G(f)}(0,r)\geq_{exc}\left(\deg G-\frac{\epsilon}{2}\right)\cdot\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}_{f}}(r)\}\geq(\deg G-\epsilon)T_{f}(r)$
-by (ii) of Proposition 8 and the fact that $\log r=o(T_{f}(r))$. This proves (ii).
-To show (i), we apply (5.6) again and obtain the following:
-$N_{G(f)}(0,r)-N^{(1)}_{G(f)}(0,r)$ $\leq N_{G(f)}(0,r)-N^{(1)}_{\tilde{G}({\mathbf{u}}),\mathcal{Y}_{f}}(0,r))$
-$=N_{G(f)}(0,r)-N_{\tilde{G}(\mathbf{u}),\mathcal{Y}_{f}}(0,r)+N_{\tilde{G}({\mathbf{u}}),\mathcal{Y}_{f}}(0,r)-N^{(1)}_{\tilde{G}(\mathbf{u}),\mathcal{Y}_{f}}(0,r))$
-$\leq_{\operatorname{exc}}\deg G\cdot T_{f}(r)-\left(\deg G-\frac{5\epsilon}{6}\right)\cdot\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}_{f}}(r)\},$
-by the First Main Theorem, Theorem 6.4, (5.4) and (5.5). Again as $\log r=o(T_{f}(r))$, it yieids
-(5.7) $N_{G(f)}(0,r)-N^{(1)}_{G(f)}(0,r)\leq\epsilon T_{f}(r)$
-by (i) of Proposition 8. Finally, we can conclude the proof by taking $Z$ to be the Zariski closure of $W$ in $\mathbb{P}^{n}({\mathbb{C}})$. ∎
+Let $\tilde{G}\in\mathbb{C}[x_{1},\ldots,x_{n}]$ be defined by $\tilde{G}(x_{1},\ldots,x_{n})=G(1,x_{1},\ldots,x_{n})$.
+Then $f_{0}^{\deg G}\cdot{\tilde{G}}({\mathbf{u}})=G(f).$ Hence it is obvious that
+(5.4) $N^{(1)}_{\tilde{G}({\mathbf{u}}),\sigma}(0,r)=N^{(1)}_{G(f),\sigma}(0,r).$
+Let $\epsilon>0$, and let $W$ be the proper Zariski closed subset $W$ of $\mathbb{A}^{n}$ determined in Theorem 26.
+We express the zero locus of $W$ as the vanishing set of a finite collection $\Sigma\subset\mathbb{C}[x_{1},\ldots,x_{n}]$, and assume that $H({\bf u})\neq 0$ for all $H\in\Sigma$.
+Then, by Theorem 26, together with (5.1) and (5.4), we obtain, after shrinking $r_{j}$ if necessary, that there exists a positive real $c_{0}$ (depending only on $\epsilon$, $n$, and $\deg G$) such that
+(5.5) $N_{G(f),\sigma}(0,r)-N^{(1)}_{G(f),\sigma}(0,r)\leq_{\operatorname{exc}}\frac{\epsilon}{3}\max_{1\leq i\leq n}\{T_{u_{i}}(r)\}+c_{0}(\mathfrak{X}_{\sigma}^{+}(r)+\log r),$
+and
+(5.6) $\left(\deg G-\frac{\epsilon}{2}\right)\cdot\max_{1\leq i\leq n}\{T_{u_{i}}(r)\}\leq_{\operatorname{exc}}N^{(1)}_{G(f),\sigma}(0,r)+c_{0}(\mathfrak{X}_{\sigma}^{+}(r)+\log r).$
+Using (5.3) and the fact that $\log r=o(T_{f}(r))$, by taking $\ell>3c_{0}(n+2)\epsilon^{-1}$, we have that
+(5.7) $c_{0}(\mathfrak{X}_{\sigma}^{+}(r)+\log r)\leq{\epsilon/3}\max_{1\leq i\leq n}\{T_{u_{i}}(r)\}.$
+Since $B_{t}^{\sigma}\subset\mathbb{D}_{t}$ for any $t\geq 1$, $N^{(1)}_{G(f),\sigma}(0,r)\leq N^{(1)}_{G(f)}(0,r)$. It follows immediately from (5.6) and (5.7) that
+(5.8) $N^{(1)}_{G(f)}(0,r)\geq_{\rm exc}\left(\deg G-\frac{5\epsilon}{6}\right)\cdot\max_{1\leq i\leq n}\{T_{u_{i}}(r)\}\geq(\deg G-\epsilon)T_{f}(r).$
+This proves (ii).
+It remains to prove (i), which follows from (ii), i.e., from (5.8) together with the First Main Theorem.
+$N_{G(f)}(0,r)-N^{(1)}_{G(f)}(0,r)$ $\leq_{\operatorname{exc}}\deg G\cdot T_{f}(r)-(\deg G-\epsilon)T_{f}(r)\leq_{\operatorname{exc}}\epsilon T_{f}(r).$
+Finally, we can conclude the proof by taking $Z$ to be the Zariski closure of $W$ in $\mathbb{P}^{n}({\mathbb{C}})$. ∎
 6. Toric Varieties
 Proof of Theorem 3.
-We recall the following setup of finding a natural finite open covering of $X$ from the proof of [13, Theorem 4.4].
+We recall the following setup of finding a natural finite open covering of $X$ from the proof of [12, Theorem 4.4].
 Let $\Sigma$ be the fan corresponding to the smooth projective toric variety $X$.
@@
 If $z_{0}\in\mathbb{C}$ is a pole of some $u_{\eta,i}$, $1\leq i\leq n$, then $f(z_{0})$ is in a component of $D_{0}$, the boundary of $X$, with multiplicity
-$-\mathop{\mathrm{min}}\{0,v_{z_{0}}(u_{\eta,1}),\ldots,v_{z_{0}}(u_{\eta,n})\}\geq\ell.$
+$\mathop{\mathrm{min}}\{0,v_{z_{0}}(u_{\eta,1}),\ldots,v_{z_{0}}(u_{\eta,n})\}\geq\ell.$
 Let $\widetilde{{i}_{\eta}}:X_{\eta}\to\mathbb{P}^{n}({\mathbb{C}})$ be the composition of $i_{\eta}$ with the natural embedding of $\mathbb{A}^{n}$ to $\mathbb{P}^{n}({\mathbb{C}})$.
@@
 Furthermore, the pullback $(\widetilde{{i}_{\eta}}^{-1})^{*}(D|_{X_{\eta}})$ of $D$ to $\mathbb{P}^{n}({\mathbb{C}})$ is defined by some nonconstant homogeneous polynomial $F_{\eta}\in\mathbb{C}[x_{0},\ldots,x_{n}]$ with no monomial and repeated factors and $\deg_{X_{i}}F_{\eta}=\deg F_{\eta}$ for each $0\leq i\leq n$ since $D$ is reduced and is in general position with the boundary of $\mathbb{G}_{m}^{n}$ in $X$.
-By [25, Proposition 12.11], there exist a non-zero constant $c_{\eta,A}$ and a proper closed subset $Z_{\eta,A}\subset X_{\eta}$, depending on $\eta$ and $A$ such that
+By [27, Proposition 12.11], there exist a non-zero constant $c_{\eta,A}$ and a proper closed subset $Z_{\eta,A}\subset X_{\eta}$, depending on $\eta$ and $A$ such that
 (6.1) $T_{\widetilde{{i}_{\eta}}(f)}(r)\leq c_{\eta,A}T_{A,f}(r)+O(1)$
@@
 (6.13) $T_{R,f}(r)\leq 2\epsilon T_{\pi^{*}A,f}(r)+O(1).$
-Since $R$ is a big divisor on $Y$, by [25, Proposition 12.11], there exists a constant $c$ and a proper Zariski-closed subset $W_{0}$ of $Y$, depending only on $A$ and $R$, such that such that
+Since $R$ is a big divisor on $Y$, by [27, Proposition 12.11], there exists a constant $c$ and a proper Zariski-closed subset $W_{0}$ of $Y$, depending only on $A$ and $R$, such that such that
 (6.14) $T_{\pi^{*}A,f}(r)\leq cT_{R,f}(r)+O(1),$
@@
 ∎
+Acknowledgements.
+We sincerely thank the referees for their careful reading of the manuscript and for their many valuable comments and suggestions.
+We also thank the first author’s student, Zhe Wang, for helpful discussions and valuable suggestions.